A Complete Bibliography of Publications in the ACM Symposia on Theory of Computing (STOC) for

2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 2.15

Title word cross-reference

(1 + ϵ, β) [EP01]. (δ + 1) [BE09]. 0 [KKMR06, MNRS08]. 1 [Kho02b]. 2 [BRSW06, DS05b, Gar05, Hås05, KdW03, Kho02b]. 2 − ϵ [Hol02]. 2^{k}(ln 2 − O(k)) [AP03]. 3 [AM02, Ale05, AK07a, DS05b, GW01, IH01]. 4 [AM02, Hol02]. 4.5n − o(n) [LR01]. ^{0} [JKS02, She07]. A [Pap07]. δ [BE09]. e/(e − 1) [KKR01]. ϵ [AES09, CGKK04, Cla06, GST^{+}01, OSS00]. H [GKP02]. k [AP03, Ale05, AAK^{+}07, AGK^{+}01, BCR01, CVV02, CGH^{+}04, CMP08, FL08, Gar05, HPM04, ITT03, KN04, Ros08, Tho08]. L [OSS00]. L_{1} [Rab03, Ind07]. l_{2} [Ind07]. L_{p} [BP03]. log^{∗}(n) [CGH^{+}04]. n^{3−Δ} [VW06]. n · 2^{o(\sqrt{(lg lg n)})} [CP07]. ≠ 2 [CCL08]. O [Bas07]. o(log^{2} k) [FL08]. O(log n) [BOPV06]. O(log n log log n) [Tri05]. O(n log log n) [Han02]. O(n log n) [Ajt05]. O(\sqrt{log n}) [ACMM05]. O^{∗}(1/ϵ) [BI04]. Ω(2^{n/4}) [BW01]. ±1 [TV05]. t [Gaf09]. Õ(mn) [HKPB07]. Ω̃(log n) [CKPR01].

-approximation [FL08, Gar05]. -bit [OSS00]. -center [CGH^{+}04]. -clique [Ros08]. -clustering [BCR01]. -CNFs [Ale05]. -colorable [AM02]. -coloring [BE09]. -colouring [GKP02]. -CSP [Hås05].

-CUT [GW01]. -DNF [Ale05]. -extension [KKMR06, MNRS08]. -hard [CGH^{+}04]. -means [HPM04]. -median [AGK^{+}01, HPM04]. -minimal [Bas07]. -MST [Gar05]. -nets [AES09, Cla06]. -node [KN04]. -optimization [OSS00]. -paths [Pap07]. -prover [Kho02b]. -query [KdW03]. -regular [Hol02]. -resiliency [Gaf09]. -restricted [ITT03]. -round [Kho02b]. -SAT [AP03]. -server [CMP08]. -source [BRSW06]. -spanner [EP01]. -vertex [CVV02, FL08]. -way [Tho08]. -wise [AAK^{+}07, ITT03].

’05 [ACM05]. ’07 [ACM07]. ’08 [ACM08]. ’09 [ACM09].

2 [Aus07]. 2-dimensional [Pat07]. 2-query [IKW09]. 2-round [DS02b]. 2-SAT [Aus07]. 2001 [ACM01]. 2002 [ACM02]. 2003 [ACM03]. 2004 [ACM04]. 2005 [ACM05]. 2006 [ACM06]. 2007 [ACM07]. 2008 [ACM08]. 2009 [ACM09]. 2CNF [ACMM05].

3 [Ach00, OW09]. 3-CSPs [OW09]. 3-manifold [AHT02]. 3-query [Efr09, GR07, Yek07]. 3-SAT [Ach00]. 33rd [ACM01]. 36th [ACM04]. 37th [ACM05]. 39th [ACM07]. 3CNF [BSHR03].

40th [ACM08].

abstract [Ach00, ASV00, AKK^{+}08, BBB^{+}00, CGGM00, CV00, CG09, CFG^{+}00, CW09a, CS00, DSY00, FKK^{+}00, FKN00, GS00a, GV00, HT04, Ist00, JKN08, KNT00, KPT09, MV00, MPSV00, NRR00, OSS00, PV00, Pei09, RW00b, Sch00, Vyg00]. AC [JKS02, She07]. Access [Gol09]. accessibility [MPSV00]. achieving [GR06, Kat07, PS04]. acknowledgement [KKR01]. ACM [ACM00, ACM01, ACM02, ACM03, ACM05, ACM06, ACM07, ACM08, ACM09, ACM10]. across [Ist00]. Adams [Dan07, CMM09, MS09b]. Adaptive [AK04, FRV06, GGP05]. Additive [BKN08a, LS08]. adequate [PS06]. Adiabatic [ATS03, Rei04]. admit [PR07]. advance [Fis04]. Advances [ABN06]. advantage [HV02]. adversarial [AKK02, MNS08, Sch05]. adversaries [HLS07]. advertising [Rag06]. Affiliation [LS09a]. Affine [BSK09, KN09]. after [Amb04]. against [Fis04, TS09, Wat06]. agents [ADTW03]. Aggregating [ACN05]. agnostic [KMV08]. Agnostically [GKK08]. agreement [Aar05, BOH05, BOPV06, GGP05, Hol05, LLR02]. algebra [CW09b]. Algebraic [KS08a, AK02, BPR05, Bas05, BC04]. algebras [CRS02]. Algebrization [AW08, IKK09]. algorithm [AJL06, AKS01, Ajt03, ABD^{+}02, AS07, Bas05, BEK^{+}03, CFJ^{+}09, CGKM09, CLL^{+}08, CH03, CD02, CJK^{+}00, CS00, DV04, EK02, FL08, GT00, GS00c, HKPB07, Imp06, IFF00, Jay05, JSV01, KS06, KM04, Kle00, KN04, MRS07, NDT09, OW08, Pla06, Rei04, RS08, RZ04, SV05b, ST01, Tri05, YYI09, MV00]. Algorithmic [Siv02, CCD^{+}03, Dev04, Gur06, NdM06]. Algorithms [CKZ01, DK08, Ind04, Pap01, ACMM05, AS04, AMS00, Amb04, AKK02, ADS09, BBCM04, BBN08, BYKS01, BHS02, BKR04, BGK^{+}07, BJR07, CL07, Cha07, Cha02b, COP03, CMM06, CLM03, CRS03, CD05, CVV02, CV05, CGM^{+}09, CF08, DS08, DKM06, DNS05, DNR^{+}09, FHL05, Fel08, FM00a, FFV05, FI00, GK04, GGI^{+}02b, GSTV07, GW01, GSVV01, GM07, GKR03, GPRS04, GS00b, Hal02, Hal05, JV01, JV02, JV07, KKL07, KR00, Kon02, LRS04, LRS06, Liu09, Lyn07, MR01b, Pla03, Rag06, Rag08, Sei00, ST01, ST04, Tal04, Vyg00, Wat01]. alignment [LMW00]. all-or-nothing [CKS04a]. All-pairs [VWY07, BHS02, Cha07, Tho05]. allocation [AS07, BNBYF^{+}00, BCG09, BKK^{+}03, wCG06, CKV02, JKK^{+}01]. allocations [BCSV00, JV02, TW07]. allows [Hås05]. Almost [AM02, CKL^{+}04, DW03, Vöc01, AAK^{+}07, BATS08, HR07b, MR06, RZ04]. almost-linear [MR06]. almost-Ramanujan [BATS08]. Alon [Fri03]. alone [LPV09]. Alpha [DGJ03]. Alpha-shapes [DGJ03]. alphabets [GI02]. alternative [OSS00]. AM [SU07]. among [GP06]. amortized [ST00]. amount [GR03]. amplification [AALV09, Din06, KW00, LP09, O’D02, SV08a, Tre05]. amplify [HVV04]. analysis [AFK^{+}01, Bac02, BCDR07, CLTW02, KM04, NRS07, O’D08, SSS02b, ST01]. analytic [APU01]. anarchy [BHLR08, CK05c, Rou02, Rou09]. Annual [ACM00, ACM01, ACM02, ACM04, ACM05, ACM06, ACM07, ACM08]. Anti [NT01]. Anti-persistence [NT01]. any [CCL08, HR07a]. application [Kle06, OS00]. Applications [JV01, AHKV03, AŠdW06, ABD^{+}02, BYKS01, BSGH^{+}04, Cha02a, CS00, FM00a, GZ03, GKK^{+}07, GV00, Gur06, HK04, IKOS04, JKS03, Nan08, PW08, VW06]. approach [AK07b, BNBYF^{+}00, BLR08, CM09, CS00, DI03, GGH^{+}08, Gur06, IKK09, JMS02]. approaches [AK04]. approximability [PV00, Tre01]. Approximate [AC06a, AJKS02, BHPI02, Dye03, MS00, Oba04, TZ01, AMM02, CD05, CGH^{+}04, CG04, DS08, DMR06b, Fel06, GGI^{+}02b, Hol02, Kho02a]. Approximately [CDR05, CD02]. Approximating [AN04b, AO09, BCR01, CLL^{+}02, FHK00, FKN00, Für00, iKM06, SL07, AJL06, Ajt03, BNBYF^{+}00, BDKR02, BEK^{+}03, CRS02, FRT03, GS00c, KR03]. Approximation [AMS00, BBCM04, BKV05, CK02, CVV02, CV05, DNS05, GW01, GM07, KN04, MR01b, Pla03, dlVKKR03, ACMM05, AAC07, AdlVKK02, AKR03, AC06b, AS07, BGK^{+}07, BCNW06, CRS03, DV07, EK02, Elk04, FL08, Fei02, FHL05, FFV05, Gar05, GMM01, GT00, GKR03, GPRS04, GK09, Hås05, HKKN01, JV01, JSV01, KKL07, KS09a, KSY00, KR00, LS08, NDT09, RS06a, SS02, Shp01, Von08, YYI09, dlVKKV05]. approximations [AM01, EGS00, IW05, JLN^{+}05]. arbitrary [FRT03, RS01, SK01]. arborescences [BCG09]. areas [Wig04]. argument [KdW03]. arguments [Aar04a, Amb00]. arithmetic [DKSS08, DSY08, Raz08, SW03, Shp07]. armed [KSU08]. arrangement [DKSV06]. arrangements [Bas02]. arrays [GV00]. arrivals [ASS02]. art [CL07]. Arthur [PV07, San07]. Artin [Gur09]. assembled [ACGH01, RW00b]. assembly [ACG^{+}02]. assignment [HV02]. assisted [NS02]. assumptions [Lin03]. Asymmetric [CGH^{+}04]. asymptotic [AEM05]. asynchronous [AC07, CKS04b, CK05b, MRR01]. Athena [Gol09]. attachment [BCDR07]. attack [KS09b]. attacks [Wat06]. attend [Wig04]. attribute [KKM05]. Auction [GK04]. auctions [AFG^{+}05, CM09, DNS05, DNS06, FGHK02, Rag06]. auditors [DS02b]. augmentation [CGKM09, CGKK04]. authenticated [LLR02]. Automata [Tra01, BC05, Vat01]. automorphisms [Gur09]. auxiliary [DKL09]. Average [Bac02, AM02, BL01, Fei02, Fei04, FFV05, GK06, PR07, Rab03, SSS02b]. average-case [PR07]. avoiding [BK09]. axiomatic [IKK09]. axis [AES09]. axis-parallel [AES09].

back [FKK^{+}00]. backup [AK07a].

Balanced [Aus07, BSW05, BCSV00, KT08, NS05, TW07, FM06]. balancing [AKK02, FS09, KL03a]. Baltimore [ACM05]. band [LMW00]. bandits [KSU08]. barrier [AW08]. based [DV07, DN07, HR07b, LRS06, LS07, Reg03, RS08, SA02]. basing [AGGM06, AGGM10]. Batch [IKOS04]. batched [Bac02]. be [Aus07, BSW05, Imp06, Nor06, Val01]. beating [BRSW06]. became [Hau08]. behavior [Ajt03, GK03]. being [DS02a]. benefit [JKK^{+}01]. Berlekamp [KS09a]. ’best [Kat07]. Bethesda [ACM09]. Better [COP03, FM00a, Gur04, GKR03]. Betti [Bas02, BPR05, Bas05]. between [AJPU02, Fei02, KRS04, WW01]. Beyond [For05, DGKS08, JKS02]. BG [Gaf09]. BG-simulation [Gaf09]. bias [ACKM05]. Biased [ESSS01, BSSVW03, DS02a]. bidders [DNS05]. bin [CJK^{+}00]. binary [CMP08]. bipartite [CG04, MV00]. bisection [FKN00, HK04]. Bit [Vio09, ATSVY00, BSW05, Hol05, OSS00]. Bit-probe [Vio09]. bits [Ajt05]. bitvectors [BMRV00]. Black [CKPR01, IKLP06]. Black-box [CKPR01, IKLP06]. block [MS00]. blocking [RW00a]. boat [BP03]. Boolean [BSW05, KST02, LR01, O’D08]. Boosted [GPRS04]. Boosting [KS03, KMV08]. both [Kat07]. bottleneck [VWY07]. bound [Aar02, Ach00, AŠdW06, BW01, CCL08, DGKS08, FRT03, FFR03, GGP05, IH01, KdW03, LR01, Rao08]. Bounded [GKRdW06, KPT05, Lin03, Pas04, Ant06, BKN08a, BCG09, BS02a, CG06, CDR05, DFKO06, DSY08, DM02, Kel04, KR00, LS08, RS01, SL07, Tre01]. Bounded-concurrent [Lin03, Pas04]. Bounded-depth [KPT05]. Bounded-error [GKRdW06]. bounded-storage [DM02]. bounds [Aar04a, AC04, Ale05, Amb00, Amb01, AT00, AC07, AHW08, BYKS01, BY03, BR00, BJR07, CCM08, CL01, CKL^{+}04, CHKX04, CK08, DW09a, Elk04, GR03, GGK03, Gur06, HR00, JKN08, JKKR03, KI03, Kel04, KR03, Kon02, LS07, OS03, PD04, Pat07, PR01b, RS01, Raz02b, Raz08, SS02, San07, SW03, She08, Vio09, Zha06]. box [CKPR01, IKLP06]. boxes [AES09]. branching [BW01]. breadth [Wig04]. British [ACM08]. broadcast [EK02, FM00b, KSY00]. Brouwer [CD05]. budgeted [GM07]. Buffer [KLM^{+}01]. buffering [AS04]. buffers [ERW07]. Building [Cla06]. bulk [CK05a]. burning [HR08]. buttons [FKK^{+}00]. buy [CK05a, FKLS06]. buy-at-bulk [CK05a]. Bypassing [Tal04]. Byzantine [BOH05, BOPV06, LLR02].

CA [ACM03]. Cache [ABD^{+}02, BF03].

Cache-oblivious [ABD^{+}02].

cache-obliviousness [BF03]. caching [BBN08]. calculus [ABSAW00]. California [ACM07]. Cambridge [ACM10]. Can [Imp06, BSW05, Val01]. Canada [ACM02, ACM08]. candidate [CH02]. cannot [Gav08]. capacitated [Pap07]. capacity [GR06]. capacity-achieving [GR06]. cares [CGL04]. cascading [CL01]. case [Ajt03, AT00, Bac02, BCSV00, CKS04b, Fei02, FFV05, Hol07, Mic02, PR07, Pei09, SSS02b, TW07, Tho05]. case/ average [Mic02]. catch [MT09]. Cauchy [OS00]. Cauchy-like [OS00]. Cayley [RSW04]. Cell [JKKR03, BR00, CDR05, Pag01]. cell-bounded [CDR05]. Cell-probe [JKKR03]. center [CGH^{+}04]. certain [GS00b]. chains [MS01b]. changing [Rag06]. channel [NS02]. characteristic [CCL08, Uma08]. characterization [Gaf09, KY00, ST00, Sha06]. cheater [Ant06]. checking [Ajt02, GGH^{+}08]. Chicago [ACM04]. choosing [GKP02]. Chow [OS08]. chromatic [AN04a, AC06b, iKM06, Zuc06]. circle [Kel04]. Circuit [KC00, San07, AACW06, KI03]. circuits [DS05b, DSY08, KPT05, LR01, RS01, Raz08, She07, Shp07, Tra01]. Clairvoyant [Gác02, BL01]. class [GT00, Hal05]. classes [BC04, FST05, iKM06, San07]. Classical [Gav08, Gur03, BYJK04, JKN08, SS04, Shi05]. classically [Val01]. classification [GT00]. Claus [BS06]. cleaning [ELM^{+}00]. Clifford [CRS02]. Clique [FRRS06, Ros08, Sri00, Zuc06]. Clique-width [FRRS06]. closed [BSS08, iKM06]. closure [BHS02]. cluster [CP01]. Clustering [CP01, Sch00, ACN05, BHPI02, BBV08, BCR01, COP03, HPM04, dlVKKR03]. CNFs [Ale05]. code [OW08]. codes [DS05b, Efr09, Elk01, GKZ08, GGR09, GS00b, Gur02, GI02, GI03, Gur04, GR05, GR06, Gur09, IKOS04, KT00, KdW03, Mei08, MW02, Reg05, TSZ01, Tar03, Yek07]. coding [BSGH^{+}04, GKY02]. cohomology [CEN09]. coin [Amb01]. collaborative [KS04]. Collective [CKS04b]. collision [Aar02]. Collusion [LMas05, HTW06]. Collusion-free [LMas05]. colorable [AM02]. coloring [BE09, CG04, DMR06b, Hay03, HVV07, Kho02a]. colors [HVV07]. Colouring [MR01a, GKP02]. colourings [Mol02a]. colours [MR01a]. Columbia [ACM08]. Combinatorial [ACG^{+}02, Bas07, EK02, Mei08, AK07b, BATS08, CGP08, DNS05, DNS06, IFF00, OSS00, Sha06]. Combining [GMP00]. commitment [DG03, HR07a]. commodity [KRV06, OSVV08]. Communication [BIW08, NN01, BYJK04, BV02a, CCM08, GKRdW06, GKK^{+}07, HM07, JKN08, JS08, Kla00, KNTSZ01, KW09, LS07, NS02, She08, Shi05]. comparison [MS00]. Compatible [Gác01]. Competitive [DKR02, FGHK02, BBN08, CGKM09, GRS03, Kon02]. complement [DNS05]. complement-free [DNS05]. Complete [GCKL08, KY00, AHT02]. Completeness [HNRR04, Kil00, Ist00]. completion [HPTV07, SSS02b]. Complex [DLS01, Für00, GW01]. complexities [SS04]. complexity [Aar05, AW08, ABM01, ABSAW00, APU01, BYJK04, BKN08b, Bas07, BDKR02, BV02a, BSS05, BK08, BKJ01, BC04, CLL^{+}02, CK08, CDD00, Dan07, DGP06, DPS02, DI06, DNR^{+}09, FPT04, Fei02, GGP05, GKRdW06, GKK^{+}07, GIKR01, GKP02, Gur03, HM07, IH01, JKN08, JKS03, KNTSZ01, KW09, LS07, Pag01, Raz02a, RW03, Ros08, RW00b, ST00, SV05a, Shi05, Siv02, Vad00, Yao01]. components [BPR05]. composability [LPV09, PS04]. composable [CLOS02]. composition [KLP05, KLR06, LLR02, Uma08]. Compressed [GV00, GSTV07]. Compression [AL00, FS08]. computable [Ajt04, BY07]. computation [AM01, CLOS02, CCM08, CDD00, CGS02, GR03, GKR08, GCKL08, HRTS00, HT04, HNRR04, IKLP06, IKOS07, Kat07, Kil00, Lin03, Pas04, SV05b, vAHL05]. computational [CHKX04, HNRR04, IKOS08, RW03, Yao01]. computations [BC04, For00]. computers [Val01]. Computing [ACM00, ACM01, ACM02, ACM03, ACM04, ACM05, ACM06, ACM07, ACM08, ACM09, ACM10, ASV00, Bas02, BPR05, FMP^{+}00, GLMM04, Gro01, Hau08, iKR07, Pap05, S̊O01, Aar04b, Bas05, CH03, DGP06, FIM^{+}03, GRS03, Gro00b, Hal05, Lyn07]. concatenated [GS00b]. concave [EdMN09]. concentration [Rao08]. Concurrent [Gol02, KLP05, KP01, CKPR01, LPV09, Lin03, Pas04]. condensers [BKS^{+}05, TSUZ01]. condition [VT07]. Conditional [DMR06b, OW09]. Conditions [MRR01]. conductors [CRVW02]. confluent [CRS03, CKL^{+}04, OS00]. congestion [AZ05a, AZ06, CKS06, CK05c, CN04, CGKT07, HTW06, KPT09, Räc08]. conjecture [Ajt04, Fri03, GOS08, iKR09, KS09b, LMS08, Sha09, Smy02]. conjectures [Gur06]. conjugate [DH07]. conjunctive [GSS01]. connected [BPR05, CVV02, FL08, KN04]. connection [Mic02, PR07]. Connectivity [KKK00, CCK08, Cha02a, CH03, PD04, Rei05, Tho00, Tri05]. connects [RW00a]. consensus [ASS02, AC07, MRR01]. consequences [NdM06]. consistency [FM00b]. Consistent [KL03a]. Constant [AKR03, CRVW02, CKS06, CD02, GGH^{+}07, GMM01, GT00, GK09, IKOS08, KK03, LRVW03, MR06, OP03, PPR07, Rao06, Ros08, Tho03a, YYI09]. constant-degree [CRVW02]. constant-depth [Ros08]. constant-factor [GK09]. constant-time [YYI09]. constrained [DSVW07, MR01b]. constraint [ART06, AKK^{+}08, BK08, BKJ01, Dev04, Mol02b, dlVKKV05]. constraints [LNSS07, LMNS09]. Constructing [BY07, DH07]. construction [BRSW06, BATS08, CH00, HKPB07, Mei08, RS09]. constructions [BKS^{+}05, EP01, GPV08, IKLP06, PR05, Reg03]. constructive [GGH^{+}08, Mos09]. contingency [CD02, CDR05]. continuous [CGM^{+}09, S̊O01, Tra01]. continuous-time [CGM^{+}09, S̊O01]. control [LRS06]. Controlling [Gol09]. convergence [EdMN09, FRV06, KLE07]. converging [CF08]. Convex [AE05, BV02b, IH01]. convolution [BHKK07]. Cooperative [CK05b, GRS03, JV01]. core [BHPI02]. core-sets [BHPI02]. Coresets [FS05, FFKN09, HPM04]. corner [LV04]. Correcting [DS05a, KT00]. correlated [Pap05]. coset [FIM^{+}03, HMR^{+}06]. cost [CVV02, FPS00, FKLS06, JV02, RS06b, Sch00, Vyg00]. cost-sharing [RS06b]. costs [CV05, GKY02, KKM05]. count [CD02]. Counting [BC04, Wei06, AJKS02, BGK^{+}07, CLX09, CDR05, Dye03, Pat07, VW09]. cover [AAA03, DGKR03, Hol02, JLN^{+}05, STT07]. covering [AK07a]. Covert [vAHL05]. Crawling [CF02]. Crete [ACM01]. critical [AJS07, KN04]. cross [RW00a]. cross-connects [RW00a]. crossing [Gro01, iKR07]. crossings [EGS00]. CRT [Bon00]. cryptographic [GPV08, Mic02, PR05, Reg03]. Cryptography [IKOS08, DKL09, DW09b, GKK^{+}07, Reg05]. cryptosystems [Pei09]. CSP [AdlVKK02, CLX09, Hås05, Rag08, Tul09]. CSPs [OW09]. cube [DFKO06]. cubic [AK07a, VWY07]. curve [AMT06]. curvelet [Liu09]. curves [AK02, CS03, Elk01]. cut [ACMM05, AAC07, AN04b, ALN05, CK06, CK07, DKSV06, MNRS08, OW08, STT07, Tho01, Tre09, FS01, GW01]. cut-norm [AN04b]. cuts [AKR03, AP09, BKL07, CEN09, Tho08]. Cutting [AKS03]. cycles [AKS03, FMS00, Gab04]. cyclotomic [Gur09].

D [AK07a, IH01]. Data

[GKS01, AJKS02, AK02, AFK^{+}01, CS03, DNR^{+}09, FS05, Ind04, IW05, KSY00, NT01, NRS07, SS02, Vio09]. Data-streams [GKS01]. database [BLR08]. deadline [BBCM04]. deadline-TSP [BBCM04]. decade [Amb04]. decentralized [KM04]. Decidability [SS01, MPSV00]. decidable [iKR09]. deciding [GGP05]. decision [DSY00, GKK08]. Decodability [DGKS08, Gur02]. decodable [DS05b, Efr09, GI02, GI03, GR06, Gur09, KdW03, Yek07]. decoding [ADS09, Bon00, DMT07, GGH^{+}07, GKZ08, GGR09, GS00b, GI02, GR05, KT00]. decomposition [GZ03, dlVKKV05]. decompositions [Räc08]. decrease [Tho03a]. decremental [BHS02]. definability [Gro00a]. defined [Bas05]. degeneracy [AC04]. Degree [DSVW07, AM02, ACKM05, BKN08a, BCG09, BSSVW03, Bog05, CRVW02, Fei04, HVV07, KR00, KR03, LNSS07, LS08, Lov08, MR01a, Mol02a, MR06, OS03, Sam07, SL07, Tre01, Zuc06]. degree-bounded [KR00]. Degree-constrained [DSVW07]. Delegating [GKR08]. deletion [ACMM05]. demand [CMV05]. demands [HKLR05]. Depth [Wig04, BS02a, DS05b, DSY08, GGH^{+}07, KPT05, RS01, Ros08, She07, Shp07]. depth-2 [She07]. depth-3 [Shp07]. Derandomization [AFG^{+}05, Siv02]. derandomized [Imp06, IJKW08]. Derandomizing [KI03, SW04]. descent [AK08]. design [ADTW03, BKN08a, BKV05, CCK08, CK05a, CM09, CV05, GKK^{+}01, GKR03, GKR09, HR08, LNSS07, LS08]. detectability [AALV09]. determinant [CCL08, Raz04, TV05]. Deterministic [Han02, KRVZ06, BGK^{+}07, CW09a, GS00c, Gur03, LRS04, Tho08]. diagrams [AMM02, CP07, IH01]. diameters [CP01]. dichotomy [KS09b]. dictionaries [ESSS01]. Dictionary [CGL04]. Diego [ACM03, ACM07]. Differential [DL09]. differentially [DNR^{+}09]. difficulty [Fis04]. digital [GGK03]. digraphs [RTV06]. dimension [ABR01, BKL07, CL01, CG06, DV07, MPP05]. dimensional [ABN^{+}01, CS03, DF08, DV01, Pat07, RS05, Tal04, Tok01]. dimensionality [Ist00, KL03b]. dimensions [APST03]. Direct [JKN08, AŠdW06, IJKW08, IKW09]. direct-product [IKW09]. directed [ACMM05, AAC07, AS03, AZ06, BKN08a, CLL^{+}08, CK06, CGKT07, CK07, EK02, HKLR05, RZ04]. discrete [BV04, CD05, CGM^{+}09, DFKO06]. discrete-time [CGM^{+}09]. discriminant [GS00c]. discriminative [BBV08]. dishonest [Pas04]. disjoint [AZ05b, CKS06]. disjointness [KNTSZ01]. disk [Bac02, GZ03]. dispersers [BKS^{+}05, BRSW06, BSK09]. distance [AO09, AMT06, BEK^{+}03, CW09a, IH01, KKMR06, OR05, SS02, TZ01]. distances [APST03, Sam07, WW01]. Distinct [APST03]. distortion [ALN05, BCIS05, KRS04, LMN02, OR05, Rab03]. Distributed [BE09, Lyn07, AK04, AK08, wCG06, Elk04, MRR01]. Distribution [Vat01, BBB^{+}00, Sel09]. distributions [ACKM05, BKR04, ITT03, RS05, Val08]. DLT [Sze06]. DNF [Ale05, HR02, KS01a, Sel09]. do [BC05]. does [MT09]. domains [FLN^{+}02]. domatic [FHK00]. don’t [CGL04]. doubling [CG06]. drawings [EGS00]. dual [AK07b, JV02, KR03, LRS04, Vyg00]. dualization [EGM02]. Dynamic [Cha02a, KMT03, KKR01, AT00, AJS07, AKK02, BKK^{+}03, CG06, DI03, Dye03, FS05, GRS03, Ind04, MPP05, PD04, RZ04, Tho00, Tho01, Tho03b, Tho05]. dynamics [EdMN09, Mol02a, WZ07].

early [GGP05]. earning [Mey01].

earthmover [KKMR06]. easy [AKK^{+}08]. Edge [CKS06, Gon05, MS01b, AZ05b, Ant06, CH03, CG04, GMM01, Sch00]. edge-coloring [CG04]. edge-cost [Sch00]. Edge-disjoint [CKS06, AZ05b]. edges [BK09, CDR03, HPTV07]. edit [AO09, BEK^{+}03, OR05]. Edmond [Gur03]. effect [HTW06]. effective [Nan08, SS08]. efficiency [GGK03, KT00]. Efficient [CGM^{+}09, FIS05, AL00, AMM02, BSSVW03, CFJ^{+}09, DI06, DNR^{+}09, KS01b, NV06, NDT09, PV07, Pla06, Tho03b]. eigenvalue [Fri03, Kel04, Tre09]. Eighth [ACM06]. Eisenberg [JV07]. election [Ant06]. Elusive [Raz08]. embeddable [Gro00a]. embedding [ABN06, Rab03, Tal04]. embeddings [ABN07, ARV04, BCIS05, Ind07, iKM08, OR05, RR06]. encodable [GI03]. encryption [GGK03, Gen09, KY00]. end [AK04, SU07]. end-to-end [AK04]. entanglement [Gur03, NS02]. entanglement-assisted [NS02]. entries [JSV01]. entropy [BRSW06, BDKR02, HRVW09, Rao06, Yao04]. environments [MNS08]. epsilon [BSSVW03, Gar05, RS09]. epsilon-biased [BSSVW03]. epsilon-net [RS09]. equally [OW08]. equation [Hal02]. equations [Gut00, HK04, KLE07, Pla06]. equilibria [DPS02, FPT04, FRV06, GLMM04, Pap05, SV08c]. equilibrium [CMV05, DGP06, DP09, Dev04, GK04, GP06, HM07, WZ07]. Equitable [JV02]. equivalent [DGJ03]. ergodic [Nan08]. Erratum [AGGM10]. error [GKRdW06, KT00, MR06]. error-correcting [KT00]. errors [CGL04, DS05a, KMS07, Reg05]. escrow [ATSVY00]. essentially [Sze06]. Estimating [CS04, WW01, Fei04]. estimation [Cha02b, FN05]. Euclidean [ALN05, LMN02]. Euler [Kos01]. evaluated [BSW05]. evaluating [RS08]. evaluation [GSS01, NN01]. Every [AS05, BSS08, CG09, Hås05, BSW05, Imp06, Rag08]. everything [BCDR07]. evolutionary [WW01]. Evolvability [Fel08]. Evolving [MP03, AP09]. Exact [For00, HR02, Sel09]. examples [BS02a]. Excellent [Elk01]. excess [CMV05]. exchange [KT08]. exchanging [KN08]. exclusion [AHW08]. existence [CKL^{+}04, Dev04]. Expander [ARV04]. Expanders [MW02, CRVW02, RSW04, TSUZ01]. expanding [AKK^{+}08]. expected [BV03]. Explicit [GR06, LR01, RS09, Ind07]. Exponential [BYJK04, CCD^{+}03, GKK^{+}07, KdW03, AJPU02, GKRdW06, KW00]. extended [Ach00, ASV00, AKK^{+}08, BBB^{+}00, CGGM00, CV00, CG09, CFG^{+}00, CW09a, CS00, DSY00, FKK^{+}00, FKN00, Gaf09, GS00a, GV00, HT04, Ist00, JKN08, KNT00, KPT09, MV00, MPSV00, NRR00, OSS00, PV00, Pei09, RW00b, Sch00, Vyg00]. extension [KKMR06, MNRS08]. extraction [BL02]. Extractor [TSZ01]. Extractors [ISW00, LRVW03, Rao06, Raz05, BKS^{+}05, DW09b, Gur04, Ind07, KRVZ06, TSUZ01, TS09, Zuc06].

face [Rag06]. facets [Tra01]. facility [AGK^{+}01, FFV05, JMS02, Mey01]. factor [AKR03, GMM01, GT00, GK09]. factoring [NRR00]. factorization [LS07, Uma08]. factors [HR07b, LRVW03, PR07]. failed [BK09]. fair [AS07]. fairness [wCG06, GMP00, GCKL08]. false [LMS08]. family [RSW04]. farm [JKK^{+}01]. farms [CKV02]. Fast [AM01, Ant06, BOH05, CF08, DKSS08, FRV06, GGI^{+}02b, Hal05, Uma08, AC06a, BHKK07, CH03, ESSS01, GSTV07, NDT09, Tho03b]. Fast-converging [CF08]. Faster [CH00, DS08, Für07]. Fault [CLPR09, vDMMS00]. Fault-tolerant [CLPR09, vDMMS00]. faults [FI04]. feedback [AK04, CLL^{+}08]. few [HPTV07, KMS07]. fewer [HVV07]. field [Hal05, SV05b]. fields [GKZ08, Gur09, Shp01]. Fifth [ACM03]. filtering [KS04]. find [KST02]. Finding [AP09, Bon00, FMS00, FM06, Gab04, GN08, KR02, VW06, VW09, CK08, MV00]. finger [BLM^{+}02]. fingerprint [Tar03]. fingerprinting [Yao03]. finite [CK05c, Shp01]. First [BCDR07, BPR05]. fitness [BCDR07]. Fitting [AK02]. five [Hay03]. fixed [CD05, CLL^{+}08]. fixed-parameter [CLL^{+}08]. flipping [Amb01]. flow [AAE05, BCES06, BL01, CGKM09, CKZ01, CK02, CKS04a, CGKK04, CKS05, CK07, DS08, DGJ03, FI00, GK06, GKK^{+}01, Vyg00]. flow-cut [CK07]. flow/min [Oba04]. flows [ARV04, CEN09, CRS03, CKL^{+}04, CDR05, DSVW07, KRV06, OSVV08]. folded [Gur09]. Folk [BCI^{+}08]. forbidden [LS09b]. forest [GK09]. forms [AMMN05]. formulas [Aar04b, HR02, Raz04, RS08]. Fourier [BHKK07, DFKO06, GGI^{+}02a]. Fourth [ACM02]. FPT [CHKX04]. fractal [GK03]. fractional [BI04, CL01]. framework [BBV08, LPV09]. Frankl [BRSW06]. free [ASS02, DNS05, Gut00, LMas05]. Frege [BS02a]. frequency [IW05]. full [BOPV06]. full-information [BOPV06]. Fully [Gen09, Tho01, RZ04, Tho00, Tho05]. Fully-dynamic [Tho01, Tho00, Tho05]. function [CMV05, FI00, Gur09, HR07a, Ist00, NN01]. functions [Ajt04, AGGM06, AGGM10, BBV08, BSW05, DW03, DFKO06, Fei06, HKKN01, IH01, IFF00, Mic02, NRR00, OS03, O’D08, PW08, RS09, Raz08, SW03, Vat01, Wee05].

Gale [JV07, KS09a]. game [KS09a, Sei00]. Games [KN08, AKK^{+}08, CMM06, CK05c, EdMN09, GW07, HTW06, JV01, KKL07, Kho02b, KPT09, LSZ08, Pap01, Pap05, PV07, Rao08, Tar04, dAM01]. gap [AALV09, CW09a, Dan07, Din06]. gaps [CMM09, CK07, DKSV06, MNRS08, STT07, Tul09]. gates [KPT05, RS01, Shp07, vDMMS00]. Gaussians [SK01]. general [AJPU02, BCIS05, CLPR09, CS00, EP01, ERW07, FLN^{+}02, GW07, KLP05, Kil00, SW04, VWY07]. generalized [BBN08, DS08, FGHK02]. Generating [KV03, EGM02]. generation [ATS03]. generators [Bog05, ISW00, Lov08, Uma02]. generic [KPT09]. genomes [WW01]. genus [AHT02, Kel04, MV00]. geometric [AMS00, ARV04, AK04, Cha02a, CLM03, FS05, GRK04, Ind04, MS01b]. geometry [ART06, Bas07, LS09b]. Girth [LMN02, Hay03, Mol02a]. given [Fis04]. Glauber [Mol02a]. global [FM00b]. Gomory [HKPB07]. goods [AS07]. gossip [KKD01]. Gowers [LMS08, ST06]. gradient [AK08, DH07]. grammar [CLL^{+}02]. Graph [BCL^{+}06, iKM08, KRV06, SS08, Yao04, AN04a, ACL00, AS05, ABD^{+}02, ARV04, CG09, EGS00, Fei04, Fis04, FN05, GZ03, HMR^{+}06, Mol02a, MRS07, Sha06, ST04, Tho00]. Graphs [BKN08b, AM02, ACKM05, AAKV01, ACL00, AS03, AMMN05, AKR03, AK07a, AKK^{+}08, BCES06, BKS^{+}05, BRSW06, BATS08, BSS08, BKL07, Cha07, CLPR09, CKS06, CX00, CK08, CK06, CGKT07, CF02, CG04, DW03, EP01, FMS00, FVV08, Gar05, GRK04, Gon05, Gro00a, HKLR05, HKPB07, Hay03, HVV07, Hol02, KL02, iKM06, Kel04, KV03, Kle06, KK09, KN04, Kos01, KK03, KL03b, LS09b, MV00, MR01a, OSVV08, RZ04, SW03, SS01, SSŠ02a, VWY07]. Greece [ACM01]. greedy [AS04, DW09a, FFV05, JMS02, Tho08]. Green [Sha09]. Grothendieck [AN04b]. group [DGKS08, HRTS00, Hal05, SV05b]. groups [BBS09, FIS05, GRS03, Gut00, SW04, Wat01]. growth [KR02]. growth-restricted [KR02]. guarantee [AC06b]. guarantees [BKN08a]. guessing [Sei00].

Hadwiger [iKR09]. halfspaces [KS08b]. Hamiltonian [FMS00]. Hard [BS02a, BHW00, Ajt05, BSHR03, CGH^{+}04, FRRS06, GPV08, HKKN01, Hol02]. Hard-Potato [BHW00]. hardest [Aus07]. Hardness [AZ05a, AZ05b, CK06, CGKT07, DSY08, Fel06, Kho02a, O’D02, SV08a, AGGM06, AZ06, CN04, CK07, DGKR03, DMR06b, DNR^{+}09, HR07b, HVV04, KS08b, MNRS08, OW09, SU07, Tre05, AGGM10]. Hardness-randomness [DSY08]. hardnesses [Uma02]. hash [DW03, Mic02]. hashing [OP03, Pag01]. heaps [KST02]. heavily [BCSV00]. Hersonissos [ACM01]. heterogeneous [CDR03, RW00a]. heuristic [AGK^{+}01]. Hidden [FIM^{+}03, GSVV01, MR05]. hiding [HR07a]. hierarchical [CMP08, Pla03, Räc08]. Hierarchies [FST05]. hierarchy [AAT05, Tul09]. high [DV01, Mol02a, RS05]. high-dimensional [DV01, RS05]. Higher [HR00, APST03, CL01, CS03]. highly [ASV00]. histogram [GGI^{+}02b]. histograms [GKS01]. history [NT01]. Hit [LV04]. Hit-and-run [LV04]. Holant [CLX09]. Holographic [CL07]. homogeneous [HK04]. Homology [CEN09]. homomorphic [Gen09]. homomorphism [BKN08b, SW04]. homomorphisms [DGKS08]. homotopy [DGJ03]. hot [CLTW02]. Hu [HKPB07]. Huffman [GKY02]. human [Hau08]. hybrids [Tra01]. hyper [Hol02]. hyper-graphs [Hol02]. Hyperbolic [Gur06]. hypercontractivity [BT03]. hypercube [AHKV03]. hypercubes [Vöc01]. hypergraph [DGKR03, EGM02, Kho02a]. hypergraphs [RS07]. hypotheses [HR02].

I. [Ist00]. ideal [Gen09, Hal02]. idempotence [AMM06]. identification [GKRdW06]. identity [DS05b, KI03, KS01b, SV08b]. II [BC04]. Illinois [ACM04]. implementations [FFR03, FFR06]. implicit [BCSV04]. importance [AMM06, DS02a]. impossibility [GGP05, Lyn07, MRS07]. Improved [AAC07, BHS02, EGS00, FHL05, FI00, Mic02, ADS09, KR00, PR05, YYI09]. Improvements [BD00]. Inaccessible [HRVW09]. Inapproximability [GJ07, SV08c, HK03, Rag08, Sri00, Zuc06]. inconsistent [ACN05]. increasing [Bac02]. independence [AAK^{+}07, AH09, BKS^{+}05, PPR07]. independent [Fei04, ITT03, NT01, Rao06, Rou02, Wei06]. indexing [CGL04, GV00]. indivisible [AS07]. inequalities [Bas05, BT03]. inequality [AN04b, GN08, Smy02]. inertia [For00]. Infeasibility [FS08]. inference [KKK00]. infinite [ASS02]. influence [MR07, ST06]. Information [KLR06, ACN05, BY03, BOPV06, CFG^{+}00, CL05, DS05a, JKS03, KN08]. Information-theoretically [KLR06]. injecting [AACW06]. input [BSW05, DKL09, MRR01]. insert [ESSS01]. insert/deletes [ESSS01]. installation [GMM01]. instance [FS08]. instances [Tre01]. Integer [Tho03a, BW01, BCP01, DKSS08, For00, Für07, SW03]. integers [Bon00, GR07]. integral [CDR05, Oba04]. Integrality [CMM09, DKSV06, FS01, STT07]. intents [AGY09]. Interaction [KNTSZ01, Gav08]. interactive [BLR08, DG03, GKR08, KW00, Vad00]. Interdomain [LSZ08]. interior [DS08, KN09]. interleaved [GGR09]. International [ACM09, ACM10]. Internet [Pap01, Rex08]. Interpolation [Shp07, OS00]. intersection [CG09, CL01, KMT03, KS08b]. Interval [HPTV07]. intervals [Bon00]. intracatability [Ist00]. Intrinsic [Rou09, KL03b]. invariance [KS08a]. invariant [Sha09]. invasiveness [Ajt02]. inventory [LRS04, LRS06]. Inverse [LMS08]. inversions [AJKS02]. IP [Adl02]. Ising [Ist00]. Isomorphism [Gro00a, Fis04, HMR^{+}06, iKM08, MRS07]. isoperimetry [MS01b]. Isotopic [BCSV04]. iterations [BI04]. Iteratively [DH07].

Johnson [AC06a, DGKS08]. Jones [AJL06]. Joy [Pap02]. Julia [BY07, RW03]. July [ACM01]. June [ACM03, ACM04, ACM07, ACM09, ACM10]. juntas [Bla09, MOS03].

kangaroo [MT09]. Key [Hol05, BBB^{+}00, DW09b, KY00, Pei09, Tho03a]. knapsack [BV03, LMNS09]. knot [AHT02]. Know [MNW04]. knowledge [ATS03, CGGM00, CKPR01, DSY00, DS02b, Gol02, IKOS07, KP01, MP06, NV06, Wat06]. Kolmogorov [CLL^{+}02]. König [CG04].

L [RTV06]. labeling [KKMR06, MNRS08, NS05]. languages [AM04, BC05, BKJ01]. Large [AAE05, Sam07]. Larry [For05]. Lasserre [Tul09]. Lattice [RR06, AKS01, Ajt03, Ajt04, GN08, Reg03]. lattice-lattice [Ajt04]. Lattices [PR07, Ajt05, GPV08, Gen09, Reg05]. law [ACKM05]. leader [Ant06]. leader-election [Ant06]. leads [WZ07]. leaking [DS05a]. Learnability [JKS02]. Learning [AACW06, KKM05, KS01a, MOS03, MR05, SK01, AK04, BKW00, BLR08, Fel06, Fel08, GKK08, GM07, HR02, KMV08, KS08b, KPT09, Reg05, Sel09]. least [Hay03]. lecture [Gol09]. legacy [For05]. lemma [AALV09, CS00, Mos09, RS07]. length [Efr09, Gab04, ISW00, NH08, Yek07]. Leslie [Wig09]. less [TSUZ01]. letter [GKY02]. level [Fel06]. Liapunov [S̊O01]. lifts [BP03]. like [Gur06, OS00]. Limitations [DN07, HMR^{+}06]. Limits [Gur02, GR05, BCL^{+}06, BF03, DMT07]. Lindenstrauss [AC06a]. line [Ajt02, BCIS05, BCES06, GST^{+}01, KS09b, Rab03]. Linear [CHKX04, GI03, KS09a, NdM06, PPR07, Zuc06, AC04, AO09, AK08, BE09, CW09b, DK08, DKSV06, DV04, Gur02, GI02, Han02, HK04, KN09, iKR07, iKM08, KS06, MR06, OP03, RS09, Raz04, Reg05, RZ04, Sha09, ST04]. linear-invariant [Sha09]. linear-time [GI02]. lines [AKS03]. linked [CKS05]. links [CH00, GLMM04]. List [GKZ08, GGR09, GS00b, Gur02, GI02, GI03, GR05, GR06, Gur09, iKM06]. list-chromatic [iKM06]. list-decodable [GI02, GR06, Gur09]. List-decoding [GKZ08]. load [AKK02, FS09, KL03a, BKK^{+}03]. loaded [BCSV00]. loads [AJS07]. Local [ABN07, AGK^{+}01, MP06, Aar04a, CS00, KT00, KR03, Mos09, Rei04, SS04, Zha06]. locality [AFG02]. Locally [DS05b, AP09, Efr09, KdW03, Mei08, Yek07]. location [AGK^{+}01, FFV05, JMS02, KKD01, Mey01, Pla03]. log [BSS05, BT03, Rei05]. log-Sobolev [BT03]. log-space [Rei05]. logarithm [KP01]. Logarithmic [AZ06, EK02, PR07]. Logconcave [FVV08]. logic [Fel06]. long [FMS00, MT09, OW08]. lookahead [MNW04]. losers [BV04]. Loss [TSUZ01]. Loss-less [TSUZ01]. lossless [CRVW02]. Lossy [PW08, DS08]. Lovasz [AAT05, CS00, Dan07, Mos09, STT07]. Low [BCIS05, KRS04, OR05, Sam07, SU07, AM01, BSSVW03, Bog05, DF08, Lov08, MR06, NdM06, NDT09, Tal04]. Low-degree [Sam07]. Low-distortion [BCIS05]. Low-end [SU07]. low-rank [NDT09]. Lower [Aar04a, AC04, Ale05, BJR07, CL01, EEST05, GR03, GGK03, Kon02, LS07, PD04, Pat07, RS01, Aar02, Ach00, Amb00, Amb01, AŠdW06, AHW08, BYKS01, BY03, BCG09, BW01, CCL08, CCM08, CHKX04, DW09a, Elk04, FFR03, GGP05, HR00, JKKR03, KI03, KdW03, LR01, PR01b, Raz02b, Raz08, SS02, San07, SW03, She08, Vio09, Zha06]. lower-bounded [BCG09]. Lower-stretch [EEST05]. LP [ADS09, DMT07, STT07].

MA [ACM10]. machine [BLM^{+}02, CN04]. machines [AE05, BL01, GK06, Kon02]. maintaining [BHS02]. maintenance [GGI^{+}02b]. majority [Pas04, SV08a, She07]. make [HLS07]. malleability [LP09, LPV09]. malleable [DG03, DW09b, PR05]. Management [AR03, KLM^{+}01]. manifold [AHT02]. manifolds [DF08]. map [iKM08]. maps [KRS04]. Market [CMV05, BCDR07, CF08, Dev04, GK04, WZ07]. markets [CGP08, JV07]. marking [Adl02, AEM05].

Markov [MS01b, MR05]. Maryland [ACM05, ACM09]. massive [ACL00]. match [JKKR03]. matches [CH02]. matching [AK07a, CH02, CGL04, GV00, MV00, MS09b]. matchings [BGK^{+}07, YYI09]. matrices [For00, Für00, OS00, TV05, Vu05]. Matrix [OS00, AM01, BBS09, FN01, JSV01, NDT09, RS01, Raz02a, She08, VT07]. Matrix-vector [OS00]. matroid [LMNS09]. matroids [MS01b]. matter [KR00]. Max [Tre09, AJS07, AS07, HK04, OW08, Oba04, STT07, Zuc06, AdlVKK02, Aus07, FS01, GW01]. MAX- [GW01]. max-bisection [HK04]. MAX-CSP [AdlVKK02]. max-cut [OW08]. max-integral-flow [Oba04]. max-integral-flow/min-multicut [Oba04]. max-min [AS07]. max-weight [AJS07]. maximal [BKJ01]. maximization [BD00, LMNS09]. maximizing [Fei06, GRS09]. maximum [HVV07, MR01a, Mol02a, VW06, YYI09]. MaxMin [BCG09]. May [ACM00, ACM02, ACM05, ACM06, ACM08, ACM09, Nor06]. means [HPM04, KI03]. measurement [Shi05]. measures [PS06]. mechanical [GSVV01]. mechanics [Ist00]. mechanism [BKV05, CM09, HR08]. mechanisms [AAM03, DNS06, DN07, GRS09, RS06b]. median [AGK^{+}01, FMP^{+}00, HPM04]. Meet [CRS03]. meets [BHKK07, KR03]. Meldable [KST02]. membership [DSY00, Fel06, Pag01]. memory [Ajt02, CK05b, FI04]. merge [CRS03]. Merlin [PV07, San07]. meshing [BCSV04]. Message [ADS09, Gav08]. metarounding [CV00]. method [AŠdW06, DH07, KN09, KLE07, She08]. methods [FRV06]. metric [ABN06, ABN07, BCR01, BLMN03, CV05, CS04, ERW07, GZ03, KKMR06, KSU08, MNRS08, NS05, PR01a]. metrical [FM00a]. metrics [BCIS05, FRT03, KR02, Rab03, Tal04]. might [Aus07]. min [ACMM05, AS07, BCR01, ITT03, Rao06, Tho01]. min-cut [Tho01]. min-entropy [Rao06]. min-sum [BCR01]. min-wise [ITT03]. minima [Rei04]. minimal [Bas07]. Minimax [Tok01]. minimization [AZ05a, AZ06, BHLR08, CN04, EdMN09, Fel06, FRRS06, FI00, KC00, KS09a, KL03a, Räc08, Sch00]. minimize [BL01, CP01, CGKK04]. Minimizing [GK06, CGKM09, CKZ01, IFF00, VW09]. Minimum [Tho08, CW09a, CVV02, CS04, DKSV06, Elk04, FKN00, FHL05, KR00, RS06a, SL07, Vyg00]. minimum-cost [CVV02]. minimum-weight [FHL05]. minor [BSS08, iKM06, LS09b]. minor-closed [BSS08, iKM06]. missing [CH00]. mixed [ASV00, GS00c]. Mixing [MS09a, BT03, MS01b, MP03, Mor05]. mixture [KS04]. mixtures [SK01]. Möbius [BHKK07]. model [ACL00, BR00, BOPV06, BKW00, CW09b, Dev04, DM02, Ist00, KLP05, MS09a, SS02, Sly09, Von08]. Models [Mol02b, CLL^{+}02, KS04, LRS06, Lyn07, MR05]. Modified [BT03]. modular [DKSS08, Elk01, Uma08]. moments [IW05]. money [HR08]. monotone [AS05, BKR04, EGM02, GRK04, HR00, KLE07, LMNS09, RS05]. Monotonicity [FLN^{+}02]. Montréal [ACM02]. Mordell [GN08]. MST [Gar05, KR03]. muggles [GKR08]. muliticast [FPS00]. Muller [GKZ08]. Multi [CGKK04, KSU08, Raz04, AR03, CLOS02, CGS02, FFR03, Jay05, Pap05, Pas04, Tok01]. Multi-armed [KSU08]. multi-dimensional [Tok01]. Multi-linear [Raz04]. multi-party [CLOS02, CGS02, Pas04]. multi-player [Pap05]. Multi-processor [CGKK04]. multi-queue [AR03]. multi-writer [FFR03, Jay05]. multicast [GIKR01]. multicasting [AL00]. Multicommodity [CKS05, CK05a, CKS04a, FKLS06, GKK^{+}01]. multicut [Oba04]. multilayered [DGKR03]. Multilinear [Aar04b]. multiparty [BV02a, CDD00, HT04, IKOS07, Kat07]. Multiple [AGY09, wCG06, GMP00, LMW00]. multiplication [BW01, DKSS08, Für07, SW03, Shp07]. Multiplicative [KPT09]. multiprocessors [SA02]. multivariate [KS01b]. multiway [MNRS08]. mutual [AHW08]. myth [BCI^{+}08].

Narrow [Nor06]. Nash [DGP06, DP09, FPT04, GLMM04, HM07, SV08c]. natural [CLL^{+}02]. NC [MV00]. NC-algorithm [MV00]. Near [ADTW03, CMM06, FS09, GGI^{+}02a, GI02, LMW00, Tho00, AO09, LRS06]. near-linear [AO09]. Near-optimal [ADTW03, CMM06, GGI^{+}02a, GI02, Tho00, LRS06]. Near-perfect [FS09]. nearest [AC06a, BR00, BV02a, KR02, MS00]. nearest-neighbor [BV02a]. Nearly [ST04, BK09, Bla09, GKP02, MR01a]. Nearly-linear [ST04]. Negative [HLS07, JSV01]. neighbor [BR00, BV02a, MNW04]. neighbors [AC06a, KR02, MS00]. net [RS09]. nets [AES09, Cla06]. Network [CCK08, Tar04, ADTW03, BKN08a, CK05a, CV05, CDR03, DSVW07, GKK^{+}01, GKR03, GKR09, LNSS07, LS08, Rou02]. networks [AHKV03, AJS07, AR03, AR04, KKK00, KT08, LS09a, MNW04, MR07, Räc08, RW00a]. Newton [KLE07]. no [Hol07, KPT09]. no-regret [KPT09]. no-signaling [Hol07]. node [KN04, MR01b, Pap07]. node-capacitated [Pap07]. nodes [Sch05]. Noise [BKW00, KS03]. Noise-tolerant [BKW00]. noisy [AK02, CS03]. Non [BL01, DG03, DW09b, LMNS09, LP09, Tre01, BLR08, BY07, CK05a, Ist00, JSV01, LPV09, PR05, RW00a].

Non-approximability [Tre01].

non-blocking [RW00a]. Non-clairvoyant [BL01]. non-computable [BY07]. Non-interactive [DG03, BLR08]. Non-malleability [LP09, LPV09]. Non-malleable [DW09b, DG03, PR05]. Non-monotone [LMNS09]. non-negative [JSV01]. non-planar [Ist00]. non-uniform [CK05a]. nonabelian [GSVV01]. nonapproximability [AAT05]. nondeterminism [HVV04]. nonlocal [Shi05]. nonsingular [MR05]. nontrivial [Hås05]. nonuniform [KR03]. nonzero [Ajt03]. norm [AN04b, BP03, RR06, Vu05, LMS08]. Normal [HRTS00]. norms [LS07, Shi05]. nothing [CKS04a]. notions [KY00, PS04]. NP [AHT02, AGGM06, FRRS06, For05, FS08, HKKN01, Ist00, O’D02, ST00, SSŠ02a, Tre05, AGGM10]. NP-complete [AHT02]. NP-completeness [Ist00]. NP-hard [FRRS06, HKKN01]. NP-hardness [AGGM06, AGGM10]. number [AN04a, AC06b, BPR05, CD02, FHK00, Hal05, iKM06, iKR07, MR01a, Rao06, SV05b, VT07, Zuc06]. numbers [Bas02, Bas05, Gro01]. numeric [BC04]. Numerical [CW09b].

obfuscating [Wee05]. objectives [wCG06, GMP00]. Oblivious [HKLR05, ABD^{+}02, ACF^{+}03, DP09, GR09]. obliviousness [BF03]. odd [iKM06]. odd-minor-closed [iKM06]. off [Ajt02]. off-line [Ajt02]. offs [PT06, RS06b]. omega [dAM01]. omega-regular [dAM01]. once [BW01, SV08b]. One [ABN^{+}01, GST^{+}01, GSTV07, AGGM06, AGGM10, ABR01, AR04, BYJK04, CF08, GKK^{+}07, HR07a, Kla00, MPP05, SL07]. One-dimensional [ABN^{+}01]. one-time [CF08]. one-way [AGGM06, AGGM10, BYJK04, GKK^{+}07, HR07a, Kla00]. ongoing [CF08]. Online [DKM06, GKR09, JKK^{+}01, Alb02, AAA03, AAM03, GMP00, Kon02, Sei00]. operations [MS00]. OPT [BKK^{+}03]. Optimal [ABR01, ACF^{+}03, BLM^{+}02, CK08, DMR06a, DV01, FM07, HR08, IW05, Räc08, Rag08, SA02, Tar03, Von08, ADTW03, BK09, BMRV00, CMM06, FFR03, GGI^{+}02a, GI02, ISW00, Jay05, LRS06, LMW00, LRVW03, MOS01, NH08, OW08, ST00, SL07, Sze06, Tho00, Vöc01]. optimality [AEM05]. optimally [Bla09]. optimization [ACG^{+}02, AAM03, BV04, GPRS04, KNT00, OSS00, Rei04, Tok01, Tre01]. optimized [IJKW08]. option [DKM06]. oracle [BK09, Von08]. oracles [TZ01]. orbit [FIM^{+}03]. order [CFJ^{+}09, LNSS07]. Oregon [ACM00]. other [AHW08, GW01, ITT03, KKR01, Wig04]. outcomes [KT08]. outer [HK04]. outerplanar [Gon05]. outlier [DV01]. outperform [KPT09]. overflow [KLM^{+}01]. overhead [CKS04b, IKOS08].

P2P [MNW04]. PAC [Fel06]. packet [Adl02, AEM05, AS04]. packing [BI04, CJK^{+}00, JvS05, Pap07, Tho08]. packings [Kel04]. paging [AFG02, PS06]. pair [GZ03]. pairs [BHS02, Cha07, DI03, Tho05, VWY07]. Parallel [Hol07, Rao08, AES09, AE05, BL01, BKL07, GLMM04, Kos01, MOS01, PV07]. Parallelization [KW00]. parameter [BCL^{+}06, CLL^{+}08]. Parameters [OS08]. parametric [Tok01]. parity [BKW00, KMV08, KK09]. partial [CFJ^{+}09, DS05a, FM00b, JKKR03]. partition [Gon05, Ist00]. partitioning [ARV04, BCP01, Kel04, KRV06, OSVV08, ST04]. partitions [NdM06]. party [CLOS02, CGS02, GCKL08, HNRR04, Kil00, Lin03, Pas04, SV05a, vAHL05]. passing [ADS09]. path [AMS00, CX00, KK03]. paths [AZ05b, BHS02, Cha07, CKS06, DI03, FMS00, Gab04, Kos01, Pap07, Tho03a, Tho05, VWY07]. pattern [She08]. PCP [DGKR03, Din06, GR07, HK04, ST00]. PCPs [BSSVW03, BSGH^{+}04, BSS05, FS08, IKW09, ST06]. peer [AHKV03]. peer-to-peer [AHKV03]. Pell [Hal02]. average-case [Mic02]. deletes [ESSS01]. min-multicut [Oba04]. Schrijver [Gur06]. write [BJR07]. percolation [Gác01]. perfect [FS09, MV00, Pag01]. performance [AS04, FFV05, PS06]. permanent [CCL08, CRS02, JSV01, Raz04]. permanents [Für00]. permutation [Vöc01]. permutations [ITT03]. persistence [NT01]. perspective [Kle00]. perspectives [OSS00, Yao01]. perturbed [VT07]. phenomena [BLMN03]. phenomenon [Kle00]. phylogenetic [DMR06a]. phylogenies [MR05]. pigeonhole [MPW00, PR01b, Raz02b]. planar [AKR03, CG09, CKS06, CX00, Gon05, HVV07, Ist00, Kle06, KK03, MV00]. plane [CG09]. player [Pap05]. Playing [KKL07]. point [CG06, DS08, KN09, KRS04, Wee05]. pointer [BLM^{+}02]. points [CD05]. poly [BSS05, KP01]. poly-log [BSS05]. poly-logarithm [KP01]. polygons [DELM03]. polyhedral [IH01, iKM08]. Polylogarithmic [HK03, CKS04b]. polymorphisms [BKN08b]. Polynomial [BBS09, Bas05, CK07, Hal02, KSY00, SV05b, AJL06, Ajt04, ACF^{+}03, BL02, BRSW06, BV03, CD02, DV04, DS05b, GJ07, GS00c, HR07b, IFF00, JSV01, KI03, KS06, KLE07, LMW00, OS03, Raz04, RS06a, Shp01, SV08b, ST01, Uma08, Val01]. Polynomial-time [BBS09, Hal02, KSY00, BL02, CD02, DV04, GS00c, IFF00, JSV01, KS06]. polynomially [Rao06, SS04]. polynomials [BSK09, Bog05, Gur06, KS01b, Lov08]. polytope [MS09b]. polytopes [KN09]. portfolio [KNT00]. Portland [ACM00]. poset [FLN^{+}02]. positive [AK08]. possible [AN04a]. Potato [BHW00]. Potts [Sly09]. power [ACKM05, Kho02b, MNW04, Yao03]. power-law [ACKM05]. precision [OSS00]. preconditioners [DH07]. predecessor [PT06]. preemptive [CK02]. preferential [BCDR07]. preliminary [FPS00]. presence [BIW08, FI04, KS03]. preserve [Vad00]. preserving [NN01]. price [AAE05, BHLR08, CK05c, DMT07, MS01a, Rou02, Rou09]. priced [CFG^{+}00, CL05]. Pricing [CGP08, wCG06, CDR03, DKM06]. Primal [KR03, LRS04, AK07b, JV02]. Primal-dual [KR03, LRS04, AK07b]. primal-dual-type [JV02]. principal [Hal02]. principle [AR04, MPW00, PR01b, Raz02b]. principles [Ind07]. priorities [KMT03]. priority [AT00, ABD^{+}02, Sze06, Tho03a]. privacy [BLR08, DMT07, DL09, GRS09]. Private [BCNW06, FFKN09, HKKN01, DNR^{+}09, GR03, GKK^{+}01, KY00, NRS07]. private-key [KY00]. probabilistic [Adl02, AEM05, KY00, Kla00, Tar03, Vat01]. probe [BR00, JKKR03, Pag01, Vio09]. probing [PPR07]. problem [Aar02, AKS01, AAA03, AZ05a, AZ05b, AZ06, BS06, BKW00, BCP01, CCL08, CKS04a, CLL^{+}08, CW09a, EK02, Elk04, FL08, FFV05, Gar05, GSVV01, GKK^{+}01, Hal02, HR07b, JKKR03, KC00, KNT00, OS08, PV00, Pei09, RTV06, Tho03a, Gur03, Von08]. problems [ART06, ACG^{+}02, ACMM05, AAC07, AMS00, AdlVKK02, AGK^{+}01, AHW08, BR00, BKN08b, BV02a, BCNW06, BI04, BK08, CLX09, COP03, CKS05, CK06, CK07, CF08, CS00, DKSV06, GW01, GP06, GMM01, GM07, GT00, Ind04, JMS02, KS09a, KKK00, LRS04, Mol02b, MR01b, MPSV00, Pla03, RR06, Tok01, Tre01, Yao04, dlVKKR03, dlVKKV05]. procedures [HM07, KT00]. proceedings [ACM05, ACM07, ACM08, ACM09, ACM00, ACM01, ACM02, ACM03, ACM04, ACM06, ACM10]. process [AHKV03]. processor [CGKK04]. product [AŠdW06, BATS08, IJKW08, IKW09, JKN08, OS00, RS01, Raz02a]. production [CFJ^{+}09]. products [GGR09]. profile [KNT00]. Profit [Mey01]. Profit-earning [Mey01]. program [ACGH01, BW01, GGH^{+}08, RS08, RW00b]. program-size [RW00b]. programming [AE05, Dye03, GW01, KN09, KS06]. Programs [Gol09, AK07b, AK08, BV02b, DV04, MOS01]. projection [DF08, Rao08]. proof [ABM01, BBB^{+}00, Dan07, DS02b, Fri03, KW00, MPW00, Mos09]. proofs [DSY00, DM02, GKR08, Gur06, Lyn07, Nor06, SV08a, Vad00]. properties [BV04, BSHR03, FN01, FN05, GRK04, JV07, PR01a, Sha06, Sha09, Val08]. Property [RS07, AS05, BSS08, KS08a]. Proportional [WZ07]. propositional [ABSAW00]. protocol [Amb01, AJS07]. protocols [Ant06, KLP05, KKD01, Kla00, KLR06, LMas05, NN01, PR05]. Provably [LRS06]. prover [Kho02b, Vad00]. provers [NV06]. proving [KI03]. Provisioning [GKK^{+}01]. proximity [BSGH^{+}04, GR09]. Pseudo [NRR00, Uma02, ISW00]. Pseudo-random [NRR00, Uma02, ISW00]. Pseudorandom [Bog05, RTV06, Lov08]. PSPACE [Gut00]. PTAS [BCES06, DP09]. Public [Pei09]. Public-key [Pei09]. pure [FPT04, SV08c]. pushdown [AM04].

QoS [AR03, KLM^{+}01]. Quadratic [AMMN05, Bas05, CCL08, Gro01]. quantifier [KK09]. Quantitative [dAM01]. Quantum [Aar02, ATSVY00, AAKV01, Amb00, Amb04, GSVV01, Kla03, Liu09, SS04, Val01, Wat01, Aar04a, Aar04b, ATS03, AJL06, AALV09, Amb01, ABN^{+}01, AŠdW06, BYJK04, BOH05, BBB^{+}00, CCD^{+}03, CGM^{+}09, CGS02, FIM^{+}03, GKRdW06, GKK^{+}07, Gav08, Gro00b, Gur03, GW07, HRTS00, Hal02, Hal05, HMR^{+}06, KdW03, KW00, Kla00, KNTSZ01, MNRS07, MRS07, NS02, Rei04, RS08, SV05b, She08, Shi05, TS09, Wat06, Yao03, Yao04, Zha06, vDMMS00]. quasi [BCES06, RS06a]. quasi-polynomial [RS06a]. quasi-PTAS [BCES06]. Québec [ACM02]. queries [CX00, DS05b, Fel06, GS00a, GSS01, KK03, Tho03b]. Query [CFG^{+}00, BSS05, BKW00, CK08, CGM^{+}09, Efr09, GR07, IKW09, KdW03, ST00, SS04, Yek07]. querying [CL05]. queue [ABD^{+}02, AR03]. queues [AT00, Tho03a].

Ramanujan [BSS09, BATS08]. Ramsey [BKS^{+}05, BRSW06, BLMN03]. Ramsey-type [BLMN03]. Random [AdlVKK02, BV03, DF08, FKK^{+}00, KN09, KL02, KK09, Ach00, AP03, AN04a, ART06, ACL00, Ale05, BV02b, BKL07, DW03, Fei04, FVV08, Gác02, GRK04, GKP02, HV02, HKLR05, ISW00, KV03, Mol02b, NRR00, Raz05, Reg05, SV05a, Sel09, TV05, Uma02, Vu05]. Randomized [BBN08, CV00, CMP08, Alb02, AC07, BJR07, DNS06, FS09, Imp06, KS06, MNW04, Sei00, Zha06]. Randomly [AH09, Hay03, HVV07, VT07]. Randomness [BSSVW03, CRVW02, KS01b, DI06, DSY08, GR03, SU07]. Randomness-efficient [BSSVW03]. range [ABR01, MPP05, Pat07]. Rank [Dan07, AM01, KMS07, NDT09]. ranking [ACN05, AGY09]. Rapid [Gro00b, MS01b]. rate [BSS05, SA02]. rate-based [SA02]. ratio [FS01]. Rational [HT04, OS00]. re [AGY09]. re-ranking [AGY09]. reachability [RZ04]. Read [SV08b, BJR07, BSW05, BW01]. Read-once [SV08b, BW01]. read/write [BJR07]. reading [CKS04b]. real [BD00]. real-time [BD00]. recognize [BC05]. Recognizing [SSŠ02a]. recommendation [DKR02]. Reconstructing [CS03]. Reconstruction [Sly09, DMR06a, HRTS00]. rectangles [AES09]. rectangular [KMT03]. Reducibility [GP06]. Reducing [AAM03]. reduction [BKL07, CW09a, DV07]. reductions [CHKX04, Tul09]. redundancy [FI04]. Reed [GKZ08, GR05]. reference [AFG02]. regression [DK08]. Regret [BHLR08, EdMN09, KPT09]. Regular [PR01b, ACKM05, AJPU02, BC05, Hol02, KV03, RTV06, dAM01]. regularity [Sha06]. Reimer [Smy02]. related [BR00, GK06, KS09a, Kon02, SS04]. Relations [Fei02]. relaxations [CMM09, FS01, MS09b, STT07]. release [DNR^{+}09]. removal [DV01, RS07]. rent [FKLS06]. rent-or-buy [FKLS06]. Reordering [ERW07]. repetition [Hol07, PV07, Rao08]. replace [Gav08]. replication [BIW08]. reporting [ABR01, MPP05]. representations [GGI^{+}02a, HRTS00]. Representing [Ajt05]. require [SV08a]. requires [CKPR01]. rescaling [DV04]. Resettable [CGGM00, KP01]. residual [KL02]. resiliency [Gaf09]. resilient [CM09]. resistance [AH09]. resistances [SS08]. Resolution [Raz02b, AJPU02, Ale05, BS02b, Nor06, NH08, PR01b]. resource [BNBYF^{+}00, CGKK04, wCG06, KKD01]. response [WZ07]. restricted [GLMM04, ITT03, KR02]. results [AAT05, CN04, Dev04, DNR^{+}09, EGM02, Kho02a, Lyn07, Pap07, Rag08, Sri00, Tre01]. Rethinking [Rex08]. reusable [DG03]. revision [GS00a]. revisited [Gol02]. rings [DW09a]. rising [BP03]. risk [KNT00]. RL [RTV06]. RMR [AHW08]. Robust [BSGH^{+}04, CCM08, DKM06, DL09]. robustness [Rou09]. role [KS08a]. rotate [Sch05]. rotations [JvS05]. round [DS02b, GIKR01, Kho02b, SV05a]. rounding [Cha02b, FS09]. rounds [BOPV06, CKPR01, KP01]. routing [AJS07, AK04, AAE05, ACF^{+}03, BBCM04, BHW00, CKS05, CGKT07, DW09a, GMP00, HKLR05, LSZ08, MS01a, Rex08, Vöc01]. rows [CD02]. run [LV04]. Running [ACGH01]. salesman [PV00]. sample [ITT03]. Sampling [BYKS01, BY03, DV07, ACKM05, AdlVKK02, DI06, FRV06, GGI^{+}02a, Gro00b, GPRS04, KL02, LRS06, NRS07, Sze06]. Sampling-based [DV07, LRS06]. San [ACM03, ACM07]. Santa [BS06]. SAT [Ach00, AP03, Aus07]. satisfaction [ART06, BK08, Mol02b, dlVKKV05]. Satisfiability [Gut00, CS00]. satisfiable [OW09]. Saving [Gar05]. scaling [BCP01]. schedulers [CLTW02]. scheduling [Alb02, AJS07, AE05, Bac02, BP03, BNBYF^{+}00, BL01, BD00, CGKK04, CN04, CS00, Gác02, GLMM04, GRS03, Kon02, Rou01, SSS02b, SA02]. scheme [KSY00, RS06a]. schemes [CK02, DG03, FKLS06, GGK03, KS09a, OSS00, dlVKKR03, dlVKKV05]. Schnorr [Ajt03]. Schrijver [AAT05, Dan07, STT07]. science [CL07]. SDP [MNRS08, OW08]. Search [MNRS07, Aar04a, AGK^{+}01, BR00, BCNW06, BLM^{+}02, KR03, PT06, Rag06, SS04, Zha06]. Searching [CG06, AT00, CL01, FI04, Tok01]. Seattle [ACM06]. Second [ACM00, Fri03]. secret [CDD00, GIKR01, HT04]. secrets [DW09b]. Secure [CGS02, CLOS02, GIKR01, GCKL08, HNRR04, IKLP06, IKOS07, KLP05, Kat07, Kil00, KLR06, Lin03, NN01, Pas04]. security [BBB^{+}00, DM02, KY00, KLR06, LPV09, PS04]. seed [ISW00, TS09]. seeds [Raz05]. segments [CG09]. selection [DK08, FM07, SV05a]. Self [vDMMS00, ACGH01, ACG^{+}02, RW00b]. self-assembled [ACGH01, RW00b]. self-assembly [ACG^{+}02]. Self-testing [vDMMS00]. Selfish [CKV02, ADTW03, CDR03, MS01a]. semantic [FST05, JS08]. semi [BPR05, Bas05]. semi-algebraic [BPR05, Bas05]. semialgebraic [BC04]. semidefinite [AK07b, FS01, GW01]. sensing [GSTV07]. sensitivity [NRS07]. separated [GZ03]. Separating [She07, KPT05, Nor06]. separation [AJPU02, BYJK04, NH08]. separations [GKRdW06, GKK^{+}07]. separators [FHL05, FM06]. sequence [DELM03, MS00]. sequences [Gác01]. series [BKL07, Kos01]. series-parallel [BKL07]. Server [BP03, CMP08, CKV02, JKK^{+}01]. set [AAA03, CLL^{+}08, GGP05, JLN^{+}05, KNTSZ01]. sets [AP09, BHPI02, BPR05, Bas05, BSSVW03, BY07, BC04, CG06, KRS04, MP03, RW03, Wei06, vDMMS00]. Setting [Ach00]. setup [Lin03, PS04]. shapes [DGJ03]. shared [CK05b]. Sharing [FPS00, CDD00, FKLS06, GIKR01, HT04, RS06b]. Sharp [BCP01, GRK04, ABM01]. sharper [Gur06]. Sherali [Dan07, CMM09, MS09b]. Shor [Amb04]. Short [KK03, TS09, BSSVW03, Bon00, GN08]. shorter [BSGH^{+}04]. Shortest [CX00, AKS01, Ajt03, AMS00, BHS02, Cha07, DI03, HR07b, Pei09, Tho03a, Tho05]. should [Wig04]. shuffle [Mor05]. sieve [AKS01, MRS07]. signaling [Hol07]. signature [GGK03]. Similarity [Cha02b, BBV08]. Simple [BGK^{+}07, BSS05, FKLS06, DW03, DV04]. Simpler [GKR03, Gur06]. simplex [KS06, ST01]. simplifications [Hol07]. simplified [IJKW08]. simulated [Val01]. Simulating [BKS^{+}05]. simulation [BL02, Gaf09, KW00]. simulations [CGM^{+}09]. single [BL01, GMM01, KRV06, OSVV08, Tho03a]. singularity [TV05]. sink [GMM01]. Size [BS02b, ACGH01, AES09, FKN00, HR00, ITT03, Kon02, MR06, Raz04, RW00b]. skepticism [Aar04b]. sketch [GSTV07]. Sketching [MNS08]. slicing [MOS01]. slivers [ELM^{+}00]. slow [Gác01]. Small [AES09, DW09a, FM06, GGI^{+}02b, GKZ08, KRVZ06, Kle00, MV00, RS09, Rao06, Uma08]. Small-size [AES09]. small-space [GGI^{+}02b, KRVZ06]. small-world [Kle00]. smaller [GI02]. smallest [CLL^{+}02, Tre09]. Smooth [NRS07, Bon00]. Smoothed [ST01]. Smoothing [ELM^{+}00]. snapshot [Jay05]. snapshots [FFR03, FFR06]. Sobolev [BT03]. social [KT08, MR07]. solid [MS09a]. solid-on-solid [MS09a]. Solomon [GR05]. solution [ART06, dAM01]. solution-space [ART06]. solvability [MRR01]. solvable [Wat01]. Solving [BV02b, BI04, DV04, Pla06, ST04]. Some [BSHR03, O’D08, Pap07, Yao01, FFV05, Nan08, RW03]. Sorting [FI04, Han02, Kla03, Yao04]. source [BRSW06, Tho03a]. sources [KRVZ06, Rao06]. Space [ABSAW00, AMM02, SS02, Tho03b, ART06, AŠdW06, AKS03, BV02a, BS02b, CS03, FFR03, FFR06, GGI^{+}02b, Han02, KRVZ06, Kla03, Nor06, NH08, OP03, PT06, Rei05, SW03, Tri05]. Space-efficient [AMM02]. space-optimal [FFR03]. spaces [ABN07, BCR01, CG06, ERW07, KSU08]. spacious [Nor06]. Span [RS08]. Span-program-based [RS08]. spanner [EP01, Kle06]. spanners [CLPR09]. spanning [CS04, Elk04, EEST05, FL08, KR00, SL07]. Sparse [Shp01, AP09, BSS08, CH02, FMS00, GGI^{+}02a]. sparsest [ALN05, DKSV06]. sparsification [ST04, SS08]. sparsifiers [BSS09]. Spatial [KKD01]. Spectral [AFK^{+}01, Kel04, Vu05, KM04]. speed [CGKM09]. speedup [CCD^{+}03]. spending [Dev04]. spread [KL03a, Sch05]. squares [ACGH01, CJK^{+}00, RW00b]. st [Tri05, Rei05]. st-connectivity [Tri05, Rei05]. stabbing [Tho03b]. Stability [AJS07, AKK02]. Stackelberg [Rou01]. stand [LPV09]. stand-alone [LPV09]. state [ATS03, GKRdW06]. Stateless [AK08]. states [ASV00, HMR^{+}06]. static [ABR01, GGP05]. Statistical [Ist00, ATS03, BKW00]. Statistically [HR07a]. Statistically-hiding [HR07a]. statistics [DL09]. Steiner [CH03, FKLS06, GK09, JLN^{+}05, MR01b]. STOC [ACM05, ACM07, ACM08, ACM09]. stochastic [AHKV03, FKLS06, GPRS04, GK09, GKR09, LRS06]. stock [KNT00]. Stockmeyer [For05]. storage [BKK^{+}03, DM02, TS09]. stories [KKR01]. strategies [CFG^{+}00, Rou01]. strategy [CL05]. stream [AJKS02, BJR07, CCM08, SS02]. streaming [COP03, CW09b]. streams [FS05, GKS01, Ind04, IW05]. stretch [EEST05]. Strict [BL02]. Strictly [RW00a]. string [GV00, SS01, SSŠ02a]. strip [JvS05]. strong [AAT05, Sri00]. stronger [HLS07]. strongly [IFF00]. structural [JV07]. structures [NT01, Vio09]. Sub [MR06, BRSW06, VWY07]. Sub-constant [MR06]. sub-cubic [VWY07]. sub-polynomial [BRSW06]. subadditive [Fei06]. subdistribution [JKN08]. subexponential [Efr09, Yek07]. subgraph [Cha02a, FL08]. subgraphs [AS03, CVV02, KN04, VW09]. subgroup [GSVV01, HRTS00]. Sublinear [BKR04, CLM03, BEK^{+}03, CS04]. sublinear-time [CS04]. submodular [FI00, IFF00, LMNS09, Von08]. submodularity [MR07]. subsequences [Bac02]. subset [BHKK07, DK08, Kle06]. subspace [BSK09, DV07]. succinct [FS08, Vio09]. suffix [CH00, FM07, GV00]. sum [BCR01, CP01, CJK^{+}00]. sum-of-squares [CJK^{+}00]. sums [Fei04]. super [Raz04]. super-polynomial [Raz04]. superpolylogarithmic [Gab04]. supported [AH09]. surface [BCSV04]. surfaces [Ist00]. Survivable [LNSS07, GKR09, LS08]. switches [AR03, KLM^{+}01]. switching [AR04]. symmetric [DW09b, For00, MW02, Val08].

Symposium [ACM00, ACM01, ACM02, ACM03, ACM04, ACM05, ACM06, ACM07, ACM08, ACM09, ACM10]. systems [AKK02, Dan07, DKR02, FM00a, JKK^{+}01, KLE07, KW00, MRR01, S̊O01, ST04].

tableaux [APU01]. tables [CD02, CDR05]. tails [DFKO06]. take [MT09]. takes [ST01]. talks [Wig04]. tall [Ree00]. Tardos [Smy02]. task [FM00a, JKK^{+}01]. tasks [Kon02]. tatonnement [CF08]. TCP [GK03, KKR01]. techniques [BKV05, Cha02b]. telephone [EK02]. telling [AAM03]. temporal [BK08, KKK00]. Tensor [HR07b, Shi05, dlVKKV05, GGR09]. Tensor-based [HR07b]. Terminal [AK07a]. terminals [CKS05]. test [BSHR03, MR06]. testable [AS05, BSS08, Mei08, Sha06]. testers [IKW09]. Testing [AS03, AAK^{+}07, Bla09, FN01, FN05, PR01a, RS05, Val08, AC04, BKR04, BCL^{+}06, DS05b, FLN^{+}02, Fis04, FIS05, GR09, Gro00a, KS08a, KS01b, RS07, Sha09, SW04, SV08b, vDMMS00]. tests [BSSVW03, KI03, OW08, Sam07]. text [GV00]. their [IKOS04, PW08]. theorem [CG04, Din06, Nan08, PV07, BCI^{+}08]. theorems [AŠdW06, CKL^{+}04, IJKW08, JKN08, Kil00, Oba04]. theoretically [KLR06]. Theory [ACM00, ACM01, ACM02, ACM03, ACM04, ACM05, ACM06, ACM07, ACM08, ACM09, ACM10, AW08, ABN06, BBS09, BY03, BLR08, GS00a, GW07, Lyn07, Siv02, Pap02]. Thirty [ACM00, ACM03, ACM06]. Thirty-Eighth [ACM06]. Thirty-Fifth [ACM03]. Thiry [ACM02]. Thiry-Fourth [ACM02]. Thorp [Mor05]. though [Gro00b]. three [APST03, CS03, Ist00]. three-dimensionality [Ist00]. threshold [ABM01, AP03, BCP01, OS03, RS09, Wei06]. thresholds [GRK04, Mol02b]. throughout [BD00]. throughput [GMP00]. thy [MNW04]. tide [BP03]. Tight [AT00, AC07, AHW08, DW09a, DM02, STT07, CKL^{+}04, FRT03, FFR03, IH01]. Tighter [BR00]. tilings [DLS01]. Time [BV02a, FFR06, PT06, SW03, Ach00, ACGH01, Ajt04, AŠdW06, AO09, ACF^{+}03, BBS09, BBCM04, BL02, BE09, Bas05, BL01, BV03, BD00, CGKM09, CP07, CKZ01, CK02, CGKK04, CGM^{+}09, CF08, CD02, CS04, DV04, Elk04, FFR03, GK06, Gro01, GI02, GI03, GS00c, Hal02, Han02, IFF00, JSV01, KS09a, iKR07, iKM08, KS06, KSY00, KW00, Kla03, KS01a, KK03, LMW00, MS09a, Mor05, NdM06, OP03, RS06a, RZ04, SSS02b, SV05b, S̊O01, ST01, ST04, Tho03a, Tra01, Val01, VW06, VWY07, YYI09]. time-approximation [Elk04]. Time-space [BV02a, FFR06, PT06, SW03, AŠdW06, Kla03]. time-windows [BBCM04]. times [Tho05]. timing [Gol02, KLP05]. tolerant [BKW00, CLPR09, vDMMS00]. top [Bas05]. topics [O’D08]. topology [Rou02]. total [BHLR08]. Touring [DELM03]. tournaments [CGP08]. traceback [Adl02]. traceroute [ACKM05]. tractable [GSS01]. trade [PT06, RS06b]. trade-offs [PT06, RS06b]. tradeoff [SW03]. Tradeoffs [Adl02, AŠdW06, BV02a, BS02b, DSY08, Elk04, FFR06, Kla03, SU07]. trading [DKM06]. traffic [CKV02]. transform [AC06a, Liu09]. transformation [Vad00]. transitive [BHS02]. translation [FIM^{+}03]. transmissions [FPS00]. transversals [EGM02]. trapdoor [PW08]. Trapdoors [GPV08]. traveling [PV00]. Tree [BC05, CH00, Elk04, FRT03, FKLS06, HKPB07, JLN^{+}05, MR01b, NdM06, Ree00, Tho08, Wei06]. Tree-walking [BC05]. tree-width [NdM06]. trees [BLM^{+}02, CMP08, CS04, DF08, EEST05, GKK08, GV00, KR00, SL07]. triangle [VW06]. triangular [AKS03]. triangulation [RS06a]. triangulations [Cla06]. trisector [AMT06]. true [GOS08, WW01]. truly [VWY07]. trusted [PS04]. truth [AAM03]. truth-telling [AAM03]. Truthful [DNS06]. TSP [BBCM04, JLN^{+}05, Kle06]. Tutte [GJ07]. Twice [BSS09]. Twice-Ramanujan [BSS09]. Two [JKS03, AN04a, CLOS02, Fel06, Gon05, GCKL08, HNRR04, KS08b, Kil00, Lin03, SV05a, Shp07, vAHL05]. two-level [Fel06]. two-party [CLOS02, GCKL08, HNRR04, Kil00, Lin03, SV05a, vAHL05]. type [BLMN03, JV02]. Typical [BV04].

UGC [MNRS08]. unbalanced [TSUZ01]. unbounded [Fei04]. uncapacitated [FFV05]. Uncertainty [Ind07, FMP^{+}00]. Unconditional [Elk04, Lov08]. uncoupled [HM07]. UnCut [ACMM05]. Undirected [Rei05, AZ05a, AZ05b, Tri05]. unequal [GKY02]. unfair [FM00a]. unified [BNBYF^{+}00, CLTW02, LPV09]. Uniform [IJKW08, OP03, CK05a, DW09a, Sel09, SU07, Tre05]. uniformity [ST06]. uniformly [GKP02]. unimodal [BKR04]. union [KST02]. union-find [KST02]. Unique [AKK^{+}08, CMM06, GI02, Kho02b]. uniqueness [Dev04]. unit [GZ03, Hal05, Kon02, SV05b]. unit-disk [GZ03]. unit-size [Kon02]. Universal [JLN^{+}05, JS08, BW01, LPV09, PS04, vDMMS00]. Universality [Ist00]. Universally [CLOS02, GRS09]. unlikely [BSW05]. unrelated [AE05]. unrelatedmachines [CGKM09]. unsplittable [AAE05, BCES06]. unweighted [HKPB07]. update [CK05b, RZ04, Tho05]. updates [KPT09]. upper [Zha06]. USA [ACM03, ACM04, ACM05, ACM06, ACM07, ACM09, ACM10]. users [CDR03]. Using [HVV04, KS04, AL00, AP09, BATS08, BW01, Bon00, Cla06, DKSS08, Gen09, HRTS00, HR02, KRV06, Liu09]. usually [ST01]. utilitarian [BKV05]. utility [Fei06, GRS09]. utility-maximizing [GRS09].

Valiant [Gur06, Wig09]. value

[Sri00, Von08]. values [AN04a, AACW06]. variables [Ach00, Fei04, ST06]. variance [Fei04]. VCG [DN07]. VCG-based [DN07]. vector [AKS01, Ajt03, HR07b, OS00, Pei09]. vectors [GN08, MRR01]. Vegas [Kla00]. vehicle [BBCM04]. verifiable [CDD00, GIKR01]. verifier [HK04]. Verifying [CH02, GGH^{+}07]. version [FPS00]. versus [BKK^{+}03, FN05]. Vertex [BKL07, Hol02, AKR03, CCK08, CLL^{+}08, CVV02, DGKR03, FL08, FHL05, STT07]. vertex-cuts [AKR03]. vertices [BK09]. via [AN04b, BHPI02, BBV08, BY03, BCG09, BSSVW03, CMV05, DS08, DH07, GGI^{+}02a, GW01, JV02, JKN08, JKK^{+}01, KdW03, KL03a, KN04, MNRS07, OSVV08, Siv02, Tho08]. Victoria [ACM08]. video [CLTW02]. view [HNRR04]. virtual [GKK^{+}01]. Visibly [AM04]. volume [GS00c]. Voroni [IH01]. Voronoi [AMM02, CP07]. VPN [GOS08]. vs [RTV06, SU07].

WA [ACM06]. Waerden [Gur06]. Waerden/Schrijver [Gur06]. Wait [ASS02]. Wait-free [ASS02]. walk [CCD^{+}03, MNRS07]. walking [BC05]. walks [AAKV01, ABN^{+}01, BV02b, BKL07, FKK^{+}00, Gác02, KN09, RTV06]. Wardrop [FRV06]. way [AGGM06, AGGM10, BYJK04, GKK^{+}07, HR07a, Kla00, Tho08]. WDM [RW00a]. weak [DW09b, Hol05, MPW00, PR01b, Raz02b, Raz05]. weakly [BEK^{+}03]. Web [CF02, Rag06]. weight [AJS07, CS04, FHL05, RS06a, VW06]. weighted [CGKM09, Cha07, CKZ01, CK02, CG04, MR01b, TW07, VW09]. weights [HLS07]. Welfare [Von08, Fei06]. Well [GZ03, CKS05]. well-linked [CKS05]. Well-separated [GZ03]. width [FRRS06, NdM06, Nor06]. wild [MT09]. wildcard [CH02]. Wilson [BRSW06]. window [BCP01]. windows [BBCM04]. Winkler [Gác01]. winners [BV04]. wires [KPT05]. wise [AAK^{+}07, ITT03]. within [GN08, HR07b, Hol02, LMW00, O’D02, SL07]. without [DS05a, FI04, Lin03, PS04]. word [Pla06]. Work [GRS03, For05, Wig09]. Work-competitive [GRS03]. world [DW09a, Kat07, Kle00]. Worst [Tho05, Ajt03, AT00, CKS04b, Mic02, PR07, Pei09]. Worst-case [Tho05, Ajt03, AT00, CKS04b, Mic02, PR07, Pei09]. worst-case/ average-case [Mic02]. writer [FFR03, Jay05].

yields [Ach00].

zag [BATS08]. Zero [IKOS07, NV06, Wat06, ATS03, AR04, CGGM00, CKPR01, DSY00, DS02b, Gol02, KP01, MP06]. Zero-knowledge [IKOS07, Wat06, CGGM00, CKPR01, DSY00, Gol02, KP01]. zero-one [AR04]. zig [BATS08]. zig-zag [BATS08].

References

Alon:2003:OSC

[AAA03] Noga Alon, Baruch Awerbuch, and Yossi Azar. The online set cover problem. In ACM [ACM03], pages 100–105. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Agarwal:2007:IAD

[AAC07] Amit Agarwal, Noga Alon, and Moses S. Charikar. Improved approximation for directed cut problems. In ACM [ACM07], pages 671–680. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Angluin:2006:LCI

[AACW06] Dana Angluin, James Aspnes, Jiang Chen, and Yinghua Wu. Learning a circuit by injecting values. In ACM [ACM06], pages 584–593. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Awerbuch:2005:LPR

[AAE05] Baruch Awerbuch, Yossi Azar, and Amir Epstein. Large the price of routing unsplittable flow. In ACM [ACM05], pages 57–66. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Alon:2007:TWA

[AAK^{+}07] Noga Alon, Alexandr Andoni, Tali Kaufman, Kevin Matulef, Ronitt Rubinfeld, and Ning Xie. Testing k-wise and almost k-wise independence. In ACM [ACM07], pages 496–505. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Aharonov:2001:QWG

[AAKV01] Dorit Aharonov, Andris Ambainis, Julia Kempe, and Umesh Vazirani. Quantum walks on graphs. In ACM [ACM01], pages 50–59. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Aharonov:2009:DLQ

[AALV09] Dorit Aharonov, Itai Arad, Zeph Landau, and Umesh Vazirani. The detectability lemma and quantum gap amplification. In ACM [ACM09], pages 417–426. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Awerbuch:2003:RTT

[AAM03] Baruch Awerbuch, Yossi Azar, and Adam Meyerson. Reducing truth-telling online mechanisms to online optimization. In ACM [ACM03], pages 503–510. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Aaronson:2002:QLB

[Aar02]  Scott Aaronson.  Quantum

lower bound for the collision problem. In ACM [ACM02], pages 635–642. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Aaronson:2004:LBL

[Aar04a] Scott Aaronson. Lower bounds for local search by quantum arguments. In ACM [ACM04], pages 465–474. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Aaronson:2004:MFS

[Aar04b]  Scott Aaronson.  Multilin-

ear formulas and skepticism of quantum computing. In ACM [ACM04], pages 118–127. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Aaronson:2005:CA

[Aar05] Scott Aaronson. The complexity of agreement. In ACM [ACM05], pages 634–643. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Alekhnovich:2005:TSN

[AAT05] Mikhail Alekhnovich, Sanjeev Arora, and Iannis Tourlakis. Towards strong nonapproximability results in the Lovasz– Schrijver hierarchy. In ACM [ACM05], pages 294–303. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Arge:2002:COP

[ABD^{+}02] Lars Arge, Michael A. Bender, Erik D. Demaine, Bryan Holland-Minkley, and J. Ian Munro. Cache-oblivious priority queue and graph algorithm applications. In ACM [ACM02], pages 268–276. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Achlioptas:2001:STP

[ABM01] Dimitris Achlioptas, Paul Beame, and Michael Molloy. A sharp threshold in proof complexity. In ACM [ACM01], pages 337–346. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Ambainis:2001:ODQ

[ABN^{+}01] Andris Ambainis, Eric Bach, Ashwin Nayak, Ashvin Vishwanath, and John Watrous.

One-dimensional quantum walks. In ACM [ACM01], pages 37– 49. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Abraham:2006:AME

[ABN06] Ittai Abraham, Yair Bartal, and Ofer Neimany. Advances in metric embedding theory. In ACM [ACM06], pages 271–286. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Abraham:2007:LEM

[ABN07] Ittai Abraham, Yair Bartal, and Ofer Neiman. Local embeddings of metric spaces. In ACM [ACM07], pages 631–640. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Alstrup:2001:OSR

[ABR01] Stephen Alstrup, Gerth Brodal, and Theis Rauhe. Optimal static range reporting in one dimension. In ACM [ACM01], pages 476–482. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Alekhnovich:2000:SCP

[ABSAW00] Michael Alekhnovich, Eli Ben-Sasson, Alexander A., and Avi Wigderson. Space complexity in propositional calculus. In ACM [ACM00], pages 358– 367. ISBN 1-58113-184-4. URL

http://www.acm.org/pubs/
articles/proceedings/stoc/

                                        335305/p358-alekhnovich/
                                        p358-alekhnovich.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        proceedings/stoc/335305/
                                        p358-alekhnovich/. ACM or-

der number 508000.

Ailon:2004:LBL

[AC04] Nir Ailon and Bernard Chazelle. Lower bounds for linear degeneracy testing. In ACM [ACM04], pages 554–560. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Ailon:2006:ANN

[AC06a] Nir Ailon and Bernard Chazelle. Approximate nearest neighbors and the fast Johnson– Lindenstrauss transform. In ACM [ACM06], pages 557–563. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

                                        //portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Arora:2006:NAG

[AC06b] Sanjeev Arora and Eden Chlamtac. New approximation guarantee for chromatic number. In ACM [ACM06], pages 215–224. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

                                        /portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Attiya:2007:TBA

[AC07] Hagit Attiya and Keren Censor. Tight bounds for asynchronous randomized consensus. In ACM [ACM07], pages 155–164. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Azar:2003:OOR

[ACF^{+}03] Yossi Azar, Edith Cohen, Amos Fiat, Haim Kaplan, and Harald Racke. Optimal oblivious routing in polynomial time. In ACM [ACM03], pages 383–388. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Adleman:2002:COP

[ACG^{+}02] Len Adleman, Qi Cheng, Ashish Goel, Ming-Deh Huang, David Kempe, Pablo Moisset de Espanés, and Paul Wilhelm Karl Rothemund. Combinatorial optimization problems in self-assembly. In ACM [ACM02], pages 23–32. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Adleman:2001:RTP

[ACGH01] Leonard Adleman, Qi Cheng, Ashish Goel, and Ming-Deh Huang. Running time and program size for self-assembled squares. In ACM [ACM01], pages 740–748. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Achlioptas:2000:SVT

[Ach00] Dimitris Achlioptas. Setting 2 variables at a time yields a new lower bound for random 3-SAT (extended abstract). In ACM [ACM00], pages 28–37. ISBN 1-58113-184-4. URL

http://www.acm.org/pubs/
articles/proceedings/stoc/
335305/p28-achlioptas/p28-
achlioptas.pdf;   http://
www.acm.org/pubs/citations/
proceedings/stoc/335305/
p28-achlioptas/. ACM order

number 508000.

Achlioptas:2005:BTS

[ACKM05] Dimitris Achlioptas, Aaron Clauset, David Kempe, and Cristopher Moore. On the bias of traceroute sampling: or, power-law degree distributions in regular graphs. In ACM [ACM05], pages 694–703. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Aiello:2000:RGM

[ACL00] William Aiello, Fan Chung, and Linyuan Lu. A random graph model for massive graphs. In ACM [ACM00], pages 171–180. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p171-aiello/p171-aiello.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p171-aiello/.

ACM order number 508000.

ACM:2000:PTS

[ACM00] ACM, editor. Proceedings of the Thirty Second Annual ACM Symposium on Theory of Computing: Portland, Oregon, May 21–23, [2000]. ACM Press, New York, NY, USA, 2000. ISBN 1-58113-184-4. ACM order number 508000.

ACM:2001:PAA

[ACM01] ACM, editor. Proceedings of the 33rd Annual ACM Symposium on Theory of Computing: Hersonissos, Crete, Greece, July 6–8, 2001. ACM Press, New York, NY, USA, 2001. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

ACM:2002:PTF

[ACM02]  ACM, editor.  Proceedings

of the Thiry-Fourth Annual ACM Symposium on Theory of Computing, Montréal, Québec, Canada, May 19–21, 2002. ACM Press, New York, NY, USA, 2002. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

ACM:2003:PTF

[ACM03] ACM, editor. Proceedings of the Thirty-Fifth ACM Symposium on Theory of Computing, San Diego, CA, USA, June 9–11, 2003. ACM Press, New York, NY, USA, 2003. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

ACM:2004:PAA

[ACM04] ACM, editor. Proceedings of the 36th Annual ACM Symposium on the Theory of Computing: Chicago, Illinois, USA, June 13–15, 2004. ACM Press, New York, NY, USA, 2004. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

ACM:2005:SPA

[ACM05] ACM, editor. STOC ’05: proceedings of the 37th Annual ACM Symposium on Theory of Computing: Baltimore, Maryland, USA, May 22–24, 2005. ACM Press, New York, NY, USA, 2005. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

ACM:2006:PTE

[ACM06] ACM, editor. Proceedings of the Thirty-Eighth Annual ACM Symposium on Theory of Computing 2006, Seattle, WA, USA, May 21–23, 2006. ACM Press, New York, NY, USA, 2006. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

ACM:2007:SPA

[ACM07] ACM, editor. STOC ’07: proceedings of the 39th Annual ACM Symposium on Theory of Computing, San Diego, California, USA, June 11–13, 2007. ACM Press, New York, NY, USA, 2007. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

ACM:2008:SPA

[ACM08] ACM, editor. STOC ’08: proceedings of the 40th Annual ACM Symposium on Theory of Computing, Victoria, British Columbia, Canada, May 17–20, 2008. ACM Press, New York, NY, USA, 2008. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

ACM:2009:SPA

[ACM09] ACM, editor. STOC ’09: proceedings of the 2009 ACM International Symposium on Theory of Computing, Bethesda, Maryland, USA, May 31–June 2, 2009. ACM Press, New York, NY, USA, 2009. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

ACM:2010:PAI

[ACM10] ACM, editor. Proceedings of the 2010 ACM International Symposium on Theory of Computing: June 5–8, 2010, Cambridge, MA, USA. ACM Press, New York, NY, USA, 2010. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL

http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Agarwal:2005:AAM

[ACMM05] Amit Agarwal, Moses Charikar, Konstantin Makarychev, and Yury Makarychev. O(\sqrt{log n}) approximation algorithms for min UnCut, min 2CNF deletion, and directed cut problems. In ACM [ACM05], pages 573–581. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Ailon:2005:AII

[ACN05] Nir Ailon, Moses Charikar, and Alantha Newman. Aggregating inconsistent information: ranking and clustering. In ACM [ACM05], pages 684–693. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Adler:2002:TPP

[Adl02] Micah Adler. Tradeoffs in probabilistic packet marking for IP traceback. In ACM [ACM02], pages 407–418. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Alon:2002:RSA

[AdlVKK02] Noga Alon, W. Fernandez de la Vega, Ravi Kannan, and Marek Karpinski. Random sampling and approximation of MAX-CSP problems. In ACM [ACM02], pages 232–239. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Arora:2009:MPA

[ADS09] Sanjeev Arora, Constantinos Daskalakis, and David Steurer. Message passing algorithms and improved LP decoding. In ACM [ACM09], pages 3–12. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Anshelevich:2003:NON

[ADTW03] Elliot Anshelevich, Anirban Dasgupta, Eva Tardos, and Tom Wexler. Near-optimal network design with selfish agents. In ACM [ACM03], pages 511–520. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Azar:2005:CPS

[AE05] Yossi Azar and Amir Epstein. Convex programming for scheduling unrelated parallel machines. In ACM [ACM05], pages 331–337. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Adler:2005:TAO

[AEM05] Micah Adler, Jeff Edmonds, and Jivri Matousek. Towards asymptotic optimality in probabilistic packet marking. In ACM [ACM05], pages 450–459. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Aronov:2009:SSN

[AES09] Boris Aronov, Esther Ezra, and Micha Shair. Small-size ϵ-nets for axis-parallel rectangles and boxes. In ACM [ACM09], pages 639–648. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Albers:2002:PLR

[AFG02] Susanne Albers, Lene M.

Favrholdt, and Oliver Giel. On paging with locality of reference. In ACM [ACM02], pages 258–267. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Aggarwal:2005:DA

[AFG^{+}05] Gagan Aggarwal, Amos Fiat, Andrew V. Goldberg, Jason D. Hartline, Nicole Immorlica, and Madhu Sudan. Derandomization of auctions. In ACM [ACM05], pages 619–625. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Azar:2001:SAD

[AFK^{+}01] Yossi Azar, Amos Fiat, Anna Karlin, Frank McSherry, and Jared Saia. Spectral analysis of data. In ACM [ACM01], pages 619–626. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Akavia:2006:BOW

[AGGM06] Adi Akavia, Oded Goldreich, Shafi Goldwasser, and Dana Moshkovitz. On basing one-way functions on NP-hardness. In ACM [ACM06], pages 701–710. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. See erratum

[AGGM10].

Akavia:2010:EBO

[AGGM10] Adi Akavia, Oded Goldreich, Shafi Goldwasser, and Dana Moshkovitz. Erratum for: On basing one-way functions on NP-hardness. In ACM [ACM10], pages 795–796. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL

http://www.gbv.de/dms/tib-
ub-hannover/63314455x.. See

[AGGM06].

Arya:2001:LSH

[AGK^{+}01] Vijay Arya, Naveen Garg, Rohit Khandekar, Kamesh Munagala, and Vinayaka Pandit. Local search heuristic for kmedian and facility location problems. In ACM [ACM01], pages 21–29. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Azar:2009:MIR

[AGY09] Yossi Azar, Iftah Gamzu, and Xiaoxin Yin. Multiple intents re-ranking. In ACM [ACM09], pages 669–678. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Austrin:2009:RSI

[AH09] Per Austrin and Johan Håstad.

Randomly supported independence and resistance. In ACM [ACM09], pages 483–492. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Adler:2003:SPH

[AHKV03] Micah Adler, Eran Halperin, Richard M. Karp, and Vijay V. Vazirani. A stochastic process on the hypercube with applications to peer-to-peer networks. In ACM [ACM03], pages 575–584. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Agol:2002:MKG

[AHT02] Ian Agol, Joel Hass, and William Thurston. 3-manifold knot genus is NP-complete. In ACM [ACM02], pages 761–766. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Attiya:2008:TRL

[AHW08] Hagit Attiya, Danny Hendler, and Philipp Woelfel. Tight RMR lower bounds for mutual exclusion and other problems. In ACM [ACM08], pages 217–226. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Ajtai:2002:ACI

[AJKS02] Miklós Ajtai, T. S. Jayram, Ravi Kumar, and D. Sivakumar. Approximate counting of inversions in a data stream. In ACM [ACM02], pages 370–379. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Aharonov:2006:PQA

[AJL06] Dorit Aharonov, Vaughan Jones, and Zeph Landau. A polynomial quantum algorithm for approximating the Jones polynomial. In ACM [ACM06], pages 427–436. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Alekhnovich:2002:ESB

[AJPU02] Michael Alekhnovich, Jan Johannsen, Toniann Pitassi, and Alasdair Urquhart. An exponential separation between regular and general resolution. In ACM [ACM02], pages 448–456. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Andrews:2007:SMW

[AJS07] Matthew Andrews, Kyomin Jung, and Alexander Stolyar. Stability of the max-weight routing and scheduling protocol in dynamic networks and at critical loads. In ACM [ACM07], pages 145–154. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Ajtai:2002:ILM

[Ajt02] Miklós Ajtai. The invasiveness of off-line memory checking. In ACM [ACM02], pages 504–513. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Ajtai:2003:WCB

[Ajt03]  Miklos Ajtai.  The worst-

case behavior of Schnorr’s algorithm approximating the shortest nonzero vector in a lattice. In ACM [ACM03], pages 396–406. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Ajtai:2004:CAP

[Ajt04] Miklós Ajtai. A conjecture about polynomial time computable lattice-lattice functions. In ACM [ACM04], pages 486–493. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Ajtai:2005:RHL

[Ajt05] Miklós Ajtai. Representing hard lattices with O(n log n) bits. In ACM [ACM05], pages 94–103. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Arora:2002:FAC

[AK02] Sanjeev Arora and Subhash Khot. Fitting algebraic curves to noisy data. In ACM [ACM02], pages 162–169. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Awerbuch:2004:ARE

[AK04] Baruch Awerbuch and Robert D. Kleinberg. Adaptive routing with end-to-end feedback: distributed learning and geometric approaches. In ACM [ACM04], pages 45–53. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Anshelevich:2007:TBM

[AK07a] Elliot Anshelevich and Adriana Karagiozova. Terminal backup, 3D matching, and covering cubic graphs. In ACM [ACM07], pages 391–400. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Arora:2007:CPD

[AK07b] Sanjeev Arora and Satyen Kale. A combinatorial, primal-dual approach to semidefinite programs. In ACM [ACM07], pages 227–236. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Awerbuch:2008:SDG

[AK08] Baruch Awerbuch and Rohit Khandekar. Stateless distributed gradient descent for positive linear programs. In ACM [ACM08], pages 691–700. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Anshelevich:2002:SLB

[AKK02] Elliot Anshelevich, David Kempe, and Jon Kleinberg. Stability of load balancing algorithms in dynamic adversarial systems. In ACM [ACM02], pages 399–406. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Arora:2008:UGE

[AKK^{+}08] Sanjeev Arora, Subhash A.

Khot, Alexandra Kolla, David Steurer, Madhur Tulsiani, and Nisheeth K. Vishnoi. Unique games on expanding constraint graphs are easy: extended abstract. In ACM [ACM08], pages 21–28. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Amir:2003:CFA

[AKR03] Eyal Amir, Robert Krauthgamer, and Satish Rao. Constant factor approximation of vertexcuts in planar graphs. In ACM [ACM03], pages 90–99. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Ajtai:2001:SAS

[AKS01] Miklós Ajtai, Ravi Kumar, and D. Sivakumar. A sieve algorithm for the shortest lattice vector problem. In ACM [ACM01], pages 601–610. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Aronov:2003:CTC

[AKS03] Boris Aronov, Vladlen Koltun, and Micha Sharir. Cutting triangular cycles of lines in space. In ACM [ACM03], pages 547–555. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Adler:2000:CUE

[AL00] Micah Adler and Tom Leighton. Compression using efficient multicasting. In ACM [ACM00], pages 153–162. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p153-adler/p153-adler.pdf;
                                         http://www.acm.org/pubs/
                                        citations/proceedings/stoc/
                                        335305/p153-adler/. ACM

order number 508000.

Albers:2002:ROS

[Alb02] Susanne Albers. On randomized online scheduling. In ACM [ACM02], pages 134–143. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Alekhnovich:2005:LBK

[Ale05] Michael Alekhnovich. Lower bounds for k-DNF resolution on random 3-CNFs. In ACM [ACM05], pages 251–256. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Arora:2005:EDS

[ALN05] Sanjeev Arora, James R. Lee, and Assaf Naor. Euclidean distortion and the sparsest cut. In ACM [ACM05], pages 553–562. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Achlioptas:2001:FCL

[AM01] Dimitris Achlioptas and Frank McSherry. Fast computation of low rank matrix approximations. In ACM [ACM01], pages 611–618. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Achlioptas:2002:AAG

[AM02] Dimitris Achlioptas and Cristopher Moore. Almost all graphs with average degree 4 are 3colorable. In ACM [ACM02], pages 199–208. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Alur:2004:VPL

[AM04] Rajeev Alur and P. Madhusudan. Visibly pushdown languages. In ACM [ACM04], pages 202–211. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Ambainis:2000:QLB

[Amb00] Andris Ambainis. Quantum lower bounds by quantum arguments. In ACM [ACM00], pages 636–643. ISBN 1-58113-184-

4. URL http://www.acm.org/
pubs/articles/proceedings/
stoc/335305/p636-ambainis/
p636-ambainis.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/335305/
p636-ambainis/. ACM order

number 508000.

Ambainis:2001:NPL

[Amb01] Andris Ambainis. A new protocol and lower bounds for quantum coin flipping. In ACM [ACM01], pages 134–142. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Ambainis:2004:QAD

[Amb04] Andris Ambainis. Quantum algorithms a decade after Shor. In ACM [ACM04], page 111. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Arya:2002:SEA

[AMM02] Sunil Arya, Theocharis Malamatos, and David M. Mount. Space-efficient approximate Voronoi diagrams. In ACM [ACM02], pages 721–730. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Arya:2006:II

[AMM06] Sunil Arya, Theocharis Malamatos, and David M. Mount. On the importance of idempotence. In ACM [ACM06], pages 564–573. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

                                        /portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Alon:2005:QFG

[AMMN05] Noga Alon, Konstantin Makarychev, Yury Makarychev, and Assaf Naor. Quadratic forms on graphs. In ACM [ACM05], pages 486–493. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Aleksandrov:2000:AAG

[AMS00] Lyudmil Aleksandrov, Anil Maheshwari, and Jörg-Rüdiger Sack. Approximation algorithms for geometric shortest path problems. In ACM [ACM00], pages 286– 295. ISBN 1-58113-184-4. URL

http://www.acm.org/pubs/
articles/proceedings/stoc/
335305/p286-aleksandrov/
p286-aleksandrov.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/335305/
p286-aleksandrov/. ACM or-

der number 508000.

Asano:2006:DTC

[AMT06] Tetsuo Asano, Jiří Matoušek, and Takeshi Tokuyama. The distance trisector curve. In ACM [ACM06], pages 336–343. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Achlioptas:2004:TPV

[AN04a] Dimitris Achlioptas and Assaf Naor. The two possible values of the chromatic number of a random graph. In ACM [ACM04], pages 587–593. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Alon:2004:ACN

[AN04b] Noga Alon and Assaf Naor. Approximating the cut-norm via Grothendieck’s inequality. In ACM [ACM04], pages 72–80. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Antonakopoulos:2006:FLE

[Ant06] Spyridon Antonakopoulos. Fast leader-election protocols with bounded cheaters’ edge. In ACM [ACM06], pages 187–196.

ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

                                        //portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Andoni:2009:AED

[AO09] Alexandr Andoni and Krzysztof Onak. Approximating edit distance in near-linear time. In ACM [ACM09], pages 199–204. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Achlioptas:2003:TRS

[AP03] Dimitris Achlioptas and Yuval Peres. The threshold for random k-SAT is 2^{k}(ln 2−O(k)). In ACM [ACM03], pages 223–231. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Andersen:2009:FSC

[AP09] Reid Andersen and Yuval Peres. Finding sparse cuts locally using evolving sets. In ACM [ACM09], pages 235–244. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Aronov:2003:DDT

[APST03] Boris Aronov, János Pach, Micha Sharir, and Gábor Tardos. Distinct distances in three and higher dimensions. In ACM [ACM03], pages 541–546. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Arai:2001:CAT

[APU01] Noriko Arai, Toniann Pitassi, and Alasdair Urquhart. The complexity of analytic tableaux. In ACM [ACM01], pages 356– 363. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Azar:2003:MMQ

[AR03] Yossi Azar and Yossi Richter.

Management of multi-queue switches in QoS networks. In ACM [ACM03], pages 82–89. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Azar:2004:ZOP

[AR04] Yossi Azar and Yossi Richter.

The zero-one principle for switching networks. In ACM [ACM04], pages 64–71. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Achlioptas:2006:SSG

[ART06] Dimitris Achlioptas and Federico Ricci-Tersenghi. On the solution-space geometry of random constraint satisfaction problems. In ACM [ACM06], pages 130–139. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Arora:2004:EFG

[ARV04] Sanjeev Arora, Satish Rao, and Umesh Vazirani. Expander flows, geometric embeddings and graph partitioning. In ACM [ACM04], pages 222–231. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Alon:2003:TSD

[AS03] Noga Alon and Asaf Shapira. Testing subgraphs in directed graphs. In ACM [ACM03], pages 700–709. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Albers:2004:PGA

[AS04] Susanne Albers and Markus Schmidt. On the performance of greedy algorithms in packet buffering. In ACM [ACM04], pages 35–44. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Alon:2005:EMG

[AS05] Noga Alon and Asaf Shapira. Every monotone graph property is testable. In ACM [ACM05], pages 128–137. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Asadpour:2007:AAM

[AS07] Arash Asadpour and Amin Saberi. An approximation algorithm for max-min fair allocation of indivisible goods. In ACM [ACM07], pages 114–121. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Ambainis:2006:NQL

[AŠdW06] Andris Ambainis, Robert Špalek, and Ronald de Wolf. A new quantum lower bound method,: with applications to direct product theorems and time-space tradeoffs. In ACM [ACM06], pages 618–633.

ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Aspnes:2002:WFC

[ASS02] James Aspnes, Gauri Shah, and Jatin Shah. Wait-free consensus with infinite arrivals. In ACM [ACM02], pages 524–533. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Ambainis:2000:CHM

[ASV00] Andris Ambainis, Leonard J.

Schulman, and Umesh V. Vazirani. Computing with highly mixed states (extended abstract). In ACM [ACM00], pages 697–704. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p697-ambainis/p697-ambainis.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p697-ambainis/

. ACM order number 508000.

Andersson:2000:TWC

[AT00] Arne A. Andersson and Mikkel Thorup. Tight(er) worst-case bounds on dynamic searching and priority queues. In ACM [ACM00], pages 335–342. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p335-andersson/p335-andersson.
pdf;   http://www.acm.org/
pubs/citations/proceedings/

                                        stoc/335305/p335-andersson/

. ACM order number 508000.

Aharonov:2003:AQS

[ATS03] Dorit Aharonov and Amnon Ta-Shma. Adiabatic quantum state generation and statistical zero knowledge. In ACM [ACM03], pages 20–29. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Aharonov:2000:QBE

[ATSVY00] Dorit Aharonov, Amnon Ta-Shma, Umesh V. Vazirani, and Andrew C. Yao. Quantum bit escrow. In ACM [ACM00], pages 705–714. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p705-aharonov/p705-aharonov.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/335305/p705-aharonov/

. ACM order number 508000.

Austrin:2007:BMS

[Aus07] Per Austrin. Balanced MAX 2-SAT might not be the hardest. In ACM [ACM07], pages 189–197. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Aaronson:2008:ANB

[AW08] Scott Aaronson and Avi Wigderson. Algebrization: a new barrier in complexity theory. In ACM [ACM08], pages 731–740. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Andrews:2005:HUC

[AZ05a] Matthew Andrews and Lisa Zhang. Hardness of the undirected congestion minimization problem. In ACM [ACM05], pages 284–293. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Andrews:2005:HUE

[AZ05b] Matthew Andrews and Lisa Zhang. Hardness of the undirected edge-disjoint paths problem. In ACM [ACM05], pages 276–283. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Andrews:2006:LHD

[AZ06] Matthew Andrews and Lisa Zhang. Logarithmic hardness of the directed congestion minimization problem. In ACM [ACM06], pages 517–526. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Bachmat:2002:ACA

[Bac02] E. Bachmat. Average case analysis for batched disk scheduling and increasing subsequences. In ACM [ACM02], pages 277–286. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Basu:2002:CBN

[Bas02]  Saugata Basu.  Computing

the Betti numbers of arrangements. In ACM [ACM02], pages 712–720. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Basu:2005:PTA

[Bas05]  Saugata Basu.  Polynomial

time algorithm for computing the top Betti numbers of semi-algebraic sets defined by quadratic inequalities. In ACM [ACM05], pages 313–322. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Basu:2007:CCM

[Bas07] Saugata Basu. Combinatorial complexity in O-minimal geometry. In ACM [ACM07], pages 47–56. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Ben-Aroya:2008:CCA

[BATS08] Avraham Ben-Aroya and Amnon Ta-Shma. A combinatorial construction of almost-Ramanujan graphs using the zig-zag product. In ACM [ACM08], pages 325–334. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Biham:2000:PSQ

[BBB^{+}00] Eli Biham, Michel Boyer, P. Oscar Boykin, Tal Mor, and Vwani Roychowdhury. A proof of the security of quantum key distribution (extended abstract). In ACM [ACM00], pages 715–724. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p715-biham/p715-biham.pdf;
 http://www.acm.org/pubs/
citations/proceedings/stoc/

335305/p715-biham/. ACM

order number 508000.

Bansal:2004:AAD

[BBCM04] Nikhil Bansal, Avrim Blum, Shuchi Chawla, and Adam Meyerson. Approximation algorithms for deadline-TSP and vehicle routing with timewindows. In ACM [ACM04], pages 166–174. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Bansal:2008:RCA

[BBN08] Nikhil Bansal, Niv Buchbinder, and Joseph (Seffi) Naor. Randomized competitive algorithms for generalized caching. In ACM [ACM08], pages 235– 244. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Babai:2009:PTT

[BBS09] László Babai, Robert Beals, and Ákos Seress. Polynomialtime theory of matrix groups. In ACM [ACM09], pages 55– 64. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Balcan:2008:DFC

[BBV08] Maria-Florina Balcan, Avrim Blum, and Santosh Vempala. A discriminative framework for clustering via similarity functions. In ACM [ACM08], pages 671–680. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Burgisser:2004:CCC

[BC04] Peter Bürgisser and Felipe Cucker. Counting complexity classes for numeric computations II: algebraic and semialgebraic sets. In ACM [ACM04], pages 475–485. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Bojanczyk:2005:TWA

[BC05] Mikolaj Bojanczyk and Thomas Colcombet. Tree-walking automata do not recognize all regular languages. In ACM [ACM05], pages 234–243. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Borgs:2007:FME

[BCDR07] Christian Borgs, Jennifer Chayes, Constantinos Daskalakis, and Sebastien Roch. First to market is not everything: an analysis of preferential attachment with fitness. In ACM [ACM07], pages 135–144. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Bansal:2006:QPU

[BCES06] Nikhil Bansal, Amit Chakrabarti, Amir Epstein, and Baruch Schieber. A quasi-PTAS for unsplittable flow on line graphs. In ACM [ACM06], pages 721–729. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Bateni:2009:MAD

[BCG09]  MohammadHossein  Bateni,

Moses Charikar, and Venkatesan Guruswami. MaxMin allocation via degree lowerbounded arborescences. In ACM [ACM09], pages 543–552. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Borgs:2008:MFT

[BCI^{+}08] Christian Borgs, Jennifer Chayes, Nicole Immorlica, Adam Tauman Kalai, Vahab Mirrokni, and Christos Papadimitriou. The myth of the Folk Theorem. In ACM [ACM08], pages 365–372. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Badoiu:2005:LDE

[BCIS05] Mihai Bădoiu, Julia Chuzhoy, Piotr Indyk, and Anastasios Sidiropoulos. Low-distortion embeddings of general metrics into the line. In ACM [ACM05], pages 225–233. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Borgs:2006:GLP

[BCL^{+}06] Christian Borgs, Jennifer Chayes, László Lovász, Vera T. Sós, Balázs Szegedy, and Katalin Vesztergombi. Graph limits and parameter testing. In ACM [ACM06], pages 261–270. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Beimel:2006:PAS

[BCNW06] Amos Beimel, Paz Carmi, Kobbi Nissim, and Enav Weinreb. Private approximation of search problems. In ACM [ACM06], pages 119–128. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Borgs:2001:STS

[BCP01] Christian Borgs, Jennifer Chayes, and Boris Pittel. Sharp threshold and scaling window for the integer partitioning problem. In ACM [ACM01], pages 330–336. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Bartal:2001:AMS

[BCR01] Yair Bartal, Moses Charikar, and Danny Raz. Approximating min-sum k-clustering in metric spaces. In ACM [ACM01], pages 11–20. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Berenbrink:2000:BAH

[BCSV00] Petra Berenbrink, Artur Czumaj, Angelika Steger, and Berthold Vöcking. Balanced allocations: the heavily loaded case. In ACM [ACM00], pages 745–754. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p745-berenbrink/p745-berenbrink.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p745-berenbrink/

. ACM order number 508000.

Boissonnat:2004:IIS

[BCSV04] Jean-Daniel Boissonnat, David Cohen-Steiner, and Gert Vegter. Isotopic implicit surface meshing. In ACM [ACM04], pages 301–309. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Berman:2000:ITM

[BD00] Piotr Berman and Bhaskar DasGupta. Improvements in throughout maximization for real-time scheduling. In ACM [ACM00], pages 680–687. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p680-berman/p680-berman.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p680-berman/.

ACM order number 508000.

Batu:2002:CAE

[BDKR02] Tuğkan Batu, Sanjoy Dasgupta, Ravi Kumar, and Ronitt Rubinfeld. The complexity of approximating entropy. In ACM [ACM02], pages 678–687. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Barenboim:2009:DCL

[BE09] Leonid Barenboim and Michael Elkin. Distributed (δ + 1)- coloring in linear (in δ ) time. In ACM [ACM09], pages 111–120. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Batu:2003:SAW

[BEK^{+}03] Tugkan Batu, Funda Ergün, Joe Kilian, Avner Magen, Sofya Raskhodnikova, Ronitt Rubinfeld, and Rahul Sami. A sublinear algorithm for weakly approximating edit distance. In ACM [ACM03], pages 316–324. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Brodal:2003:LCO

[BF03] Gerth Stølting Brodal and Rolf Fagerberg. On the limits of cache-obliviousness. In ACM [ACM03], pages 307–315. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Bayati:2007:SDA

[BGK^{+}07] Mohsen Bayati, David Gamarnik, Dimitriy Katz, Chandra Nair, and Prasad Tetali. Simple deterministic approximation algorithms for counting matchings. In ACM [ACM07], pages 122–127. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Bjorklund:2007:FMM

[BHKK07] Andreas Björklund, Thore Husfeldt, Petteri Kaski, and Mikko Koivisto. Fourier meets Möbius: fast subset convolution. In ACM [ACM07], pages 67–74. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Blum:2008:RMP

[BHLR08] Avrim Blum, MohammadTaghi Hajiaghayi, Katrina Ligett, and Aaron Roth. Regret minimization and the price of total anarchy. In ACM [ACM08], pages 373–382. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Badoiu:2002:ACC

[BHPI02] Mihai Bădoiu, Sariel Har-Peled, and Piotr Indyk. Approximate clustering via core-sets. In ACM [ACM02], pages 250–257. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Baswana:2002:IDA

[BHS02] Surender Baswana, Ramesh Hariharan, and Sandeep Sen. Improved decremental algorithms for maintaining transitive closure and all-pairs shortest paths. In ACM [ACM02], pages 117–123. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Busch:2000:HPR

[BHW00] Costas Busch, Maurice Herlihy, and Roger Wattenhofer. Hard-potato routing. In ACM [ACM00], pages 278–285. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p278-busch/p278-busch.pdf;
 http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p278-busch/. ACM

order number 508000.

Bienstock:2004:SFP

[BI04] D. Bienstock and G. Iyengar.

Solving fractional packing problems in O^{∗}(1/ϵ) iterations. In ACM [ACM04], pages 146–155. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Barkol:2008:CPR

[BIW08] Omer Barkol, Yuval Ishai, and Enav Weinreb. Communication in the presence of replication. In ACM [ACM08], pages 661–670. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Beame:2007:LBR

[BJR07] Paul Beame, T. S. Jayram, and Atri Rudra. Lower bounds for randomized read/write stream algorithms. In ACM [ACM07], pages 689–698. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Bodirsky:2008:CTC

[BK08] Manuel Bodirsky and Jan Kara. The complexity of temporal constraint satisfaction problems. In ACM [ACM08], pages 29–38. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Bernstein:2009:NOO

[BK09] Aaron Bernstein and David Karger. A nearly optimal oracle for avoiding failed vertices and edges. In ACM [ACM09], pages 101–110. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Bulatov:2001:CMC

[BKJ01] Andrei Bulatov, Andrei Krokhin, and Peter Jeavons. The complexity of maximal constraint languages. In ACM [ACM01], pages 667–674. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Buchsbaum:2003:OVL

[BKK^{+}03] Adam L. Buchsbaum, Howard Karloff, Claire Kenyon, Nick Reingold, and Mikkel Thorup. OPT versus LOAD in dynamic storage allocation. In ACM [ACM03], pages 556–564. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Brinkman:2007:VCR

[BKL07] Bo Brinkman, Adriana Karagiozova, and James R. Lee. Vertex cuts, random walks, and dimension reduction in seriesparallel graphs. In ACM [ACM07], pages 621–630. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Bansal:2008:AGD

[BKN08a] Nikhil Bansal, Rohit Khandekar, and Viswanath Nagarajan. Additive guarantees for degree bounded directed network design. In ACM [ACM08], pages 769–778. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Barto:2008:GPC

[BKN08b] Libor Barto, Marcin Kozik, and Todd Niven. Graphs, polymorphisms and the complexity of homomorphism problems. In ACM [ACM08], pages 789–796. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Batu:2004:SAT

[BKR04] Tugkan Batu, Ravi Kumar, and Ronitt Rubinfeld. Sublinear algorithms for testing monotone and unimodal distributions. In ACM [ACM04], pages 381–390. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Barak:2005:SIN

[BKS^{+}05] Boaz Barak, Guy Kindler, Ronen Shaltiel, Benny Sudakov, and Avi Wigderson. Simulating independence: new constructions of condensers, Ramsey graphs, dispersers, and extractors. In ACM [ACM05], pages 1–10. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Briest:2005:ATU

[BKV05] Patrick Briest, Piotr Krysta, and Berthold Vöcking. Approximation techniques for utilitarian mechanism design. In ACM [ACM05], pages 39–48. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Blum:2000:NTL

[BKW00] Avrim Blum, Adam Kalai, and Hal Wasserman. Noisetolerant learning, the parity problem, and the statistical query model. In ACM [ACM00], pages 435–440. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p435-blum/p435-blum.pdf;
                                        http://www.acm.org/pubs/
                                        citations/proceedings/stoc/

335305/p435-blum/. ACM or-

der number 508000.

Becchetti:2001:NCS

[BL01] Luca Becchetti and Stefano Leonardi. Non-clairvoyant scheduling to minimize the average flow time on single and parallel machines. In ACM [ACM01], pages 94–103. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Barak:2002:SPT

[BL02] Boaz Barak and Yehuda Lindell. Strict polynomial-time in simulation and extraction. In ACM [ACM02], pages 484–493. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Blais:2009:TJN

[Bla09] Eric Blais. Testing juntas nearly optimally. In ACM [ACM09], pages 151–158. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Brodal:2002:OFS

[BLM^{+}02] Gerth Stølting Brodal, George Lagogiannis, Christos Makris, Athanasios Tsakalidis, and Kostas Tsichlas. Optimal finger search trees in the pointer machine. In ACM [ACM02], pages 583–591. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Bartal:2003:MRT

[BLMN03] Yair Bartal, Nathan Linial, Manor Mendel, and Assaf Naor.

On metric Ramsey-type phenomena. In ACM [ACM03], pages 463–472. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Blum:2008:LTA

[BLR08] Avrim Blum, Katrina Ligett, and Aaron Roth. A learning theory approach to noninteractive database privacy. In ACM [ACM08], pages 609–618. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Buhrman:2000:BO

[BMRV00] H. Buhrman, P. B. Miltersen, J. Radhakrishnan, and S. Venkatesh. Are bitvectors optimal? In ACM [ACM00], pages 449–458. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p449-buhrman/p449-buhrman.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p449-buhrman/.

ACM order number 508000.

Bar-Noy:2000:UAA

[BNBYF^{+}00] Amotz Bar-Noy, Reuven Bar-Yehuda, Ari Freund, Joseph Naor, and Baruch Schieber. A unified approach to approximating resource allocation and scheduling. In ACM [ACM00], pages 735–744. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p735-bar-noy/p735-bar-noy.
pdf;   http://www.acm.org/

pubs/citations/proceedings/
stoc/335305/p735-bar-noy/.

ACM order number 508000.

Bogdanov:2005:PGL

[Bog05] Andrej Bogdanov. Pseudorandom generators for low degree polynomials. In ACM [ACM05], pages 21–30. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Ben-Or:2005:FQB

[BOH05] Michael Ben-Or and Avinatan Hassidim. Fast quantum Byzantine agreement. In ACM [ACM05], pages 481–485. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Boneh:2000:FSI

[Bon00] Dan Boneh. Finding smooth integers in short intervals using CRT decoding. In ACM [ACM00], pages 265–272. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p265-boneh/p265-boneh.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p265-boneh/. ACM

order number 508000.

Ben-Or:2006:BAF

[BOPV06] Michael Ben-Or, Elan Pavlov, and Vinod Vaikuntanathan. Byzantine agreement in the full-information model in O(log n) rounds. In ACM [ACM06], pages 179–186. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.

cfm?id=1132516. ACM order

number 508060.

Bansal:2003:SSN

[BP03] Nikhil Bansal and Kirk Pruhs. Server scheduling in the L_{p} norm: a rising tide lifts all boat. In ACM [ACM03], pages 242–250. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Basu:2005:CFB

[BPR05] Saugata Basu, Richard Pollack, and Marie-Françoise Roy. Computing the first Betti number and the connected components of semi-algebraic sets. In ACM [ACM05], pages 304–312. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Barkol:2000:TBN

[BR00] Omer Barkol and Yuval Rabani. Tighter bounds for nearest neighbor search and related problems in the cell probe model. In ACM [ACM00], pages 388–396. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p388-barkol/p388-barkol.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p388-barkol/.

ACM order number 508000.

Barak:2006:SDS

[BRSW06] Boaz Barak, Anup Rao, Ronen Shaltiel, and Avi Wigderson. 2-source dispersers for sub-polynomial entropy and Ramsey graphs beating the Frankl–Wilson construction. In ACM [ACM06], pages 671–680. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Ben-Sasson:2002:HEB

[BS02a] Eli Ben-Sasson. Hard examples for bounded depth Frege. In ACM [ACM02], pages 563–572. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Ben-Sasson:2002:SST

[BS02b] Eli Ben-Sasson. Size space tradeoffs for resolution. In ACM [ACM02], pages 457–464. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Bansal:2006:SCP

[BS06] Nikhil Bansal and Maxim Sviridenko. The Santa Claus problem. In ACM [ACM06], pages 31–40. ISBN 1-59593-134-1. LCCN QA75.5 .A22

2006. URL http://portal.
acm.org/citation.cfm?id=

1132516. ACM order number 508060.

Ben-Sasson:2004:RPP

[BSGH^{+}04] Eli Ben-Sasson, Oded Goldreich, Prahladh Harsha, Madhu Sudan, and Salil Vadhan. Robust PCPs of proximity, shorter PCPs and applications to coding. In ACM [ACM04], pages 1– 10. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Ben-Sasson:2003:SPH

[BSHR03] Eli Ben-Sasson, Prahladh Harsha, and Sofya Raskhodnikova. Some 3CNF properties are hard to test. In ACM [ACM03], pages 345–354. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Ben-Sasson:2009:ADS

[BSK09] Eli Ben-Sasson and Swastik Kopparty. Affine dispersers from subspace polynomials. In ACM [ACM09], pages 65–74. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Ben-Sasson:2005:SPP

[BSS05] Eli Ben-Sasson and Madhu Sudan. Simple PCPs with polylog rate and query complexity. In ACM [ACM05], pages 266–275. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Benjamini:2008:EMC

[BSS08] Itai Benjamini, Oded Schramm, and Asaf Shapira. Every minorclosed property of sparse graphs is testable. In ACM [ACM08], pages 393–402. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Batson:2009:TRS

[BSS09] Joshua D. Batson, Daniel A. Spielman, and Nikhil Srivastava. Twice-Ramanujan sparsifiers. In ACM [ACM09], pages 255–262. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Ben-Sasson:2003:REL

[BSSVW03] Eli Ben-Sasson, Madhu Sudan, Salil Vadhan, and Avi Wigderson. Randomness-efficient low degree tests and short PCPs via epsilon-biased sets. In ACM [ACM03], pages 612–621. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Benjamini:2005:BBF

[BSW05] Itai Benjamini, Oded Schramm, and David B. Wilson. Balanced Boolean functions that can be evaluated so that every input bit is unlikely to be read. In ACM [ACM05], pages 244–250. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Bobkov:2003:MLS

[BT03] Sergey Bobkov and Prasad Tetali. Modified log-Sobolev inequalities, mixing and hypercontractivity. In ACM [ACM03], pages 287–296. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Beame:2002:TST

[BV02a] Paul Beame and Erik Vee.

Time-space tradeoffs, multiparty communication complexity, and nearest-neighbor problems. In ACM [ACM02], pages 688–697. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Bertsimas:2002:SCP

[BV02b] Dimitris Bertsimas and Santosh Vempala. Solving convex programs by random walks. In ACM [ACM02], pages 109–115. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Beier:2003:RKE

[BV03] Rene Beier and Berthold Vöcking. Random knapsack in expected polynomial time. In ACM [ACM03], pages 232–241. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Beier:2004:TPW

[BV04] Rene Beier and Berthold Vöcking. Typical properties of winners and losers in discrete optimization. In ACM [ACM04], pages 343–352. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Bollig:2001:ROB

[BW01] Beate Bollig and Philipp Woelfel. A read-once branching program lower bound of Ω(2^{n/4}) for integer multiplication using universal. In ACM [ACM01], pages 419–424. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Bar-Yossef:2003:SLB

[BY03] Ziv Bar-Yossef. Sampling lower bounds via information theory. In ACM [ACM03], pages 335–344. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Braverman:2007:CNC

[BY07] Mark Braverman and Michael Yampolsky. Constructing noncomputable Julia sets. In ACM [ACM07], pages 709–716. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Bar-Yossef:2004:ESQ

[BYJK04] Ziv Bar-Yossef, T. S. Jayram, and Iordanis Kerenidis. Exponential separation of quantum and classical one-way communication complexity. In ACM [ACM04], pages 128–137. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Bar-Yossef:2001:SAL

[BYKS01] Ziv Bar-Yossef, Ravi Kumar, and D. Sivakumar. Sampling algorithms: lower bounds and applications. In ACM [ACM01], pages 266–275. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Childs:2003:EAS

[CCD^{+}03] Andrew M. Childs, Richard Cleve, Enrico Deotto, Edward Farhi, Sam Gutmann, and Daniel A. Spielman. Exponential algorithmic speedup by a quantum walk. In ACM [ACM03], pages 59–68. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Chakraborty:2008:NDV

[CCK08] Tanmoy Chakraborty, Julia Chuzhoy, and Sanjeev Khanna. Network design for vertex connectivity. In ACM [ACM08], pages 167–176. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Cai:2008:QLB

[CCL08] Jin-Yi Cai, Xi Chen, and Dong Li. A quadratic lower bound for the permanent and determinant problem over any characteristic ≠ 2. In ACM [ACM08], pages 491–498. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Chakrabarti:2008:RLB

[CCM08] Amit Chakrabarti, Graham Cormode, and Andrew Mc-Gregor. Robust lower bounds for communication and stream computation. In ACM [ACM08], pages 641–650. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Cryan:2002:PTA

[CD02] Mary Cryan and Martin Dyer. A polynomial-time algorithm to approximately count contingency tables when the number of rows is constant. In ACM [ACM02], pages 240–249. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Chen:2005:ADA

[CD05] Xi Chen and Xiaotie Deng. On algorithms for discrete and approximate Brouwer fixed points. In ACM [ACM05], pages 323–330. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Cramer:2000:CVS

[CDD00] Ronald Cramer, Ivan Damgård, and Stefan Dziembowski. On the complexity of verifiable secret sharing and multiparty computation. In ACM [ACM00], pages 325–334. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p325-cramer/p325-cramer.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p325-cramer/.

ACM order number 508000. Cole:2003:PNE

[CDR03] Richard Cole, Yevgeniy Dodis, and Tim Roughgarden. Pricing network edges for heterogeneous selfish users. In ACM [ACM03], pages 521–530. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Cryan:2005:ACI

[CDR05] Mary Cryan, Martin Dyer, and Dana Randall. Approximately counting integral flows and cellbounded contingency tables. In ACM [ACM05], pages 413–422. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Chambers:2009:HFC

[CEN09] Erin W. Chambers, Jeff Erickson, and Amir Nayyeri. Homology flows, cohomology cuts. In ACM [ACM09], pages 273–282.

ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Cooper:2002:CWG

[CF02] Colin Cooper and Alan Frieze. Crawling on Web graphs. In ACM [ACM02], pages 419–427. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Cole:2008:FCT

[CF08] Richard Cole and Lisa Fleischer. Fast-converging tatonnement algorithms for onetime and ongoing market problems. In ACM [ACM08], pages 315–324. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Charikar:2000:QSP

[CFG^{+}00] Moses Charikar, Ronald Fagin, Venkatesan Guruswami, Jon Kleinberg, Prabhakar Raghavan, and Amit Sahai. Query strategies for priced information (extended abstract). In ACM [ACM00], pages 582– 591. ISBN 1-58113-184-4.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/335305/p582-charikar/
                                        p582-charikar.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        proceedings/stoc/335305/
                                        p582-charikar/. ACM order

number 508000.

Cardinal:2009:EAP

[CFJ^{+}09] Jean Cardinal, Samuel Fiorini, Gwenaël Joret, Raphaël M. Jungers, and J. Ian Munro. An efficient algorithm for partial order production. In ACM [ACM09], pages 93–100. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Correa:2004:AKT

[CG04] José R. Correa and Michel X.

Goemans. An approximate König’s theorem for edgecoloring weighted bipartite graphs. In ACM [ACM04], pages 398–406. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Cole:2006:SDP

[CG06] Richard Cole and Lee-Ad Gottlieb. Searching dynamic point sets in spaces with bounded doubling dimension. In ACM [ACM06], pages 574–583. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Chalopin:2009:EPG

[CG09] Jérémie Chalopin and Daniel Gonçalves. Every planar graph is the intersection graph of segments in the plane: extended abstract. In ACM [ACM09], pages 631–638. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Canetti:2000:RZK

[CGGM00] Ran Canetti, Oded Goldreich, Shafi Goldwasser, and Silvio Micali. Resettable zero-knowledge (extended abstract). In ACM [ACM00], pages 235–244. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p235-canetti/p235-canetti.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/335305/p235-canetti/.

ACM order number 508000. Chuzhoy:2004:ACH

[CGH^{+}04] Julia Chuzhoy, Sudipto Guha, Eran Halperi, Sanjeev Khanna, Guy Kortsarz, and Joseph (Seffi) Nao. Asymmetric k-center is log^{∗}(n)-hard to approximate. In ACM [ACM04], pages 21– 27. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Chekuri:2004:MPS

[CGKK04] Chandra Chekuri, Ashish Goel, Sanjeev Khanna, and Amit Kumar. Multi-processor scheduling to minimize flow time with ϵ resource augmentation. In ACM [ACM04], pages 363–372. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Chadha:2009:CAM

[CGKM09] Jivitej S. Chadha, Naveen Garg, Amit Kumar, and V. N. Muralidhara. A competitive algorithm for minimizing weighted flow time on unrelatedmachines with speed augmentation. In ACM [ACM09], pages 679–684. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Chuzhoy:2007:HRC

[CGKT07] Julia Chuzhoy, Venkatesan Guruswami, Sanjeev Khanna, and Kunal Talwar. Hardness of routing with congestion in directed graphs. In ACM [ACM07], pages 165–178. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Cole:2004:DMI

[CGL04] Richard Cole, Lee-Ad Gottlieb, and Moshe Lewenstein. Dictionary matching and indexing with errors and don’t cares. In ACM [ACM04], pages 91–100. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Cleve:2009:EDT

[CGM^{+}09] Richard Cleve, Daniel Gottesman, Michele Mosca, Rolando D. Somma, and David Yonge-Mallo. Efficient discrete-time simulations of continuous-time quantum query algorithms. In ACM [ACM09], pages 409–416. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Chen:2008:PCM

[CGP08] Yiling Chen, Sharad Goel, and David M. Pennock. Pricing combinatorial markets for tournaments. In ACM [ACM08], pages 305–314. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Crepeau:2002:SMP

[CGS02] Claude Crépeau, Daniel Gottesman, and Adam Smith. Secure multi-party quantum computation. In ACM [ACM02], pages 643–652. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Cole:2000:FST

[CH00] Richard Cole and Ramesh Hariharan. Faster suffix tree construction with missing suffix links. In ACM [ACM00], pages 407–415. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p407-cole/p407-cole.pdf;
                                        http://www.acm.org/pubs/
                                        citations/proceedings/stoc/
                                        335305/p407-cole/. ACM or-

der number 508000.

Cole:2002:VCM

[CH02] Richard Cole and Ramesh Hariharan. Verifying candidate matches in sparse and wildcard matching. In ACM [ACM02], pages 592–601. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Cole:2003:FAC

[CH03] Richard Cole and Ramesh Hariharan. A fast algorithm for computing Steiner edge connectivity. In ACM [ACM03], pages 167–176. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Chan:2002:DSC

[Cha02a] Timothy M. Chan. Dynamic subgraph connectivity with geometric applications. In ACM [ACM02], pages 7–13. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Charikar:2002:SET

[Cha02b] Moses S. Charikar. Similarity estimation techniques from rounding algorithms. In ACM [ACM02], pages 380–388. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Chan:2007:MAA

[Cha07] Timothy M. Chan. More algorithms for all-pairs shortest paths in weighted graphs. In ACM [ACM07], pages 590–598. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Chen:2004:LFR

[CHKX04] Jianer Chen, Xiuzhen Huang, Iyad A. Kanj, and Ge Xia. Linear FPT reductions and computational lower bounds. In ACM [ACM04], pages 212–221. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Csirik:2000:SSA

[CJK^{+}00] Janos Csirik, David S. Johnson, Claire Kenyon, James B. Orlin, Peter W. Shor, and Richard R. Weber. On the sum-of-squares algorithm for bin packing. In ACM [ACM00], pages 208–217. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p208-csirik/p208-csirik.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p208-csirik/.

ACM order number 508000.

Chekuri:2002:ASP

[CK02] Chandra Chekuri and Sanjeev Khanna. Approximation schemes for preemptive weighted flow time. In ACM [ACM02], pages 297–305. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Charikar:2005:NUM

[CK05a] Moses Charikar and Adriana Karagiozova. On non-uniform multicommodity buy-at-bulk network design. In ACM [ACM05], pages 176–182. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Chlebus:2005:CAU

[CK05b] Bogdan S. Chlebus and Dariusz R. Kowalski. Cooperative asynchronous update of shared memory. In ACM [ACM05], pages 733–739. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Christodoulou:2005:PAF

[CK05c] George Christodoulou and Elias Koutsoupias. The price of anarchy of finite congestion games. In ACM [ACM05], pages 67– 73. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Chuzhoy:2006:HCP

[CK06] Julia Chuzhoy and Sanjeev Khanna. Hardness of cut problems in directed graphs. In ACM [ACM06], pages 527–536. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

                                        //portal.acm.org/citation.

cfm?id=1132516. ACM order

number 508060.

Chuzhoy:2007:PFC

[CK07] Julia Chuzhoy and Sanjeev Khanna. Polynomial flowcut gaps and hardness of directed cut problems. In ACM [ACM07], pages 179–188. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Choi:2008:OQC

[CK08] Sung-Soon Choi and Jeong Han Kim. Optimal query complexity bounds for finding graphs. In ACM [ACM08], pages 749–758. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Chen:2004:ATB

[CKL^{+}04] Jiangzhuo Chen, Robert D.

Kleinberg, László Lovász, Rajmohan Rajaraman, Ravi Sundaram, and Adrian Vetta. (almost) tight bounds and existence theorems for confluent flows. In ACM [ACM04], pages 529–538. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Canetti:2001:BBC

[CKPR01] Ran Canetti, Joe Kilian, Erez Petrank, and Alon Rosen. Black-box concurrent zeroknowledge requires Ω̃(log n) rounds. In ACM [ACM01], pages 570–579. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Chekuri:2004:ANM

[CKS04a] Chandra Chekuri, Sanjeev Khanna, and F. Bruce Shepherd. The all-or-nothing multicommodity flow problem. In ACM [ACM04], pages 156–165. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Chlebus:2004:CAR

[CKS04b] Bogdan S. Chlebus, Dariusz R. Kowalski, and Alexander A. Shvartsman. Collective asynchronous reading with polylogarithmic worst-case overhead. In ACM [ACM04], pages 321–330. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Chekuri:2005:MFW

[CKS05] Chandra Chekuri, Sanjeev Khanna, and F. Bruce Shepherd. Multicommodity flow, well-linked terminals, and routing problems. In ACM [ACM05], pages 183–192. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Chekuri:2006:EDP

[CKS06] Chandra Chekuri, Sanjeev Khanna, and F. Bruce Shepherd. Edge-disjoint paths in planar graphs with constant congestion. In ACM [ACM06], pages 757–766. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Czumaj:2002:STA

[CKV02] Artur Czumaj, Piotr Krysta, and Berthold Vöcking. Selfish traffic allocation for server farms. In ACM [ACM02], pages 287–296. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Chekuri:2001:AMW

[CKZ01] Chandra Chekuri, Sanjeev Khanna, and An Zhu. Algorithms for minimizing weighted flow time. In ACM [ACM01], pages 84–93. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Chazelle:2001:LBI

[CL01] Bernard Chazelle and Ding Liu.

Lower bounds for intersection searching and fractional cascading in higher dimension. In ACM [ACM01], pages 322–329. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Cicalese:2005:NSQ

[CL05] Ferdinando Cicalese and Eduardo Sany Laber. A new strategy for querying priced information. In ACM [ACM05], pages 674–683. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Cai:2007:HAA

[CL07] Jin-Yi Cai and Pinyan Lu.

Holographic algorithms: from art to science. In ACM [ACM07], pages 401–410. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Clarkson:2006:BTU

[Cla06] Kenneth L. Clarkson. Building triangulations using ϵ-nets. In ACM [ACM06], pages 326–335. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Charikar:2002:ASG

[CLL^{+}02] Moses Charikar, Eric Lehman, Ding Liu, Rina Panigrahy, Manoj Prabhakaran, April Rasala, Amit Sahai, and Abhi Shelat. Approximating the smallest grammar: Kolmogorov complexity in natural models. In ACM [ACM02], pages 792–801. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Chen:2008:FPA

[CLL^{+}08] Jianer Chen, Yang Liu, Songjian Lu, Barry O’Sullivan, and Igor Razgon. A fixed-parameter algorithm for the directed feedback vertex set problem. In ACM [ACM08], pages 177–186. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Chazelle:2003:SGA

[CLM03] Bernard Chazelle, Ding Liu, and Avner Magen. Sublinear geometric algorithms. In ACM [ACM03], pages 531–540. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Canetti:2002:UCT

[CLOS02] Ran Canetti, Yehuda Lindell, Rafail Ostrovsky, and Amit Sahai. Universally composable two-party and multi-party secure computation. In ACM [ACM02], pages 494–503. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Chechik:2009:FTS

[CLPR09] S. Chechik, M. Langberg, David Peleg, and L. Roditty. Fault-tolerant spanners for general graphs. In ACM [ACM09], pages 435–444. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Chan:2002:UAH

[CLTW02] Wun-Tat Chan, Tak-Wah Lam, Hing-Fung Ting, and Wai-Ha Wong. A unified analysis of hot video schedulers. In ACM [ACM02], pages 179–188. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Cai:2009:HPC

[CLX09] Jin-Yi Cai, Pinyan Lu, and Mingji Xia. Holant problems and counting CSP. In ACM [ACM09], pages 715–724. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Chen:2009:NAA

[CM09] Jing Chen and Silvio Micali. A new approach to auctions and resilient mechanism design. In ACM [ACM09], pages 503–512.

ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Charikar:2006:NOA

[CMM06] Moses Charikar, Konstantin Makarychev, and Yury Makarychev. Near-optimal algorithms for unique games. In ACM [ACM06], pages 205–214. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

                                        /portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Charikar:2009:IGS

[CMM09] Moses Charikar, Konstantin Makarychev, and Yury Makarychev. Integrality gaps for Sherali– Adams relaxations. In ACM [ACM09], pages 283–292. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Cote:2008:RSH

[CMP08] Aaron Coté, Adam Meyerson, and Laura Poplawski. Randomized k-server on hierarchical binary trees. In ACM [ACM08], pages 227–234. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Codenotti:2005:MEE

[CMV05] Bruno Codenotti, Benton Mc-Cune, and Kasturi Varadarajan. Market equilibrium via the excess demand function. In ACM [ACM05], pages 74–83. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Chuzhoy:2004:NHR

[CN04] Julia Chuzhoy and Joseph (Seffi)

Naor. New hardness results for congestion minimization and machine scheduling. In ACM [ACM04], pages 28–34. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Charikar:2003:BSA

[COP03] Moses Charikar, Liadan O’Callaghan, and Rina Panigrahy. Better streaming algorithms for clustering problems. In ACM [ACM03], pages 30–39. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Charikar:2001:CMS

[CP01] Moses Charikar and Rina Panigrahy. Clustering to minimize the sum of cluster diameters. In ACM [ACM01], pages 1– 10. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Chan:2007:VDC

[CP07] Timothy M. Chan and Mihai Pătraşcu. Voronoi diagrams in n · 2^{o(\sqrt{(lg lg n)})} time. In ACM [ACM07], pages 31–39. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Chien:2002:CAA

[CRS02] Steve Chien, Lars Rasmussen, and Alistair Sinclair. Clifford algebras and approximating the permanent. In ACM [ACM02], pages 222–231. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Chen:2003:MMA

[CRS03] Jiangzhuo Chen, Rajmohan Rajaraman, and Ravi Sundaram. Meet and merge: approximation algorithms for confluent flows. In ACM [ACM03], pages 373–382. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Capalbo:2002:RCC

[CRVW02] Michael Capalbo, Omer Reingold, Salil Vadhan, and Avi Wigderson. Randomness conductors and constant-degree lossless expanders. In ACM [ACM02], pages 659–668. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Czumaj:2000:NAA

[CS00] Artur Czumaj and Christian Scheideler. A new algorithm approach to the general Lovász local lemma with applications to scheduling and satisfiability problems (extended abstract). In ACM [ACM00], pages 38–47. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p38-czumaj/p38-czumaj.pdf;
 http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p38-czumaj/. ACM

order number 508000.

Coppersmith:2003:RCT

[CS03] Don Coppersmith and Madhu Sudan. Reconstructing curves in three (and higher) dimensional space from noisy data. In ACM [ACM03], pages 136–142. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Czumaj:2004:EWM

[CS04] Artur Czumaj and Christian Sohler. Estimating the weight of metric minimum spanning trees in sublinear-time. In ACM [ACM04], pages 175–183. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Carr:2000:RME

[CV00] Robert Carr and Santosh Vempala. Randomized metarounding (extended abstract). In ACM [ACM00], pages 58–62. ISBN 1-58113-184-4. URL

http://www.acm.org/pubs/
articles/proceedings/stoc/
335305/p58-carr/p58-carr.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p58-carr/. ACM

order number 508000.

Cheriyan:2005:AAN

[CV05] Joseph Cheriyan and Adrian Vetta. Approximation algorithms for network design with metric costs. In ACM [ACM05], pages 167–175. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Cheriyan:2002:AAM

[CVV02] Joseph Cheriyan, Santosh Vempala, and Adrian Vetta. Approximation algorithms for minimum-cost k-vertex connected subgraphs. In ACM [ACM02], pages 306–312. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Cheng:2009:DRG

[CW09a] Qi Cheng and Daqing Wan. A deterministic reduction for the gap minimum distance problem: [extended abstract]. In ACM [ACM09], pages 33–38. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Clarkson:2009:NLA

[CW09b] Kenneth L. Clarkson and David P. Woodruff. Numerical linear algebra in the streaming model. In ACM [ACM09], pages 205–214. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Chen:2000:SPQ

[CX00] Danny Z. Chen and Jinhui Xu. Shortest path queries in planar graphs. In ACM [ACM00], pages 469–478. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p469-chen/p469-chen.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p469-chen/. ACM or-

der number 508000.

deAlfaro:2001:QSO

[dAM01] Luca de Alfaro and Rupak Majumdar. Quantitative solution of omega-regular games. In ACM [ACM01], pages 675–683. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Dantchev:2007:RCG

[Dan07] Stefan S. Dantchev. Rank complexity gap for Lovász–Schrijver and sherali–adams proof systems. In ACM [ACM07], pages 311–317. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Dror:2003:TSP

[DELM03] Moshe Dror, Alon Efrat, Anna Lubiw, and Joseph S. B. Mitchell. Touring a sequence of polygons. In ACM [ACM03], pages 473–482. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Devanur:2004:SCM

[Dev04] Nikhil R. Devanur. The spending constraint model for market equilibrium: algorithmic, existence and uniqueness results. In ACM [ACM04], pages 519–528. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Dasgupta:2008:RPT

[DF08] Sanjoy Dasgupta and Yoav Freund. Random projection trees and low dimensional manifolds. In ACM [ACM08], pages 537–546. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Dinur:2006:FTB

[DFKO06] Irit Dinur, Ehud Friedgut, Guy Kindler, and Ryan O’Donnell. On the Fourier tails of bounded functions over the discrete cube. In ACM [ACM06], pages 437–446. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Damgard:2003:NIR

[DG03] Ivan Damgard and Jens Groth. Non-interactive and reusable non-malleable commitment schemes. In ACM [ACM03], pages 426–

437. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Dey:2003:ASF

[DGJ03] Tamal K. Dey, Joachim Giesen, and Matthias John. Alphashapes and flow shapes are homotopy equivalent. In ACM [ACM03], pages 493–502. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Dinur:2003:NMP

[DGKR03] Irit Dinur, Venkatesan Guruswami, Subhash Khot, and Oded Regev. A new multilayered PCP and the hardness of hypergraph vertex cover. In ACM [ACM03], pages 595–601. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Dinur:2008:DGH

[DGKS08] Irit Dinur, Elena Grigorescu, Swastik Kopparty, and Madhu Sudan. Decodability of group homomorphisms beyond the Johnson bound. In ACM [ACM08], pages 275–284. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Daskalakis:2006:CCN

[DGP06] Constantinos Daskalakis, Paul W.

Goldberg, and Christos H. Papadimitriou. The complexity of computing a Nash equilibrium. In ACM [ACM06], pages 71– 78. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Dunagan:2007:ICP

[DH07] John Dunagan and Nicholas J. A. Harvey. Iteratively constructing preconditioners via the conjugate gradient method. In ACM [ACM07], pages 207– 216. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Demetrescu:2003:NAD

[DI03] Camil Demetrescu and Giuseppe F.

Italiano. A new approach to dynamic all pairs shortest paths. In ACM [ACM03], pages 159–166. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Dubrov:2006:RCE

[DI06] Bella Dubrov and Yuval Ishai.

On the randomness complexity of efficient sampling. In ACM [ACM06], pages 711–720. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Dinur:2006:PTG

[Din06] Irit Dinur. The PCP theorem by gap amplification. In ACM [ACM06], pages 241–250. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Das:2008:ASS

[DK08] Abhimanyu Das and David Kempe. Algorithms for subset selection in linear regression. In ACM [ACM08], pages 45– 54. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Dodis:2009:CAI

[DKL09] Yevgeniy Dodis, Yael Tauman Kalai, and Shachar Lovett. On cryptography with auxiliary input. In ACM [ACM09], pages 621–630. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

DeMarzo:2006:OTA

[DKM06] Peter DeMarzo, Ilan Kremer, and Yishay Mansour. Online trading algorithms and robust option pricing. In ACM [ACM06], pages 477–486. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.

cfm?id=1132516. ACM order

number 508060.

Drineas:2002:CRS

[DKR02] Petros Drineas, Iordanis Kerenidis, and Prabhakar Raghavan. Competitive recommendation systems. In ACM [ACM02], pages 82–90. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

De:2008:FIM

[DKSS08] Anindya De, Piyush P. Kurur, Chandan Saha, and Ramprasad Saptharishi. Fast integer multiplication using modular arithmetic. In ACM [ACM08], pages 499–506. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Devanur:2006:IGS

[DKSV06] Nikhil R. Devanur, Subhash A. Khot, Rishi Saket, and Nisheeth K. Vishnoi. Integrality gaps for sparsest cut and minimum linear arrangement problems. In ACM [ACM06], pages 537–546. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Dwork:2009:DPR

[DL09] Cynthia Dwork and Jing Lei.

Differential privacy and robust statistics. In ACM [ACM09], pages 371–380. ISBN 1-60558-613-7. LCCN QA75.5 .A22

Durand:2001:CT

[DLS01] Bruno Durand, Leonid Levin, and Alexander Shen. Complex tilings. In ACM [ACM01], pages 732–739. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

delaVega:2003:ASC

[dlVKKR03] W. Fernandez de la Vega, Marek Karpinski, Claire Kenyon, and Yuval Rabani. Approximation schemes for clustering problems. In ACM [ACM03], pages 50–58. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

delaVega:2005:TDA

[dlVKKV05] W. Fernandez de la Vega, Marek Karpinski, Ravi Kannan, and Santosh Vempala. Tensor decomposition and approximation schemes for constraint satisfaction problems. In ACM [ACM05], pages 747–754. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Dziembowski:2002:TSP

[DM02] Stefan Dziembowski and Ueli Maurer. Tight security proofs for the bounded-storage model. In ACM [ACM02], pages 341– 350. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Daskalakis:2006:OPR

[DMR06a] Constantinos Daskalakis, Elchanan Mossel, and Sébastien Roch.

Optimal phylogenetic reconstruction. In ACM [ACM06], pages 159–168. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Dinur:2006:CHA

[DMR06b] Irit Dinur, Elchanan Mossel, and Oded Regev. Conditional hardness for approximate coloring. In ACM [ACM06], pages 344–353. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Dwork:2007:PPL

[DMT07] Cynthia Dwork, Frank McSherry, and Kunal Talwar. The price of privacy and the limits of LP decoding. In ACM [ACM07], pages 85–94. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Dobzinski:2007:LVB

[DN07] Shahar Dobzinski and Noam Nisan. Limitations of VCGbased mechanisms. In ACM [ACM07], pages 338–344. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Dwork:2009:CDP

[DNR^{+}09] Cynthia Dwork, Moni Naor, Omer Reingold, Guy N. Rothblum, and Salil Vadhan. On the complexity of differentially private data release: efficient algorithms and hardness results. In ACM [ACM09], pages 381–390. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Dobzinski:2005:AAC

[DNS05] Shahar Dobzinski, Noam Nisan, and Michael Schapira. Approximation algorithms for combinatorial auctions with complement-free bidders. In ACM [ACM05], pages 610–618. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Dobzinski:2006:TRM

[DNS06] Shahar Dobzinski, Noam Nisan, and Michael Schapira. Truthful randomized mechanisms for combinatorial auctions. In ACM [ACM06], pages 644–652. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

                                        //portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Daskalakis:2009:OPN

[DP09] Constantinos Daskalakis and Christos H. Papadimitriou. On oblivious PTAS’s for Nash equilibrium. In ACM [ACM09], pages 75–84. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Deng:2002:CE

[DPS02] Xiaotie Deng, Christos Papadimitriou, and Shmuel Safra. On the complexity of equilibria. In ACM [ACM02], pages 67– 71. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Dinur:2002:IBB

[DS02a] Irit Dinur and Shmuel Safra.

The importance of being biased. In ACM [ACM02], pages 33– 42. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Dwork:2002:RZK

[DS02b] Cynthia Dwork and Larry Stockmeyer. 2-round zero knowledge and proof auditors. In ACM [ACM02], pages 322–331. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Dodis:2005:CEL

[DS05a] Yevgeniy Dodis and Adam Smith. Correcting errors without leaking partial information. In ACM [ACM05], pages 654–663. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Dvir:2005:LDC

[DS05b] Zeev Dvir and Amir Shpilka.

Locally decodable codes with 2 queries and polynomial identity testing for depth 3 circuits. In ACM [ACM05], pages 592–601. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Daitch:2008:FAL

[DS08] Samuel I. Daitch and Daniel A.

Spielman. Faster approximate lossy generalized flow via interior point algorithms. In ACM [ACM08], pages 451–460. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Donovan:2007:DCN

[DSVW07] P. Donovan, B. Shepherd, A. Vetta, and G. Wilfong. Degree-constrained network flows. In ACM [ACM07], pages 681–688. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

DiCrescenzo:2000:ZKP

[DSY00] Giovanni Di Crescenzo, Kouichi Sakurai, and Moti Yung. On zero-knowledge proofs (extended abstract): “from membership to decision”. In ACM [ACM00], pages 255– 264. ISBN 1-58113-184-4. URL

http://www.acm.org/pubs/
articles/proceedings/stoc/
335305/p255-di_crescenzo/
p255-di_crescenzo.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p255-di_crescenzo/.

ACM order number 508000.

Dvir:2008:HRT

[DSY08] Zeev Dvir, Amir Shpilka, and Amir Yehudayoff. Hardnessrandomness tradeoffs for bounded depth arithmetic circuits. In ACM [ACM08], pages 741–748. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Dunagan:2001:OOR

[DV01] John Dunagan and Santosh Vempala. Optimal outlier removal in high-dimensional. In ACM [ACM01], pages 627–636. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Dunagan:2004:SPT

[DV04] John Dunagan and Santosh Vempala. A simple polynomialtime rescaling algorithm for solving linear programs. In ACM [ACM04], pages 315–320. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Deshpande:2007:SBD

[DV07] Amit Deshpande and Kasturi Varadarajan. Sampling-based dimension reduction for subspace approximation. In ACM [ACM07], pages 641–650. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Dietzfelbinger:2003:ARG

[DW03] Martin Dietzfelbinger and Philipp Woelfel. Almost random graphs with simple hash functions. In ACM [ACM03], pages 629–638. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Dietzfelbinger:2009:TLB

[DW09a] Martin Dietzfelbinger and Philipp Woelfel. Tight lower bounds for greedy routing in uniform small world rings. In ACM [ACM09], pages 591–600. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Dodis:2009:NME

[DW09b] Yevgeniy Dodis and Daniel Wichs. Non-malleable extractors and symmetric key cryptography from weak secrets. In ACM [ACM09], pages 601–610.

ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Dyer:2003:ACD

[Dye03] Martin Dyer. Approximate counting by dynamic programming. In ACM [ACM03], pages 693–699. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Even-dar:2009:CRM

[EdMN09] Eyal Even-dar, Yishay Mansour, and Uri Nadav. On the convergence of regret minimization dynamics in concave games. In ACM [ACM09], pages 523–532. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Elkin:2005:LSS

[EEST05] Michael Elkin, Yuval Emek, Daniel A. Spielman, and Shang-Hua Teng. Lower-stretch spanning trees. In ACM [ACM05], pages 494–503. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Efremenko:2009:QLD

[Efr09] Klim Efremenko. 3-query locally decodable codes of subexponential length. In ACM [ACM09], pages 39–44. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Eiter:2002:NRM

[EGM02] Thomas Eiter, Georg Gottlob, and Kazuhisa Makino. New results on monotone dualization and generating hypergraph transversals. In ACM [ACM02], pages 14–22. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Even:2000:IAC

[EGS00] Guy Even, Sudipto Guha, and Baruch Schieber. Improved approximations of crossings in graph drawings. In ACM [ACM00], pages 296–305. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p296-even/p296-even.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p296-even/. ACM or-

der number 508000.

Elkin:2002:CLA

[EK02] Michael Elkin and Guy Kortsarz. Combinatorial logarithmic approximation algorithm for directed telephone broadcast problem. In ACM [ACM02], pages 438–447. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Elkies:2001:ECM

[Elk01] Noam D. Elkies. Excellent codes from modular curves. In ACM [ACM01], pages 200–208. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Elkin:2004:ULB

[Elk04] Michael Elkin. Unconditional lower bounds on the timeapproximation tradeoffs for the distributed minimum spanning tree problem. In ACM [ACM04], pages 331–340. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Edelsbrunner:2000:SCS

[ELM^{+}00] Herbert Edelsbrunner, Xiang-Yang Li, Gary Miller, Andreas Stathopoulos, Dafna Talmor, Shang-Hua Teng, Alper Üngör, and Noel Walkington. Smoothing and cleaning up slivers. In ACM [ACM00], pages 273– 277. ISBN 1-58113-184-4. URL

http://www.acm.org/pubs/
articles/proceedings/stoc/
335305/p273-edelsbrunner/
p273-edelsbrunner.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p273-edelsbrunner/.

ACM order number 508000.

Elkin:2001:EBS

[EP01] Michael Elkin and David Peleg. (1 + ϵ, β)-spanner constructions for general graphs. In ACM [ACM01], pages 173–182. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Englert:2007:RBG

[ERW07] Matthias Englert, Harald Räcke, and Matthias Westermann. Reordering buffers for general metric spaces. In ACM [ACM07], pages 556–564. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Ergun:2001:BDF

[ESSS01] Funda Ergun, S. Cenk Sahinalp, Jonathan Sharp, and Rakesh Sinha. Biased dictionaries with fast insert/deletes. In ACM [ACM01], pages 483–491. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Feige:2002:RBA

[Fei02] Uriel Feige. Relations between average case complexity and approximation complexity. In ACM [ACM02], pages 534–543. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Feige:2004:SIR

[Fei04] Uriel Feige. On sums of independent random variables with unbounded variance, and estimating the average degree in a graph. In ACM [ACM04], pages 594–603. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Feige:2006:MWW

[Fei06] Uriel Feige. On maximizing welfare when utility functions are subadditive. In ACM [ACM06], pages 41–50. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Feldman:2006:HAT

[Fel06]  Vitaly Feldman.  Hardness

of approximate two-level logic minimization and PAC learning with membership queries. In ACM [ACM06], pages 363–372. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Feldman:2008:ELA

[Fel08] Vitaly Feldman. Evolvability from learning algorithms. In ACM [ACM08], pages 619–628. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Feldman:2009:PC

[FFKN09] Dan Feldman, Amos Fiat, Haim Kaplan, and Kobbi Nissim. Private coresets. In ACM [ACM09], pages 361–370. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Fatourou:2003:TTL

[FFR03] Panagiota Fatourou, Faith Fich, and Eric Ruppert. A tight time lower bound for space-optimal implementations of multi-writer snapshots. In ACM [ACM03], pages 259–268. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Fatourou:2006:TST

[FFR06] Panagiota Fatourou, Faith Ellen Fich, and Eric Ruppert. Timespace tradeoffs for implementations of snapshots. In ACM [ACM06], pages 169–178. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.

cfm?id=1132516. ACM order

number 508060.

Flaxman:2005:ACP

[FFV05] Abraham D. Flaxman, Alan M.

Frieze, and Juan C. Vera. On the average case performance of some greedy approximation algorithms for the uncapacitated facility location problem. In ACM [ACM05], pages 441–449. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Fiat:2002:CGA

[FGHK02] Amos Fiat, Andrew V. Goldberg, Jason D. Hartline, and Anna R. Karlin. Competitive generalized auctions. In ACM [ACM02], pages 72–81. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Feige:2000:ADN

[FHK00] Uriel Feige, Magnús M. Halldórsson, and Guy Kortsarz. Approximating the domatic number. In ACM [ACM00], pages 134–143. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p134-feige/p134-feige.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p134-feige/. ACM

order number 508000.

Feige:2005:IAA

[FHL05] Uriel Feige, MohammadTaghi Hajiaghayi, and James R. Lee. Improved approximation algorithms for minimum-weight vertex separators. In ACM [ACM05], pages 563–572. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Fleischer:2000:IAS

[FI00] Lisa Fleischer and Satoru Iwata. Improved algorithms for submodular function minimization and submodular flow. In ACM [ACM00], pages 107–116. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p107-fleischer/p107-fleischer.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p107-fleischer/

. ACM order number 508000.

Finocchi:2004:SSP

[FI04] Irene Finocchi and Giuseppe F. Italiano. Sorting and searching in the presence of memory faults (without redundancy). In ACM [ACM04], pages 101–110. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Friedl:2003:HTO

[FIM^{+}03] Katalin Friedl, Gábor Ivanyos, Frédéric Magniez, Miklos Santha, and Pranab Sen. Hidden translation and orbit coset in quantum computing. In ACM [ACM03], pages 1–9. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Fischer:2004:DTI

[Fis04] Eldar Fischer. The difficulty of testing for isomorphism against a graph that is given in advance. In ACM [ACM04], pages 391–397. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Friedl:2005:ETG

[FIS05] Katalin Friedl, Gábor Ivanyos, and Miklos Santha. Efficient testing of groups. In ACM [ACM05], pages 157–166. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Fagin:2000:RWB

[FKK^{+}00] Ronald Fagin, Anna R. Karlin, Jon Kleinberg, Prabhakar Raghavan, Sridhar Rajagopalan, Ronitt Rubinfeld, Madhu Sudan, and Andrew Tomkins. Random walks with “back buttons” (extended abstract). In ACM [ACM00], pages 484–493. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p484-fagin/p484-fagin.pdf;
 http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p484-fagin/. ACM

order number 508000.

Fleischer:2006:SCS

[FKLS06] Lisa Fleischer, Jochen Könemann, Stefano Leonardi, and Guido Schäfer. Simple cost sharing schemes for multicommodity rent-or-buy and stochastic Steiner tree. In ACM [ACM06], pages 663–670. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.

                                        cfm?id=1132516. ACM order

number 508060.

Feige:2000:AMB

[FKN00] Uriel Feige, Robert Krauthgamer, and Kobbi Nissim. Approximating the minimum bisection size (extended abstract). In ACM [ACM00], pages 530–536. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p530-feige/p530-feige.pdf;
                                         http://www.acm.org/pubs/
                                        citations/proceedings/stoc/
                                        335305/p530-feige/. ACM

order number 508000.

Fakcharoenphol:2008:AAV

[FL08] Jittat Fakcharoenphol and Bundit Laekhanukit. An o(log^{2} k)-approximation algorithm for the k-vertex connected spanning subgraph problem. In ACM [ACM08], pages 153–158. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Fischer:2002:MTG

[FLN^{+}02] Eldar Fischer, Eric Lehman, Ilan Newman, Sofya Raskhodnikova, Ronitt Rubinfeld, and Alex Samorodnitsky. Monotonicity testing over general poset domains. In ACM [ACM02], pages 474–483. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Fiat:2000:BAU

[FM00a] Amos Fiat and Manor Mendel.

Better algorithms for unfair metrical task systems and applications. In ACM [ACM00], pages 725–734. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p725-fiat/p725-fiat.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p725-fiat/. ACM or-

der number 508000.

Fitzi:2000:PCG

[FM00b] Mattias Fitzi and Ueli Maurer. From partial consistency to global broadcast. In ACM [ACM00], pages 494–503. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p494-fitzi/p494-fitzi.pdf;
 http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p494-fitzi/. ACM

order number 508000.

Feige:2006:FSB

[FM06] Uriel Feige and Mohammad Mahdian. Finding small balanced separators. In ACM [ACM06], pages 375–384. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Franceschini:2007:OSS

[FM07] Gianni Franceschini and S. Muthukrishnan. Optimal suffix selection. In ACM [ACM07], pages 328–337. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Feder:2000:CMU

[FMP^{+}00] Tomas Feder, Rajeev Motwani, Rina Panigrahy, Chris Olston, and Jennifer Widom. Computing the median with uncertainty. In ACM [ACM00], pages 602–607. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p602-feder/p602-feder.pdf;
                                         http://www.acm.org/pubs/
                                        citations/proceedings/stoc/
                                        335305/p602-feder/. ACM

order number 508000.

Feder:2000:FLP

[FMS00] Tomas Feder, Rajeev Motwani, and Carlos Subi. Finding long paths and cycles in sparse Hamiltonian graphs. In ACM [ACM00], pages 524–529. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p524-feder/p524-feder.pdf;
                                         http://www.acm.org/pubs/
                                        citations/proceedings/stoc/
                                        335305/p524-feder/. ACM

order number 508000.

Fischer:2001:TMP

[FN01] Eldar Fischer and Ilan Newman. Testing of matrix properties. In ACM [ACM01], pages 286–295. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Fischer:2005:TVE

[FN05] Eldar Fischer and Ilan Newman. Testing versus estimation of graph properties. In ACM [ACM05], pages 138–146. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Fortune:2000:ECI

[For00] Steven Fortune. Exact computations of the inertia symmetric integer matrices. In ACM [ACM00], pages 556–564. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p556-fortune/p556-fortune.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p556-fortune/.

ACM order number 508000.

Fortnow:2005:BNW

[For05] Lance Fortnow. Beyond NP:

the work and legacy of Larry Stockmeyer. In ACM [ACM05], pages 120–127. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Feigenbaum:2000:SCM

[FPS00] Joan Feigenbaum, Christos Papadimitriou, and Scott Shenker. Sharing the cost of muliticast transmissions (preliminary version). In ACM [ACM00], pages 218–227. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p218-feigenbaum/p218-feigenbaum.
pdf;   http://www.acm.org/
pubs/citations/proceedings/

                                        stoc/335305/p218-feigenbaum/

. ACM order number 508000.

Fabrikant:2004:CPN

[FPT04] Alex Fabrikant, Christos Papadimitriou, and Kunal Talwar. The complexity of pure Nash equilibria. In ACM [ACM04], pages 604–612. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Friedman:2003:PAS

[Fri03] Joel Friedman. A proof of Alon’s second eigenvalue conjecture. In ACM [ACM03], pages 720–724. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Fellows:2006:CWM

[FRRS06] Michael R. Fellows, Frances A. Rosamond, Udi Rotics, and Stefan Szeider. Clique-width minimization is NP-hard. In ACM [ACM06], pages 354–362. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

                                        //portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Fakcharoenphol:2003:TBA

[FRT03] Jittat Fakcharoenphol, Satish Rao, and Kunal Talwar. A tight bound on approximating arbitrary metrics by tree metrics. In ACM [ACM03], pages 448–455. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Fischer:2006:FCW

[FRV06] Simon Fischer, Harald Räcke, and Berthold Vöcking. Fast convergence to Wardrop equilibria by adaptive sampling methods. In ACM [ACM06], pages 653–662. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Feige:2001:IRS

[FS01] Uriel Feige and Gideon Schechtman. On the integrality ratio of semidefinite relaxations of MAX CUT. In ACM [ACM01], pages 433–442. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Frahling:2005:CDG

[FS05] Gereon Frahling and Christian Sohler. Coresets in dynamic geometric data streams. In ACM [ACM05], pages 209–217. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Fortnow:2008:IIC

[FS08] Lance Fortnow and Rahul Santhanam. Infeasibility of instance compression and succinct PCPs for NP. In ACM [ACM08], pages 133–142. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Friedrich:2009:NPL

[FS09] Tobias Friedrich and Thomas Sauerwald. Near-perfect load balancing by randomized rounding. In ACM [ACM09], pages 121–130. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Fortnow:2005:HSC

[FST05] Lance Fortnow, Rahul Santhanam, and Luca Trevisan. Hierarchies for semantic classes. In ACM [ACM05], pages 348– 355. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Furer:2000:APC

[Für00] Martin Fürer. Approximating permanents of complex matrices. In ACM [ACM00], pages 667–669. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p667-furer/p667-furer.pdf;
                                         http://www.acm.org/pubs/
                                        citations/proceedings/stoc/
                                        335305/p667-furer/. ACM

order number 508000.

Furer:2007:FIM

[Für07] Martin Fürer. Faster integer multiplication. In ACM [ACM07], pages 57–66. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Frieze:2008:LRG

[FVV08] Alan Frieze, Santosh Vempala, and Juan Vera. Logconcave random graphs. In ACM [ACM08], pages 779–788. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Gabow:2004:FPC

[Gab04] Harold N. Gabow. Finding paths and cycles of superpolylogarithmic length. In ACM [ACM04], pages 407–416. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Gacs:2001:CSS

[Gác01] Peter Gács. Compatible sequences and a slow Winkler percolation. In ACM [ACM01], pages 694–703. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Gacs:2002:CSR

[Gác02]  Peter Gács.  Clairvoyant

scheduling of random walks. In ACM [ACM02], pages 99–108. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Gafni:2009:EBS

[Gaf09] Eli Gafni. The extended BGsimulation and the characterization of t-resiliency. In ACM [ACM09], pages 85–92. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Garg:2005:SEA

[Gar05] Naveen Garg. Saving an epsilon: a 2-approximation for the k-MST problem in graphs. In ACM [ACM05], pages 396–402. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Gavinsky:2008:CIC

[Gav08] Dmitry Gavinsky. Classical interaction cannot replace a quantum message. In ACM [ACM08], pages 95–102. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Gordon:2008:CFS

[GCKL08] Dov S. Gordon, Hazay Carmit, Jonathan Katz, and Yehuda Lindell. Complete fairness in secure two-party computation. In ACM [ACM08], pages 413–422. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Gentry:2009:FHE

[Gen09] Craig Gentry. Fully homomorphic encryption using ideal lattices. In ACM [ACM09], pages 169–178. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Goldwasser:2007:VDC

[GGH^{+}07] Shafi Goldwasser, Dan Gutfreund, Alexander Healy, Tali Kaufman, and Guy N. Rothblum. Verifying and decoding in constant depth. In ACM [ACM07], pages 440–449. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Goldwasser:2008:CAP

[GGH^{+}08] Shafi Goldwasser, Dan Gutfreund, Alexander Healy, Tali Kaufman, and Guy N. Rothblum. A (de)constructive approach to program checking. In ACM [ACM08], pages 143–152. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Gilbert:2002:NOS

[GGI^{+}02a] A. C. Gilbert, S. Guha, P. Indyk, S. Muthukrishnan, and M. Strauss. Near-optimal sparse Fourier representations via sampling. In ACM [ACM02], pages 152–161. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Gilbert:2002:FSS

[GGI^{+}02b] Anna C. Gilbert, Sudipto Guha, Piotr Indyk, Yannis Kotidis, S. Muthukrishnan, and Martin J. Strauss. Fast, small-space algorithms for approximate histogram maintenance. In ACM [ACM02], pages 389–398. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Gennaro:2003:LBE

[GGK03] Rosario Gennaro, Yael Gertner, and Jonathan Katz. Lower bounds on the efficiency of encryption and digital signature schemes. In ACM [ACM03], pages 417–425. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Gafni:2005:SIA

[GGP05] Eli Gafni, Rachid Guerraoui, and Bastian Pochon. From a static impossibility to an adaptive lower bound: the complexity of early deciding set agreement. In ACM [ACM05], pages 714–722. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Gopalan:2009:LDT

[GGR09] Parikshit Gopalan, Venkatesan Guruswami, and Prasad Raghavendra. List decoding tensor products and interleaved codes. In ACM [ACM09], pages 13–22. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Guruswami:2002:NOL

[GI02] Venkatesan Guruswami and Piotr Indyk. Near-optimal lineartime codes for unique decoding and new list-decodable codes over smaller alphabets. In ACM [ACM02], pages 812–821. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Guruswami:2003:LTE

[GI03] Venkatesan Guruswami and Piotr Indyk. Linear time encodable and list decodable codes. In ACM [ACM03], pages 126–135. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Gennaro:2001:RCV

[GIKR01] Rosario Gennaro, Yuval Ishai, Eyal Kushilevitz, and Tal Rabin. The round complexity of verifiable secret sharing and secure multicast. In ACM [ACM01], pages 580–589. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Goldberg:2007:ITP

[GJ07] Leslie Ann Goldberg and Mark Jerrum. Inapproximability of the Tutte polynomial. In ACM [ACM07], pages 459–468. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Gilbert:2003:FBT

[GK03] Anna Gilbert and Howard Karloff. On the fractal behavior of TCP. In ACM [ACM03], pages 297–306. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Garg:2004:AAM

[GK04] Rahul Garg and Sanjiv Kapoor.

Auction algorithms for market equilibrium. In ACM [ACM04], pages 511–518. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Garg:2006:MAF

[GK06] Naveen Garg and Amit Kumar. Minimizing average flow time on related machines. In ACM [ACM06], pages 730–738. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Gupta:2009:CFA

[GK09] Anupam Gupta and Amit Kumar. A constant-factor approximation for stochastic Steiner forest. In ACM [ACM09], pages 659–668. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Gupta:2001:PVP

[GKK^{+}01] Anupam Gupta, Jon Kleinberg, Amit Kumar, Rajeev Rastogi, and Bulent Yener. Provisioning a virtual private network: a network design problem for multicommodity flow. In ACM [ACM01], pages 389–398. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Gavinsky:2007:ESO

[GKK^{+}07] Dmitry Gavinsky, Julia Kempe, Iordanis Kerenidis, Ran Raz, and Ronald de Wolf. Exponential separations for oneway quantum communication complexity, with applications to cryptography. In ACM [ACM07], pages 516–525. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Gopalan:2008:ALD

[GKK08] Parikshit Gopalan, Adam Tauman Kalai, and Adam R. Klivans. Agnostically learning decision trees. In ACM [ACM08], pages 527–536. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Goldberg:2002:CCC

[GKP02] Leslie Ann Goldberg, Steven Kelk, and Mike Paterson. The complexity of choosing an Hcolouring (nearly) uniformly at random. In ACM [ACM02], pages 53–62. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Gupta:2003:SBA

[GKR03] Anupam Gupta, Amit Kumar, and Tim Roughgarden. Simpler and better approximation algorithms for network design. In ACM [ACM03], pages 365–372. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Goldwasser:2008:DCI

[GKR08] Shafi Goldwasser, Yael Tauman Kalai, and Guy N. Rothblum. Delegating computation: interactive proofs for muggles. In ACM [ACM08], pages 113–122. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Gupta:2009:OSS

[GKR09] Anupam Gupta, Ravishankar Krishnaswamy, and R. Ravi. Online and stochastic survivable network design. In ACM [ACM09], pages 685–694. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Gavinsky:2006:BEQ

[GKRdW06] Dmitry Gavinsky, Julia Kempe, Oded Regev, and Ronald de Wolf. Bounded-error quantum state identification and exponential separations in communication complexity. In ACM [ACM06], pages 594–603. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Guha:2001:DSH

[GKS01] Sudipto Guha, Nick Koudas, and Kyuseok Shim. Datastreams and histograms. In ACM [ACM01], pages 471–475. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Golin:2002:HCU

[GKY02] Mordecai J. Golin, Claire Kenyon, and Neal E. Young. Huffman coding with unequal letter costs. In ACM [ACM02], pages 785–791. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Gopalan:2008:LDR

[GKZ08] Parikshit Gopalan, Adam R. Klivans, and David Zuckerman. List-decoding Reed– Muller codes over small fields. In ACM [ACM08], pages 265– 274. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Gairing:2004:CNE

[GLMM04] Martin Gairing, Thomas Lücking, Marios Mavronicolas, and Burkhard Monien. Computing Nash equilibria for scheduling on restricted parallel links. In ACM [ACM04], pages 613–622. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Guha:2007:AAB

[GM07] Sudipto Guha and Kamesh Munagala. Approximation algorithms for budgeted learning problems. In ACM [ACM07], pages 104–113. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Guha:2001:CFA

[GMM01] Sudipto Guha, Adam Meyerson, and Kamesh Munagala. A constant factor approximation for the single sink edge installation problems. In ACM [ACM01], pages 383–388. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Goel:2000:CFT

[GMP00] Ashish Goel, Adam Meyerson, and Serge Plotkin. Combining fairness with throughput: online routing with multiple objectives. In ACM [ACM00], pages 670–679. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p670-goel/p670-goel.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p670-goel/. ACM or-

der number 508000.

Gama:2008:FSL

[GN08] Nicolas Gama and Phong Q.

Nguyen. Finding short lattice vectors within Mordell’s inequality. In ACM [ACM08], pages 207–216. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Goldreich:2002:CZK

[Gol02] Oded Goldreich. Concurrent zero-knowledge with timing, revisited. In ACM [ACM02], pages 332–340. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Goldwasser:2009:ALC

[Gol09] Shafi Goldwasser. Athena lecture: Controlling access to programs? In ACM [ACM09], pages 167–168. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Goncalves:2005:EPP

[Gon05] Daniel Gonçalves. Edge partition of planar graphs into two outerplanar graphs. In ACM [ACM05], pages 504–512. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Goyal:2008:VCT

[GOS08] Navin Goyal, Neil Olver, and F. B. Shepherd. The VPN conjecture is true. In ACM [ACM08], pages 443–450. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Goldberg:2006:RAE

[GP06] Paul W. Goldberg and Christos H. Papadimitriou. Reducibility among equilibrium problems. In ACM [ACM06], pages 61–70. ISBN 1-59593-134-1. LCCN QA75.5 .A22

                                        2006. URL http://portal.
                                        acm.org/citation.cfm?id=

1132516. ACM order number 508060.

Gupta:2004:BSA

[GPRS04] Anupam Gupta, Martin Pál, R. Ravi, and Amitabh Sinha. Boosted sampling: approximation algorithms for stochastic optimization. In ACM [ACM04], pages 417–426. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Gentry:2008:THL

[GPV08] Craig Gentry, Chris Peikert, and Vinod Vaikuntanathan.

Trapdoors for hard lattices and new cryptographic constructions. In ACM [ACM08], pages 197–206. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Gal:2003:LBA

[GR03] Anna Gal and Adi Rosen.

Lower bounds on the amount of randomness in private computation. In ACM [ACM03], pages 659–666. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Guruswami:2005:LLD

[GR05] Venkatesan Guruswami and Atri Rudra. Limits to list decoding Reed–Solomon codes. In ACM [ACM05], pages 602–609. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Guruswami:2006:ECA

[GR06] Venkatesan Guruswami and Atri Rudra. Explicit capacityachieving list-decodable codes. In ACM [ACM06], pages 1– 10. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Guruswami:2007:QPI

[GR07] Venkatesan Guruswami and Prasad Raghavendra. A 3query PCP over integers. In ACM [ACM07], pages 198–206. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Goldreich:2009:POT

[GR09] Oded Goldreich and Dana Ron. On proximity oblivious testing. In ACM [ACM09], pages 141–150. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Goel:2004:STM

[GRK04] Ashish Goel, Sanatan Rai, and Bhaskar Krishnamachari. Sharp thresholds for monotone properties in random geometric graphs. In ACM [ACM04], pages 580–586. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Grohe:2000:ITE

[Gro00a] Martin Grohe. Isomorphism testing for embeddable graphs through definability. In ACM [ACM00], pages 63–72. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p63-grohe/p63-grohe.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p63-grohe/. ACM or-

der number 508000.

Grover:2000:RST

[Gro00b] Lov K. Grover. Rapid sampling though quantum computing. In ACM [ACM00], pages 618–626. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p618-grover/p618-grover.
pdf;   http://www.acm.org/
pubs/citations/proceedings/

stoc/335305/p618-grover/.

ACM order number 508000.

Grohe:2001:CCN

[Gro01] Martin Grohe. Computing crossing numbers in quadratic time. In ACM [ACM01], pages 231–236. ISBN 1-58113-349-

9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Georgiou:2003:WCS

[GRS03] Chryssis Georgiou, Alexander Russell, and Alex A. Shvartsman. Work-competitive scheduling for cooperative computing with dynamic groups. In ACM [ACM03], pages 251–258. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Ghosh:2009:UUM

[GRS09] Arpita Ghosh, Tim Roughgarden, and Mukund Sundararajan. Universally utilitymaximizing privacy mechanisms. In ACM [ACM09], pages 351–360. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Goldsmith:2000:MTR

[GS00a] Judy Goldsmith and Robert H.

Sloan. More theory revision with queries (extended abstract). In ACM [ACM00], pages 441–448. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p441-goldsmith/p441-goldsmith.
pdf;   http://www.acm.org/
pubs/citations/proceedings/

stoc/335305/p441-goldsmith/

. ACM order number 508000.

Guruswami:2000:LDA

[GS00b] Venkatesan Guruswami and Madhu Sudan. List decoding algorithms for certain concatenated codes. In ACM [ACM00], pages 181–190. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p181-guruswami/p181-guruswami.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p181-guruswami/

. ACM order number 508000.

Gurvits:2000:DPT

[GS00c] Leonid Gurvits and Alex Samorodnitsky. A deterministic polynomial-time algorithm for approximating mixed discriminant and mixed volume. In ACM [ACM00], pages 48–57. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p48-gurvits/p48-gurvits.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p48-gurvits/.

ACM order number 508000.

Groher:2001:WEC

[GSS01] Martin Groher, Thomas Schwentick, and Luc Segoufin. When is the evaluation of conjunctive queries tractable? In ACM [ACM01], pages 657–666. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Gartner:2001:OLE

[GST^{+}01] Bernd Gärtner, József Solymosi, Falk Tschirschnitz, Emo Welzl, and Pavel Valtr. One line and ϵ. In ACM [ACM01], pages 306–315. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Gilbert:2007:OSA

[GSTV07] A. C. Gilbert, M. J. Strauss, J. A. Tropp, and R. Vershynin. One sketch for all: fast algorithms for compressed sensing. In ACM [ACM07], pages 237–246. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Grigni:2001:QMA

[GSVV01] Michelangelo Grigni, Leonard Schulman, Monica Vazirani, and Umesh Vazirani. Quantum mechanical algorithms for the nonabelian hidden subgroup problem. In ACM [ACM01], pages 68–74. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Gupta:2000:CFA

[GT00] Anupam Gupta and Éva Tardos. A constant factor approximation algorithm for a class of classification problems. In ACM [ACM00], pages 652–658. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p652-gupta/p652-gupta.pdf;
 http://www.acm.org/pubs/
citations/proceedings/stoc/

                                        335305/p652-gupta/. ACM

order number 508000.

Guruswami:2002:LLD

[Gur02] Venkatesan Guruswami. Limits to list decodability of linear codes. In ACM [ACM02], pages 802–811. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Gurvits:2003:CDC

[Gur03] Leonid Gurvits. Classical deterministic complexity of Edmonds’ Problem and quantum entanglement. In ACM [ACM03], pages 10–19. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Guruswami:2004:BEB

[Gur04] Venkatesan Guruswami. Better extractors for better codes? In ACM [ACM04], pages 436–444. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Gurvits:2006:HPA

[Gur06] Leonid Gurvits. Hyperbolic polynomials approach to Van der Waerden/Schrijver–Valiant like conjectures: sharper bounds, simpler proofs and algorithmic applications. In ACM [ACM06], pages 417–426. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

                                        //portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Guruswami:2009:AAC

[Gur09] Venkatesan Guruswami. Artin automorphisms, cyclotomic function fields, and folded listdecodable codes. In ACM [ACM09], pages 23–32. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Gutierrez:2000:SEF

[Gut00]  Claudio Gutiérrez.  Satisfi-

ability of equations in free groups is in PSPACE. In ACM [ACM00], pages 21–27. ISBN 1-58113-184-4. URL

http://www.acm.org/pubs/
articles/proceedings/stoc/
335305/p21-gutierrez/p21-
gutierrez.pdf;   http://
www.acm.org/pubs/citations/
proceedings/stoc/335305/
p21-gutierrez/. ACM order

number 508000.

Grossi:2000:CSA

[GV00] Roberto Grossi and Jeffrey Scott Vitter. Compressed suffix arrays and suffix trees with applications to text indexing and string matching (extended abstract). In ACM [ACM00], pages 397–406. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p397-grossi/p397-grossi.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p397-grossi/.

ACM order number 508000.

Goemans:2001:AAM

[GW01] Michel X. Goemans and David Williamson. Approximation algorithms for MAX-3-CUT and other problems via complex semidefinite programming. In ACM [ACM01], pages 443–452. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Gutoski:2007:TGT

[GW07] Gus Gutoski and John Watrous. Toward a general theory of quantum games. In ACM [ACM07], pages 565–574. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Gao:2003:WSP

[GZ03] Jie Gao and Li Zhang. Wellseparated pair decomposition for the unit-disk graph metric and its applications. In ACM [ACM03], pages 483–492. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Hallgren:2002:PTQ

[Hal02] Sean Hallgren. Polynomialtime quantum algorithms for Pell’s equation and the principal ideal problem. In ACM [ACM02], pages 653–658. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Hallgren:2005:FQA

[Hal05] Sean Hallgren. Fast quantum algorithms for computing the unit group and class group of a number field. In ACM [ACM05], pages 468–474. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Han:2002:DST

[Han02] Yijie Han. Deterministic sorting in O(n log log n) time and linear space. In ACM [ACM02], pages 602–608. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Haastad:2005:ECA

[Hås05] Johan Håstad. Every 2-CSP allows nontrivial approximation. In ACM [ACM05], pages 740–746. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Haussler:2008:CHW

[Hau08] David Haussler. Computing how we became human. In ACM [ACM08], pages 639–640. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Hayes:2003:RCG

[Hay03] Thomas P. Hayes. Randomly coloring graphs of girth at least five. In ACM [ACM03], pages 269–278. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Halperin:2003:PI

[HK03] Eran Halperin and Robert Krauthgamer. Polylogarithmic inapproximability. In ACM [ACM03], pages 585–594. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Holmerin:2004:NPO

[HK04] Jonas Holmerin and Subhash Khot. A new PCP outer verifier with applications to homogeneous linear equations and max-bisection. In ACM [ACM04], pages 11–20. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Halevi:2001:PAN

[HKKN01] Shai Halevi, Robert Krauthgamer, Eyal Kushilevitz, and Kobbi Nissim. Private approximation of NP-hard functions. In ACM [ACM01], pages 550–559. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Hajiaghayi:2005:ORD

[HKLR05] MohammadTaghi Hajiaghayi, Jeong Han Kim, Tom Leighton, and Harald Räcke. Oblivious routing in directed graphs with random demands. In ACM [ACM05], pages 193–201. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Hariharan:2007:MGH

[HKPB07] Ramesh Hariharan, Telikepalli Kavitha, Debmalya Panigrahi, and Anand Bhalgat. An Õ(mn) Gomory–Hu tree construction algorithm for unweighted graphs. In ACM [ACM07], pages 605–614. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Hoyer:2007:NWM

[HLS07] Peter Hoyer, Troy Lee, and Robert Spalek. Negative weights make adversaries stronger. In ACM [ACM07], pages 526– 535. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Hart:2007:CCU

[HM07] Sergiu Hart and Yishay Mansour. The communication complexity of uncoupled Nash equilibrium procedures. In ACM [ACM07], pages 345–353. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Hallgren:2006:LQC

[HMR^{+}06] Sean Hallgren, Cristopher Moore, Martin Rötteler, Alexander Russell, and Pranab Sen. Limitations of quantum coset states for graph isomorphism. In ACM [ACM06], pages 604–617. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Harnik:2004:CTP

[HNRR04] Danny Harnik, Moni Naor, Omer Reingold, and Alon Rosen. Completeness in twoparty secure computation: a computational view. In ACM [ACM04], pages 252–261. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Holmerin:2002:VCR

[Hol02] Jonas Holmerin. Vertex cover on 4-regular hyper-graphs is hard to approximate within 2 − ϵ. In ACM [ACM02], pages 544–552. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Holenstein:2005:KAW

[Hol05] Thomas Holenstein. Key agreement from weak bit agreement. In ACM [ACM05], pages 664–673. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Holenstein:2007:PRS

[Hol07] Thomas Holenstein. Parallel repetition: simplifications and the no-signaling case. In ACM [ACM07], pages 411–419. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Har-Peled:2004:CKM

[HPM04] Sariel Har-Peled and Soham Mazumdar. On coresets for k-means and k-median clustering. In ACM [ACM04], pages 291–300. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Heggernes:2007:ICF

[HPTV07] Pinar Heggernes, Christophe Paul, Jan Arne Telle, and Yngve Villanger. Interval completion with few edges. In ACM [ACM07], pages 374–381. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Harnik:2000:HLB

[HR00] Danny Harnik and Ran Raz. Higher lower bounds on monotone size. In ACM [ACM00], pages 378–387. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/

p378-harnik/p378-harnik.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p378-harnik/.

ACM order number 508000. Hellerstein:2002:ELD

[HR02] Lisa Hellerstein and Vijay Raghavan. Exact learning of DNF formulas using DNF hypotheses. In ACM [ACM02], pages 465–473. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Haitner:2007:SHC

[HR07a] Iftach Haitner and Omer Reingold. Statistically-hiding commitment from any one-way function. In ACM [ACM07], pages 1–10. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Haviv:2007:TBH

[HR07b] Ishay Haviv and Oded Regev.

Tensor-based hardness of the shortest vector problem to within almost polynomial factors. In ACM [ACM07], pages 469–477. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Hartline:2008:OMD

[HR08] Jason D. Hartline and Tim Roughgarden. Optimal mechanism design and money burning. In ACM [ACM08], pages 75–84. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Hallgren:2000:NSR

[HRTS00] Sean Hallgren, Alexander Russell, and Amnon Ta-Shma.

Normal subgroup reconstruction and quantum computation using group representations. In ACM [ACM00], pages 627–635. ISBN 1-58113-184-

4. URL http://www.acm.org/
pubs/articles/proceedings/
stoc/335305/p627-hallgren/
p627-hallgren.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/335305/
p627-hallgren/. ACM order

number 508000.

Haitner:2009:IE

[HRVW09] Iftach Haitner, Omer Reingold, Salil Vadhan, and Hoeteck Wee. Inaccessible entropy. In ACM [ACM09], pages 611–620. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Halpern:2004:RSS

[HT04] Joseph Halpern and Vanessa Teague. Rational secret sharing and multiparty computation: extended abstract. In ACM [ACM04], pages 623–632. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Hayrapetyan:2006:ECC

[HTW06] Ara Hayrapetyan, Éva Tardos, and Tom Wexler. The effect of collusion in congestion games. In ACM [ACM06], pages 89– 98. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Haastad:2002:ARA

[HV02] Johan Håstad and S. Venkatesh.

On the advantage over a random assignment. In ACM [ACM02], pages 43–52. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Healy:2004:UNA

[HVV04] Alexander Healy, Salil Vadhan, and Emanuele Viola. Using nondeterminism to amplify hardness. In ACM [ACM04], pages 192–201. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Hayes:2007:RCP

[HVV07] Thomas P. Hayes, Juan C.

Vera, and Eric Vigoda. Randomly coloring planar graphs with fewer colors than the maximum degree. In ACM [ACM07], pages 450–458. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Iwata:2000:CSP

[IFF00] Satoru Iwata, Lisa Fleischer, and Satoru Fujishige. A combinatorial, strongly polynomialtime algorithm for minimizing submodular functions. In ACM [ACM00], pages 97–106. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p97-iwata/p97-iwata.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p97-iwata/. ACM or-

der number 508000.

Icking:2001:TBC

[IH01] Christian Icking and Lihong Ha. A tight bound for the complexity of Voroni diagrams under polyhedral convex distance functions in 3D. In ACM [ACM01], pages 316–321. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Impagliazzo:2008:UDP

[IJKW08] Russell Impagliazzo, Ragesh Jaiswal, Valentine Kabanets, and Avi Wigderson. Uniform direct product theorems: simplified, optimized, and derandomized. In ACM [ACM08], pages 579–588. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Impagliazzo:2009:AAA

[IKK09] Russell Impagliazzo, Valentine Kabanets, and Antonina Kolokolova. An axiomatic approach to algebrization. In ACM [ACM09], pages 695–704. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Ishai:2006:BBC

[IKLP06] Yuval Ishai, Eyal Kushilevitz, Yehuda Lindell, and Erez Petrank. Black-box constructions for secure computation. In ACM [ACM06], pages 99–108. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

                                        //portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Kawarabayashi:2006:ALC

[iKM06] Ken ichi Kawarabayashi and Bojan Mohar. Approximating the list-chromatic number and the chromatic number in minor-closed and odd-minor-closed classes of graphs. In ACM [ACM06], pages 401–416. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Kawarabayashi:2008:GMI

[iKM08] Ken ichi Kawarabayashi and Bojan Mohar. Graph and map isomorphism and all polyhedral embeddings in linear time. In ACM [ACM08], pages 471–480. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Ishai:2004:BCT

[IKOS04] Yuval Ishai, Eyal Kushilevitz, Rafail Ostrovsky, and Amit Sahai. Batch codes and their applications. In ACM [ACM04], pages 262–271. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Ishai:2007:ZKS

[IKOS07] Yuval Ishai, Eyal Kushilevitz, Rafail Ostrovsky, and Amit Sahai. Zero-knowledge from secure multiparty computation. In ACM [ACM07], pages 21– 30. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Ishai:2008:CCC

[IKOS08] Yuval Ishai, Eyal Kushilevitz, Rafail Ostrovsky, and Amit Sahai. Cryptography with constant computational overhead. In ACM [ACM08], pages 433–442. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Kawarabayashi:2007:CCN

[iKR07] Ken ichi Kawarabayashi and Buce Reed. Computing crossing number in linear time. In ACM [ACM07], pages 382–390. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Kawarabayashi:2009:HCD

[iKR09] Ken ichi Kawarabayashi and Bruce Reed. Hadwiger’s conjecture is decidable. In ACM [ACM09], pages 445–454. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Impagliazzo:2009:NDP

[IKW09] Russell Impagliazzo, Valentine Kabanets, and Avi Wigderson. New direct-product testers and 2-query PCPs. In ACM [ACM09], pages 131–140. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Impagliazzo:2006:CER

[Imp06] Russell Impagliazzo. Can every randomized algorithm be derandomized? In ACM [ACM06], pages 373–374. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

                                        /portal.acm.org/citation.

cfm?id=1132516. ACM order

number 508060.

Indyk:2004:ADG

[Ind04] Piotr Indyk. Algorithms for dynamic geometric problems over data streams. In ACM [ACM04], pages 373–380. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Indyk:2007:UPE

[Ind07] Piotr Indyk. Uncertainty principles, extractors, and explicit embeddings of l_{2} into l_{1}. In ACM [ACM07], pages 615–620. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Istrail:2000:SMT

[Ist00] Sorin Istrail. Statistical mechanics, three-dimensionality and NP-completeness: I. Universality of intracatability for the partition function of the Ising model across nonplanar surfaces (extended abstract). In ACM [ACM00], pages 87–96. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p87-istrail/p87-istrail.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p87-istrail/.

ACM order number 508000.

Impagliazzo:2000:EPR

[ISW00] Russell Impagliazzo, Ronen Shaltiel, and Avi Wigderson. Extractors and pseudorandom generators with optimal seed length. In ACM [ACM00], pages 1–10. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p1-impagliazzo/p1-impagliazzo.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/335305/p1-impagliazzo/

. ACM order number 508000.

Itoh:2003:SSR

[ITT03] Toshiya Itoh, Yoshinori Takei, and Jun Tarui. On the sample size of k-restricted min-wise independent permutations and other k-wise distributions. In ACM [ACM03], pages 710–719. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Indyk:2005:OAF

[IW05] Piotr Indyk and David Woodruff. Optimal approximations of the frequency moments of data streams. In ACM [ACM05], pages 202–208. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Jayanti:2005:OMW

[Jay05]  Prasad Jayanti.  An opti-

mal multi-writer snapshot algorithm. In ACM [ACM05], pages 723–732. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Jayram:2001:OSA

[JKK^{+}01] T. S. Jayram, Tracy Kimbrel, Robert Krauthgamer, Baruch Schieber, and Maxim Sviridenko. Online server allocation in a server farm via benefit task systems. In ACM [ACM01], pages 540–549. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Jayram:2003:CPL

[JKKR03] T. S. Jayram, Subhash Khot, Ravi Kumar, and Yuval Rabani. Cell-probe lower bounds for the partial match problem. In ACM [ACM03], pages 667–672. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Jain:2008:DPT

[JKN08] Rahul Jain, Hartmut Klauck, and Ashwin Nayak. Direct product theorems for classical communication complexity via subdistribution bounds: extended abstract. In ACM [ACM08], pages 599–608. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Jackson:2002:LBA

[JKS02] Jeffrey C. Jackson, Adam R.

Klivans, and Rocco A. Servedio. Learnability beyond AC^{0}. In ACM [ACM02], pages 776–784. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Jayram:2003:TAI

[JKS03] T. S. Jayram, Ravi Kumar, and D. Sivakumar. Two applications of information complexity. In ACM [ACM03], pages 673–682. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Jia:2005:UAT

[JLN^{+}05] Lujun Jia, Guolong Lin, Guevara Noubir, Rajmohan Rajaraman, and Ravi Sundaram. Universal approximations for TSP, Steiner tree, and set cover. In ACM [ACM05], pages 386–395. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Jain:2002:NGA

[JMS02] Kamal Jain, Mohammad Mahdian, and Amin Saberi. A new greedy approach for facility location problems. In ACM [ACM02], pages 731–740. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Juba:2008:USC

[JS08] Brendan Juba and Madhu Sudan. Universal semantic communication I. In ACM [ACM08], pages 123–132. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Jerrum:2001:PTA

[JSV01] Mark Jerrum, Alistair Sinclair, and Eric Vigoda. A polynomialtime approximation algorithm for the permanent of a matrix with non-negative entries. In ACM [ACM01], pages 712–721. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Jain:2001:AAA

[JV01] Kamal Jain and Vijay Vazirani. Applications of approximation algorithms to cooperative games. In ACM [ACM01], pages 364–372. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Jain:2002:ECA

[JV02] Kamal Jain and Vijay V. Vazirani. Equitable cost allocations via primal-dual-type algorithms. In ACM [ACM02], pages 313–321. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Jain:2007:EGM

[JV07] Kamal Jain and Vijay V. Vazirani. Eisenberg–Gale markets: algorithms and structural properties. In ACM [ACM07], pages 364–373. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Jansen:2005:SPR

[JvS05] Klaus Jansen and Rob van Stee.

On strip packing with rotations. In ACM [ACM05], pages 755–761. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Katz:2007:ABB

[Kat07] Jonathan Katz. On achieving the ’best of both worlds’ in secure multiparty computation. In ACM [ACM07], pages 11– 20. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Kabanets:2000:CMP

[KC00] Valentine Kabanets and Jin-Yi Cai. Circuit minimization problem. In ACM [ACM00], pages 73–79. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p73-kabanets/p73-kabanets.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p73-kabanets/.

ACM order number 508000.

Kerenidis:2003:ELB

[KdW03] Iordanis Kerenidis and Ronald de Wolf. Exponential lower bound for 2-query locally decodable codes via a quantum argument. In ACM [ACM03], pages 106–115. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Kelner:2004:SPE

[Kel04] Jonathan A. Kelner. Spectral partitioning, eigenvalue bounds, and circle packings for graphs of bounded genus. In ACM [ACM04], pages 455–464. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Khot:2002:HRA

[Kho02a] Subhash Khot. Hardness results for approximate hypergraph coloring. In ACM [ACM02], pages 351–359. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Khot:2002:PUP

[Kho02b] Subhash Khot. On the power of unique 2-prover 1-round games. In ACM [ACM02], pages 767–775. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Kabanets:2003:DPI

[KI03] Valentine Kabanets and Russell Impagliazzo. Derandomizing polynomial identity tests means proving circuit lower bounds. In ACM [ACM03], pages 355–364. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Kilian:2000:MGC

[Kil00]  Joe Kilian.  More gen-

eral completeness theorems for secure two-party computation. In ACM [ACM00], pages 316–324. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p316-kilian/p316-kilian.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p316-kilian/.

ACM order number 508000.

Kowalik:2003:SPQ

[KK03] Lukasz Kowalik and Maciej Kurowski. Short path queries in planar graphs in constant time. In ACM [ACM03], pages 143–148. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Kolaitis:2009:RGP

[KK09] Phokion G. Kolaitis and Swastik Kopparty. Random graphs and the parity quantifier. In ACM [ACM09], pages 705–714. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Kempe:2001:SGR

[KKD01] David Kempe, Jon Kleinberg, and Alan Demers. Spatial gossip and resource location protocols. In ACM [ACM01], pages 163–172. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Kempe:2000:CIP

[KKK00] David Kempe, Jon Kleinberg, and Amit Kumar. Connectivity and inference problems for temporal networks. In ACM [ACM00], pages 504–513. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p504-kempe/p504-kempe.pdf;
                                         http://www.acm.org/pubs/
                                        citations/proceedings/stoc/
                                        335305/p504-kempe/. ACM

order number 508000.

Kakade:2007:PGA

[KKL07] Sham M. Kakade, Adam Tauman Kalai, and Katrina Ligett. Playing games with approximation algorithms. In ACM [ACM07], pages 546–555. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Kaplan:2005:LAC

[KKM05] Haim Kaplan, Eyal Kushilevitz, and Yishay Mansour. Learning with attribute costs. In ACM [ACM05], pages 356–365. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Karloff:2006:EDM

[KKMR06] Howard Karloff, Subhash Khot, Aranyak Mehta, and Yuval Rabani. On earthmover distance, metric labeling, and 0extension. In ACM [ACM06], pages 547–556. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Karlin:2001:DTA

[KKR01] Anna R. Karlin, Claire Kenyon, and Dana Randall. Dynamic TCP acknowledgement and other stories about e/(e−1). In ACM [ACM01], pages 502–509. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Karger:2002:RSR

[KL02] David R. Karger and Matthew S.

Levine. Random sampling in residual graphs. In ACM [ACM02], pages 63–66. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Kleinberg:2003:CLB

[KL03a] Robert Kleinberg and Tom Leighton. Consistent load balancing via spread minimization. In ACM [ACM03], pages 565–574. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Krauthgamer:2003:IDG

[KL03b] Robert Krauthgamer and James R.

Lee. The intrinsic dimensionality of graphs. In ACM [ACM03], pages 438–447. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Klauck:2000:QPC

[Kla00] Hartmut Klauck. On quantum and probabilistic communication: Las Vegas and one-way protocols. In ACM [ACM00], pages 644–651. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p644-klauck/p644-klauck.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p644-klauck/.

ACM order number 508000.

Klauck:2003:QTS

[Kla03] Hartmut Klauck. Quantum time-space tradeoffs for sorting. In ACM [ACM03], pages 69–

76. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Kleinberg:2000:SWP

[Kle00] Jon Kleinberg. The smallworld phenomenon: an algorithm perspective. In ACM [ACM00], pages 163–170. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p163-kleinberg/p163-kleinberg.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p163-kleinberg/

. ACM order number 508000.

Klein:2006:SSP

[Kle06] Philip N. Klein. A subset spanner for planar graphs, with application to subset TSP. In ACM [ACM06], pages 749–756. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Kiefer:2007:CNM

[KLE07] Stefan Kiefer, Michael Luttenberger, and Javier Esparza. On the convergence of Newton’s method for monotone systems of polynomial equations. In ACM [ACM07], pages 217–226. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Kesselman:2001:BOM

[KLM^{+}01] Alexander Kesselman, Zvi Lotker, Yishay Mansour, Boaz Patt-Shamir, Baruch Schieber, and Maxim Sviridenko. Buffer overflow management in QoS switches. In ACM [ACM01], pages 520–529. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Kalai:2005:CGC

[KLP05] Yael Tauman Kalai, Yehuda Lindell, and Manoj Prabhakaran. Concurrent general composition of secure protocols in the timing model. In ACM [ACM05], pages 644–653. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Kushilevitz:2006:ITS

[KLR06] Eyal Kushilevitz, Yehuda Lindell, and Tal Rabin. Informationtheoretically secure protocols and security under composition. In ACM [ACM06], pages 109–118. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

                                        /portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Kempe:2004:DAS

[KM04] David Kempe and Frank McSherry. A decentralized algorithm for spectral analysis. In ACM [ACM04], pages 561–568. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Kenyon-Mathieu:2007:HRF

[KMS07] Claire Kenyon-Mathieu and Warren Schudy. How to rank with few errors. In ACM [ACM07], pages 95–103. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Kaplan:2003:DRI

[KMT03] Haim Kaplan, Eyal Molad, and Robert E. Tarjan. Dynamic rectangular intersection with priorities. In ACM [ACM03], pages 639–648. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Kalai:2008:ABP

[KMV08] Adam Tauman Kalai, Yishay Mansour, and Elad Verbin. On agnostic boosting and parity learning. In ACM [ACM08], pages 629–638. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Kortsarz:2004:AAK

[KN04] G. Kortsarz and Z. Nutov. Approximation algorithm for knode connected subgraphs via critical graphs. In ACM [ACM04], pages 138–145. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Kol:2008:GEI

[KN08] Gillat Kol and Moni Naor.

Games for exchanging information. In ACM [ACM08], pages 423–432. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Kannan:2009:RWP

[KN09] Ravi Kannan and Hariharan Narayanan. Random walks on polytopes and an affine interior point method for linear programming. In ACM [ACM09], pages 561–570. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Kao:2000:RPP

[KNT00] Ming-Yang Kao, Andreas Nolte, and Stephen R. Tate. The risk profile problem for stock portfolio optimization (extended abstract). In ACM [ACM00], pages 228–234. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p228-kao/p228-kao.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/

                                        335305/p228-kao/. ACM or-

der number 508000.

Klauck:2001:IQC

[KNTSZ01] Hartmut Klauck, Ashwin Nayak, Amnon Ta-Shma, and David Zuckerman. Interaction in quantum communication and the complexity of set disjointness. In ACM [ACM01], pages 124–133. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Kontogiannis:2002:LBC

[Kon02] Spyros Kontogiannis. Lower bounds & competitive algorithms for online scheduling of unit-size tasks to related machines. In ACM [ACM02], pages 124–133. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Kosaraju:2001:EPS

[Kos01] S. Rao Kosaraju. Euler paths in series parallel graphs. In ACM [ACM01], pages 237–240. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Kilian:2001:CRZ

[KP01] Joe Kilian and Erez Petrank. Concurrent and resettable zeroknowledge in poly-logarithm rounds. In ACM [ACM01], pages 560–569. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Koucky:2005:BDC

[KPT05] Michal Koucký, Pavel Pudlák, and Denis Thérien. Boundeddepth circuits: separating wires from gates. In ACM [ACM05], pages 257–265. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Kleinberg:2009:MUO

[KPT09] Robert Kleinberg, Georgios Piliouras, and Eva Tardos. Multiplicative updates outperform generic no-regret learning in congestion games: extended abstract. In ACM [ACM09], pages 533–542. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Konemann:2000:MDI

[KR00] J. Könemann and R. Ravi.

A matter of degree: improved approximation algorithms for degree-bounded minimum spanning trees. In ACM [ACM00], pages 537– 546. ISBN 1-58113-184-4.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/335305/p537-konemann/
p537-konemann.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/335305/
p537-konemann/. ACM order

number 508000.

Karger:2002:FNN

[KR02] David R. Karger and Matthias Ruhl. Finding nearest neighbors in growth-restricted metrics. In ACM [ACM02], pages 741–750. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Konemann:2003:PDM

[KR03] Jochen Könemann and R. Ravi. Primal-dual meets local search: approximating MST’s with nonuniform degree bounds. In ACM [ACM03], pages 389–395. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Kenyon:2004:LDM

[KRS04] Claire Kenyon, Yuval Rabani, and Alistair Sinclair. Low distortion maps between point sets. In ACM [ACM04], pages 272–280. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Khandekar:2006:GPU

[KRV06] Rohit Khandekar, Satish Rao, and Umesh Vazirani. Graph partitioning using single commodity flows. In ACM [ACM06], pages 385–390. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

                                        /portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Kamp:2006:DES

[KRVZ06] Jesse Kamp, Anup Rao, Salil Vadhan, and David Zuckerman. Deterministic extractors for small-space sources. In ACM [ACM06], pages 691–700. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

                                        //portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Klivans:2001:LDT

[KS01a] Adam R. Klivans and Rocco Servedio. Learning DNF in time. In ACM [ACM01], pages 258–265. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Klivans:2001:REI

[KS01b] Adam R. Klivans and Daniel Spielman. Randomness efficient identity testing of multivariate polynomials. In ACM [ACM01], pages 216–223. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Kalai:2003:BPN

[KS03] Adam Kalai and Rocco A.

Servedio. Boosting in the presence of noise. In ACM [ACM03], pages 195–205. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Kleinberg:2004:UMM

[KS04] Jon Kleinberg and Mark Sandler. Using mixture models for collaborative filtering. In ACM [ACM04], pages 569–578. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Kelner:2006:RPT

[KS06] Jonathan A. Kelner and Daniel A. Spielman. A randomized polynomial-time simplex algorithm for linear programming. In ACM [ACM06], pages 51–60. ISBN 1-59593-134-1. LCCN QA75.5 .A22

                                        2006. URL http://portal.
                                        acm.org/citation.cfm?id=

1132516. ACM order number 508060.

Kaufman:2008:APT

[KS08a] Tali Kaufman and Madhu Sudan. Algebraic property testing: the role of invariance. In ACM [ACM08], pages 403–412. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Khot:2008:HLI

[KS08b] Subhash Khot and Rishi Saket. On hardness of learning intersection of two halfspaces. In ACM [ACM08], pages 345–354. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Karpinski:2009:LTA

[KS09a] Marek Karpinski and Warren Schudy. Linear time approximation schemes for the Gale– Berlekamp game and related minimization problems. In ACM [ACM09], pages 313–322. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Kun:2009:NLA

[KS09b] Gábor Kun and Mario Szegedy. A new line of attack on the dichotomy conjecture. In ACM [ACM09], pages 725–734. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Kaplan:2002:MHB

[KST02] Haim Kaplan, Nira Shafrir, and Robert E. Tarjan. Meldable heaps and Boolean unionfind. In ACM [ACM02], pages 573–582. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Kleinberg:2008:MAB

[KSU08] Robert Kleinberg, Aleksandrs Slivkins, and Eli Upfal. Multiarmed bandits in metric spaces. In ACM [ACM08], pages 681– 690. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Kenyon:2000:PTA

[KSY00] Claire Kenyon, Nicolas Schabanel, and Neal Young. Polynomial-time approximation scheme for data broadcast. In ACM [ACM00], pages 659–666. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p659-kenyon/p659-kenyon.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p659-kenyon/.

ACM order number 508000.

Katz:2000:ELD

[KT00] Jonathan Katz and Luca Trevisan. On the efficiency of local decoding procedures for error-correcting codes. In ACM [ACM00], pages 80–86. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p80-katz/p80-katz.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p80-katz/. ACM or-

der number 508000.

Kleinberg:2008:BOS

[KT08] Jon Kleinberg and Éva Tardos. Balanced outcomes in social exchange networks. In ACM [ACM08], pages 295–304. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Kim:2003:GRR

[KV03] Jeong Han Kim and Van H. Vu. Generating random regular graphs. In ACM [ACM03], pages 213–222. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Kitaev:2000:PAE

[KW00] Alexei Kitaev and John Watrous. Parallelization, amplification, and exponential time simulation of quantum interactive proof systems. In ACM [ACM00], pages 608–617. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p608-kitaev/p608-kitaev.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/335305/p608-kitaev/.

ACM order number 508000.

Kushilevitz:2009:CCC

[KW09] Eyal Kushilevitz and Enav Weinreb. On the complexity of communication complexity. In ACM [ACM09], pages 465–474. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Katz:2000:CCS

[KY00] Jonathan Katz and Moti Yung.

Complete characterization of security notions for probabilistic private-key encryption. In ACM [ACM00], pages 245–254. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p245-katz/p245-katz.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p245-katz/. ACM or-

der number 508000.

Lindell:2003:BCS

[Lin03]  Yehuda Lindell.  Bounded-

concurrent secure two-party computation without setup assumptions. In ACM [ACM03], pages 683–692. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Liu:2009:QAU

[Liu09] Yi-Kai Liu. Quantum algorithms using the curvelet transform. In ACM [ACM09], pages 391–400. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Lindell:2002:CAB

[LLR02] Yehuda Lindell, Anna Lysyanskaya, and Tal Rabin. On the composition of authenticated Byzantine agreement. In ACM [ACM02], pages 514–523. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Lepinksi:2005:CFP

[LMas05] Matt Lepinksi, Silvio Micali, and abhi shelat. Collusion-free protocols. In ACM [ACM05], pages 543–552. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Linial:2002:GED

[LMN02] Nathan Linial, Avner Magen, and Assaf Naor. Girth and Euclidean distortion. In ACM [ACM02], pages 705–711. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Lee:2009:NMS

[LMNS09] Jon Lee, Vahab S. Mirrokni, Viswanath Nagarajan, and Maxim Sviridenko. Nonmonotone submodular maximization under matroid and knapsack constraints. In ACM [ACM09], pages 323–332. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Lovett:2008:ICG

[LMS08] Shachar Lovett, Roy Meshulam, and Alex Samorodnitsky. Inverse conjecture for the Gowers norm is false. In ACM [ACM08], pages 547–556. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Li:2000:NOM

[LMW00] Ming Li, Bin Ma, and Lusheng Wang. Near optimal multiple alignment within a band in polynomial time. In ACM [ACM00], pages 425– 434. ISBN 1-58113-184-4.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/335305/p425-li/p425-
li.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/335305/p425-li/. ACM

order number 508000.

Lau:2007:SND

[LNSS07] Lap Chi Lau, Joseph (Seffi)

Naor, Mohammad R. Salavatipour, and Mohit Singh. Survivable network design with degree or order constraints. In ACM [ACM07], pages 651–660. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Lovett:2008:UPG

[Lov08] Shachar Lovett. Unconditional pseudorandom generators for low degree polynomials. In ACM [ACM08], pages 557–562. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Lin:2009:NMA

[LP09] Huijia Lin and Rafael Pass.

Non-malleability amplification. In ACM [ACM09], pages 189–

198. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Lin:2009:UFC

[LPV09] Huijia Lin, Rafael Pass, and Muthuramakrishnan Venkitasubramaniam. A unified framework for concurrent security: universal composability from stand-alone non-malleability. In ACM [ACM09], pages 179–188. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Lachish:2001:ELB

[LR01] Oded Lachish and Ran Raz. Explicit lower bound of 4.5n − o(n) for Boolean circuits. In ACM [ACM01], pages 399–408. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Levi:2004:PDA

[LRS04] Retsef Levi, Robin Roundy, and David B. Shmoys. Primal-dual algorithms for deterministic inventory problems. In ACM [ACM04], pages 353–362. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Levi:2006:PNO

[LRS06] Retsef Levi, Robin O. Roundy, and David B. Shmoys. Provably near-optimal samplingbased algorithms for stochastic inventory control models. In ACM [ACM06], pages 739–748. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

                                        //portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Lu:2003:EOC

[LRVW03] Chi-Jen Lu, Omer Reingold, Salil Vadhan, and Avi Wigderson. Extractors: optimal up to constant factors. In ACM [ACM03], pages 602–611. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Linial:2007:LBC

[LS07] Nati Linial and Adi Shraibman. Lower bounds in communication complexity based on factorization norms. In ACM [ACM07], pages 699–708. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Lau:2008:AAB

[LS08] Lap Chi Lau and Mohit Singh.

Additive approximation for bounded degree survivable network design. In ACM [ACM08], pages 759–768. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Lattanzi:2009:AN

[LS09a] Silvio Lattanzi and D. Sivakumar. Affiliation networks. In ACM [ACM09], pages 427–434. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Lee:2009:GGF

[LS09b] James R. Lee and Anastasios Sidiropoulos. On the geometry of graphs with a forbidden minor. In ACM [ACM09], pages 245–254. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Levin:2008:IRG

[LSZ08] Hagay Levin, Michael Schapira, and Aviv Zohar. Interdomain routing and games. In ACM [ACM08], pages 57–66. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Lovasz:2004:HRC

[LV04] László Lovász and Santosh Vempala. Hit-and-run from a corner. In ACM [ACM04], pages 310–314. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Lynch:2007:DCT

[Lyn07] Nancy A. Lynch. Distributed computing theory: algorithms, impossibility results, models, and proofs. In ACM [ACM07], page 247. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Meir:2008:CCL

[Mei08] Or Meir. Combinatorial construction of locally testable codes. In ACM [ACM08], pages 285–294. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Meyerson:2001:PEF

[Mey01] Adam Meyerson. Profit-earning facility location. In ACM [ACM01], pages 30–36. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Micciancio:2002:ICH

[Mic02] Daniele Micciancio. Improved cryptographic hash functions with worst-case/average-case connection. In ACM [ACM02], pages 609–618. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Magniez:2007:SQW

[MNRS07] Frederic Magniez, Ashwin Nayak, Jeremie Roland, and Miklos Santha. Search via quantum walk. In ACM [ACM07], pages 575–584. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Manokaran:2008:SGU

[MNRS08] Rajsekar Manokaran, Joseph (Seffi)

Naor, Prasad Raghavendra, and Roy Schwartz. SDP gaps and UGC hardness for multiway cut, 0-extension, and metric labeling. In ACM [ACM08], pages 11–20. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Mironov:2008:SAE

[MNS08] Ilya Mironov, Moni Naor, and Gil Segev. Sketching in adversarial environments. In ACM [ACM08], pages 651–660. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Manku:2004:KTN

[MNW04] Gurmeet Singh Manku, Moni Naor, and Udi Wieder. Know thy neighbor’s neighbor: the power of lookahead in randomized P2P networks. In ACM [ACM04], pages 54–63. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Molloy:2002:GDC

[Mol02a] Michael Molloy. The Glauber dynamics on colourings of a graph with high girth and maximum degree. In ACM [ACM02], pages 91–98. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Molloy:2002:MTR

[Mol02b] Michael Molloy. Models and thresholds for random constraint satisfaction problems. In ACM [ACM02], pages 209–217. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Morris:2005:MTT

[Mor05] Ben Morris. The mixing time of the Thorp shuffle. In ACM [ACM05], pages 403–412. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Muller-Olm:2001:OSP

[MOS01] Markus Müller-Olm and Helmut Seidl. On optimal slicing of parallel programs. In ACM [ACM01], pages 647–656. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Mossel:2003:LJ

[MOS03] Elchanan Mossel, Ryan O’Donnell, and Rocco P. Servedio. Learning juntas. In ACM [ACM03], pages 206–212. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Moser:2009:CPL

[Mos09] Robin A. Moser. A constructive proof of the Lovász local lemma. In ACM [ACM09], pages 343–350. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Morris:2003:ESM

[MP03] Ben Morris and Yuval Peres. Evolving sets and mixing. In ACM [ACM03], pages 279–286. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Micali:2006:LZK

[MP06] Silvio Micali and Rafael Pass.

Local zero knowledge. In ACM [ACM06], pages 306–315. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Mortensen:2005:DRR

[MPP05] Christian Worm Mortensen, Rasmus Pagh, and Mihai Pătraşcu. On dynamic range reporting in one dimension. In ACM [ACM05], pages 104–111. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Motwani:2000:DAP

[MPSV00] Rajeev Motwani, Rina Panigrahy, Vijay Saraswat, and Suresh Ventkatasubramanian. On the decidability of accessibility problems (extended abstract). In ACM [ACM00], pages 306–315. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p306-motwani/p306-motwani.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p306-motwani/.

ACM order number 508000.

Maciel:2000:NPW

[MPW00] Alexis Maciel, Toniann Pitassi, and Alan R. Woods. A new proof of the weak pigeonhole principle. In ACM [ACM00], pages 368–377. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p368-maciel/p368-maciel.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p368-maciel/.

ACM order number 508000.

Molloy:2001:CGW

[MR01a] Michael Molloy and Bruce Reed. Colouring graphs when the number of colours is nearly the maximum degree. In ACM [ACM01], pages 462–470. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Moss:2001:AAC

[MR01b] Anna Moss and Yuval Rabani. Approximation algorithms for constrained for constrained node weighted Steiner tree problems. In ACM [ACM01], pages 373–382. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Mossel:2005:LNP

[MR05] Elchanan Mossel and Sébastien Roch. Learning nonsingular phylogenies and hidden Markov models. In ACM [ACM05], pages 366–375. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Moshkovitz:2006:SCE

[MR06] Dana Moshkovitz and Ran Raz. Sub-constant error low degree test of almost-linear size. In ACM [ACM06], pages 21–30. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Mossel:2007:SIS

[MR07] Elchanan Mossel and Sebastien Roch. On the submodularity of influence in social networks. In ACM [ACM07], pages 128–134. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Mostefaoui:2001:CIV

[MRR01] Achour Mostefaoui, Sergio Rajsbaum, and Michel Raynal. Conditions on input vectors for consensus solvability in asynchronous distributed systems. In ACM [ACM01], pages 153–162. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Moore:2007:IQS

[MRS07] Cristopher Moore, Alexander Russell, and Piotr Sniady. On the impossibility of a quantum sieve algorithm for graph isomorphism. In ACM [ACM07], pages 536–545. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Muthukrishnan:2000:ANN

[MS00] S. Muthukrishnan and Süleyman Cenk Sahinalp. Approximate nearest neighbors and sequence comparison with block operations. In ACM [ACM00], pages 416– 424. ISBN 1-58113-184-4. URL

http://www.acm.org/pubs/
articles/proceedings/stoc/
335305/p416-muthukrishnan/

                                        p416-muthukrishnan.pdf;
                                        http://www.acm.org/pubs/
                                        citations/proceedings/stoc/
                                        335305/p416-muthukrishnan/

. ACM order number 508000.

Mavronicolas:2001:PSR

[MS01a] Marios Mavronicolas and Paul Spirakis. The price of selfish routing. In ACM [ACM01], pages 510–519. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Montenegro:2001:EIR

[MS01b] Ravi Montenegro and Jung-Bae Son. Edge isoperimetry and rapid mixing on matroids and geometric Markov chains. In ACM [ACM01], pages 704–711. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Martinelli:2009:MTS

[MS09a] Fabio Martinelli and Alistair Sinclair. Mixing time for the solid-on-solid model. In ACM [ACM09], pages 571–580. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Mathieu:2009:SAR

[MS09b] Claire Mathieu and Alistair Sinclair. Sherali–Adams relaxations of the matching polytope. In ACM [ACM09], pages 293–302. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Montenegro:2009:HLD

[MT09] Ravi Montenegro and Prasad Tetali. How long does it take to catch a wild kangaroo? In ACM [ACM09], pages 553–560. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Mahajan:2000:NNA

[MV00] Meena Mahajan and Kasturi R.

Varadarajan. A new NCalgorithm for finding a perfect matching in bipartite planar and small genus graphs (extended abstract). In ACM [ACM00], pages 351–357. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p351-mahajan/p351-mahajan.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p351-mahajan/.

ACM order number 508000.

Meshulam:2002:ESC

[MW02] Roy Meshulam and Avi Wigderson. Expanders from symmetric codes. In ACM [ACM02], pages 669–677. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Nandakumar:2008:EET

[Nan08] Satyadev Nandakumar. An effective ergodic theorem and some applications. In ACM [ACM08], pages 39–44. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Nesetril:2006:LTL

[NdM06] Jaroslav Nešetřil and Patrice Ossona de Mendez. Linear time low tree-width partitions and algorithmic consequences. In ACM [ACM06], pages 391–400.

ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

                                        //portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Nguyen:2009:FEA

[NDT09] Nam H. Nguyen, Thong T. Do, and Trac D. Tran. A fast and efficient algorithm for low-rank approximation of a matrix. In ACM [ACM09], pages 215–224. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Nordstrom:2008:TOS

[NH08] Jakob Nordström and Johan Håstad. Towards an optimal separation of space and length in resolution. In ACM [ACM08], pages 701–710. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Naor:2001:CPP

[NN01] Moni Naor and Kobbi Nissim. Communication preserving protocols for secure function evaluation. In ACM [ACM01], pages 590–599. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Nordstrom:2006:NPM

[Nor06]  Jakob Nordström.  Narrow

proofs may be spacious: separating space and width in resolution. In ACM [ACM06], pages 507–516. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

                                        /portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Naor:2000:PRF

[NRR00] Moni Naor, Omer Reingold, and Alon Rosen. Pseudorandom functions and factoring (extended abstract). In ACM [ACM00], pages 11–20. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p11-naor/p11-naor.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p11-naor/. ACM or-

der number 508000.

Nissim:2007:SSS

[NRS07] Kobbi Nissim, Sofya Raskhodnikova, and Adam Smith. Smooth sensitivity and sampling in private data analysis. In ACM [ACM07], pages 75– 84. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Nayak:2002:CEA

[NS02] Ashwin Nayak and Julia Salzman. On communication over an entanglement-assisted quantum channel. In ACM [ACM02], pages 698–704. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Naor:2005:BML

[NS05] Joseph (Seffi) Naor and Roy Schwartz. Balanced metric labeling. In ACM [ACM05], pages 582–591. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Naor:2001:APH

[NT01] Moni Naor and Vanessa Teague. Anti-persistence: history independent data structures. In ACM [ACM01], pages 492–501. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Nguyen:2006:ZKE

[NV06] Minh-Huyen Nguyen and Salil Vadhan. Zero knowledge with efficient provers. In ACM [ACM06], pages 287–295. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

                                        /portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Obata:2004:AMI

[Oba04]  Kenji Obata.  Approxi-

mate max-integral-flow/min-multicut theorems. In ACM [ACM04], pages 539–545. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

ODonnell:2002:HAW

[O’D02] Ryan O’Donnell. Hardness amplification within NP. In ACM [ACM02], pages 751–760. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

ODonnell:2008:STA

[O’D08] Ryan O’Donnell. Some topics in analysis of Boolean functions. In ACM [ACM08], pages 569–578. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Ostlin:2003:UHC

[OP03] Anna Ostlin and Rasmus Pagh.

Uniform hashing in constant time and linear space. In ACM [ACM03], pages 622–628. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Ostrovsky:2005:LDE

[OR05] Rafail Ostrovsky and Yuval Rabani. Low distortion embeddings for edit distance. In ACM [ACM05], pages 218–224. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Olshevsky:2000:MVP

[OS00] Vadim Olshevsky and Amin Shokrollahi. Matrix-vector product for confluent Cauchylike matrices with application to confluent rational interpolation. In ACM [ACM00], pages 573–581. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p573-olshevsky/p573-olshevsky.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p573-olshevsky/

. ACM order number 508000.

ODonnell:2003:NDB

[OS03] Ryan O’Donnell and Rocco A.

Servedio. New degree bounds for polynomial threshold functions. In ACM [ACM03], pages 325–334. ISBN 1-58113-674-

9. LCCN QA75.5 .A22 2003. ACM order number 508030.

ODonnell:2008:CPP

[OS08] Ryan O’Donnell and Rocco A. Servedio. The Chow Parameters problem. In ACM [ACM08], pages 517–526. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Orlin:2000:OSB

[OSS00] James B. Orlin, Andreas S. Schulz, and Sudipta Sengupta. ϵ-optimization schemes and Lbit precision (extended abstract): alternative perspectives in combinatorial optimization. In ACM [ACM00], pages 565–572. ISBN 1-58113-184-4. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/335305/
                                        p565-orlin/p565-orlin.pdf;
                                         http://www.acm.org/pubs/
                                        citations/proceedings/stoc/
                                        335305/p565-orlin/. ACM

order number 508000.

Orecchia:2008:PGS

[OSVV08] Lorenzo Orecchia, Leonard J. Schulman, Umesh V. Vazirani, and Nisheeth K. Vishnoi. On partitioning graphs via single commodity flows. In ACM [ACM08], pages 461–470. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

ODonnell:2008:OSA

[OW08] Ryan O’Donnell and Yi Wu. An optimal SDP algorithm for max-cut, and equally optimal long code tests. In ACM [ACM08], pages 335–344. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

ODonnell:2009:CHS

[OW09] Ryan O’Donnell and Yi Wu.

Conditional hardness for satisfiable 3-CSPs. In ACM [ACM09], pages 493–502. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Pagh:2001:CPC

[Pag01] Rasmus Pagh. On the cell probe complexity of membership and perfect hashing. In ACM [ACM01], pages 425–432. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Papadimitriou:2001:AGI

[Pap01] Christos Papadimitriou. Algorithms, games, and the Internet. In ACM [ACM01], pages 749–753. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Papadimitriou:2002:JT

[Pap02] Christos H. Papadimitriou. The Joy of Theory. In ACM [ACM02], page 116. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Papadimitriou:2005:CCE

[Pap05] Christos H. Papadimitriou.

Computing correlated equilibria in multi-player games. In ACM [ACM05], pages 49–56. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Pap:2007:SNR

[Pap07] Gyula Pap. Some new results on node-capacitated packing of A-paths. In ACM [ACM07], pages 599–604. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Pass:2004:BCS

[Pas04]  Rafael Pass.  Bounded-

concurrent secure multi-party computation with a dishonest majority. In ACM [ACM04], pages 232–241. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Patrascu:2007:LBD

[Pat07] Mihai Patrascu. Lower bounds for 2-dimensional range counting. In ACM [ACM07], pages 40–46. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Patrascu:2004:LBD

[PD04] Mihai Pătraşcu and Erik D. Demaine. Lower bounds for dynamic connectivity. In ACM [ACM04], pages 546–553. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Peikert:2009:PKC

[Pei09] Chris Peikert. Public-key cryptosystems from the worst-case shortest vector problem: extended abstract. In ACM [ACM09], pages 333–342. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Plaxton:2003:AAH

[Pla03] C. Greg Plaxton. Approximation algorithms for hierarchical location problems. In ACM [ACM03], pages 40–49. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Plandowski:2006:EAS

[Pla06] Wojciech Plandowski. An efficient algorithm for solving word equations. In ACM [ACM06], pages 467–476. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Pagh:2007:LPC

[PPR07] Anna Pagh, Rasmus Pagh, and Milan Ruzic. Linear probing with constant independence. In ACM [ACM07], pages 318–327. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Parnas:2001:TMP

[PR01a] Michal Parnas and Dana Ron.

Testing metric properties. In ACM [ACM01], pages 276–285. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Pitassi:2001:RRL

[PR01b] Toniann Pitassi and Ran Raz. Regular resolution lower bounds for the weak pigeonhole principle. In ACM [ACM01], pages 347–355. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Pass:2005:NIC

[PR05] Rafael Pass and Alon Rosen. New and improved constructions of non-malleable cryptographic protocols. In ACM [ACM05], pages 533–542. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Peikert:2007:LAL

[PR07] Chris Peikert and Alon Rosen. Lattices that admit logarithmic worst-case to average-case connection factors. In ACM [ACM07], pages 478–487. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Prabhakaran:2004:NNS

[PS04] Manoj Prabhakaran and Amit Sahai. New notions of security: achieving universal composability without trusted setup. In ACM [ACM04], pages 242–251. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Panagiotou:2006:APM

[PS06] Konstantinos Panagiotou and Alexander Souza. On adequate performance measures for paging. In ACM [ACM06], pages 487–496. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

                                        /portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Patrascu:2006:TST

[PT06] Mihai Pãtraşcu and Mikkel Thorup. Time-space tradeoffs for predecessor search. In ACM [ACM06], pages 232–240. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Papadimitriou:2000:ATS

[PV00] Christos H. Papadimitriou and Santosh Vempala. On the approximability of the traveling salesman problem (extended abstract). In ACM [ACM00], pages 126– 133. ISBN 1-58113-184-4. URL

http://www.acm.org/pubs/
articles/proceedings/stoc/
335305/p126-papadimitriou/
p126-papadimitriou.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p126-papadimitriou/

. ACM order number 508000.

Pass:2007:EPR

[PV07] Rafael Pass and Muthuramakrishnan Venkitasubramaniam. An efficient parallel repetition theorem for Arthur–Merlin games. In ACM [ACM07], pages 420–429. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Peikert:2008:LTF

[PW08] Chris Peikert and Brent Waters. Lossy trapdoor functions and their applications. In ACM [ACM08], pages 187–196. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Rabinovich:2003:ADE

[Rab03] Yuri Rabinovich. On average distortion of embedding metrics into the line and into L_{1}. In ACM [ACM03], pages 456–462. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Racke:2008:OHD

[Räc08] Harald Räcke. Optimal hierarchical decompositions for congestion minimization in networks. In ACM [ACM08], pages 255–264. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Raghavan:2006:CFW

[Rag06]  Prabhakar Raghavan.  The

changing face of Web search: algorithms, auctions and advertising. In ACM [ACM06], page 129. ISBN 1-59593-134-1. LCCN QA75.5 .A22

                                        2006. URL http://portal.
                                        acm.org/citation.cfm?id=

1132516. ACM order number 508060.

Raghavendra:2008:OAI

[Rag08] Prasad Raghavendra. Optimal algorithms and inapproximability results for every CSP? In ACM [ACM08], pages 245–254. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Rao:2006:ECN

[Rao06] Anup Rao. Extractors for a constant number of polynomially small min-entropy independent sources. In ACM [ACM06], pages 497–506. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Rao:2008:PRP

[Rao08] Anup Rao. Parallel repetition in projection games and a concentration bound. In ACM [ACM08], pages 1–10. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Raz:2002:CMP

[Raz02a] Ran Raz. On the complexity of matrix product. In ACM [ACM02], pages 144–151. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Raz:2002:RLB

[Raz02b] Ran Raz. Resolution lower bounds for the weak pigeonhole principle. In ACM [ACM02], pages 553–562. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Raz:2004:MLF

[Raz04] Ran Raz. Multi-linear formulas for permanent and determinant are of super-polynomial size. In ACM [ACM04], pages 633–641. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Raz:2005:EWR

[Raz05] Ran Raz. Extractors with weak random seeds. In ACM [ACM05], pages 11–20. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Raz:2008:EFL

[Raz08] Ran Raz. Elusive functions and lower bounds for arithmetic circuits. In ACM [ACM08], pages 711–720. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Reed:2000:HTT

[Ree00] Bruce Reed. How tall is a tree? In ACM [ACM00], pages 479–483. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p479-reed/p479-reed.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p479-reed/. ACM or-

der number 508000.

Regev:2003:NLB

[Reg03] Oded Regev. New lattice based cryptographic constructions. In ACM [ACM03], pages 407–416. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Regev:2005:LLE

[Reg05] Oded Regev. On lattices, learning with errors, random linear codes, and cryptography. In ACM [ACM05], pages 84–93. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Reichardt:2004:QAO

[Rei04] Ben W. Reichardt. The quantum adiabatic optimization algorithm and local minima. In ACM [ACM04], pages 502–510.

ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Reingold:2005:USC

[Rei05] Omer Reingold. Undirected STconnectivity in log-space. In ACM [ACM05], pages 376–385. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Rexford:2008:RIR

[Rex08] Jennifer Rexford. Rethinking Internet routing. In ACM [ACM08], pages 55–56. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Rossman:2008:CDC

[Ros08] Benjamin Rossman. On the constant-depth complexity of kclique. In ACM [ACM08], pages 721–730. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Roughgarden:2001:SSS

[Rou01] Tim Roughgarden. Stackelberg scheduling strategies. In ACM [ACM01], pages 104–113. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Roughgarden:2002:PAI

[Rou02] Tim Roughgarden. The price of anarchy is independent of the network topology. In ACM [ACM02], pages 428–437. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Roughgarden:2009:IRP

[Rou09] Tim Roughgarden. Intrinsic robustness of the price of anarchy. In ACM [ACM09], pages 513–522. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Regev:2006:LPN

[RR06] Oded Regev and Ricky Rosen. Lattice problems and norm embeddings. In ACM [ACM06], pages 447–456. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Raz:2001:LBM

[RS01] Ran Raz and Amir Shpilka. Lower bounds for matrix product, in bounded depth circuits with arbitrary gates. In ACM [ACM01], pages 409–418. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Rubinfeld:2005:TMH

[RS05] Ronitt Rubinfeld and Rocco A. Servedio. Testing monotone high-dimensional distributions. In ACM [ACM05], pages 147– 156. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Remy:2006:QPT

[RS06a] Jan Remy and Angelika Steger. A quasi-polynomial time approximation scheme for minimum weight triangulation. In ACM [ACM06], pages 316–325. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Roughgarden:2006:NTO

[RS06b] Tim Roughgarden and Mukund Sundararajan. New trade-offs in cost-sharing mechanisms. In ACM [ACM06], pages 79–88. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Rodl:2007:PTH

[RS07] V. Rödl and M. Schacht. Property testing in hypergraphs and the removal lemma. In ACM [ACM07], pages 488–495. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Reichardt:2008:SPB

[RS08] Ben W. Reichardt and Robert Spalek. Span-program-based quantum algorithm for evaluating formulas. In ACM [ACM08], pages 103–112. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Rabani:2009:ECS

[RS09] Yuval Rabani and Amir Shpilka. Explicit construction of a small epsilon-net for linear threshold functions. In ACM [ACM09], pages 649–658. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Rozenman:2004:NFC

[RSW04] Eyal Rozenman, Aner Shalev, and Avi Wigderson. A new family of Cayley expanders (?). In ACM [ACM04], pages 445–454. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Reingold:2006:PWR

[RTV06] Omer Reingold, Luca Trevisan, and Salil Vadhan. Pseudorandom walks on regular digraphs and the RL vs. L problem. In ACM [ACM06], pages 457–466. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Rasala:2000:SNB

[RW00a] April Rasala and Gordon Wilfong. Strictly non-blocking WDM cross-connects for heterogeneous networks. In ACM [ACM00], pages 514–523. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p514-rasala/p514-rasala.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p514-rasala/.

ACM order number 508000.

Rothemund:2000:PSC

[RW00b] Paul W. K. Rothemund and Erik Winfree. The program-size complexity of self-assembled squares (extended abstract). In ACM [ACM00], pages 459– 468. ISBN 1-58113-184-4. URL

http://www.acm.org/pubs/
articles/proceedings/stoc/
335305/p459-rothemund/p459-
rothemund.pdf;   http://
www.acm.org/pubs/citations/
proceedings/stoc/335305/
p459-rothemund/. ACM order

number 508000.

Rettinger:2003:CCS

[RW03] Robert Rettinger and Klaus Weihrauch. The computational complexity of some Julia sets. In ACM [ACM03], pages 177–185. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Roditty:2004:FDR

[RZ04] Liam Roditty and Uri Zwick.

A fully dynamic reachability algorithm for directed graphs with an almost linear update time. In ACM [ACM04], pages 184–191. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Srinivasan:2002:ORB

[SA02] Anand Srinivasan and James H.

Anderson. Optimal ratebased scheduling on multiprocessors. In ACM [ACM02], pages 189–198. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Samorodnitsky:2007:LDT

[Sam07] Alex Samorodnitsky. Lowdegree tests at large distances. In ACM [ACM07], pages 506– 515. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Santhanam:2007:CLB

[San07]  Rahul Santhanam.  Circuit

lower bounds for Merlin–Arthur classes. In ACM [ACM07], pages 275–283. ISBN 1-59593-631-9. LCCN QA75.5 .A22

Schulman:2000:CEC

[Sch00] Leonard J. Schulman. Clustering for edge-cost minimization (extended abstract). In ACM [ACM00], pages 547– 555. ISBN 1-58113-184-4.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/335305/p547-schulman/
p547-schulman.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/335305/
p547-schulman/. ACM order

number 508000.

Scheideler:2005:HSA

[Sch05] Christian Scheideler. How to spread adversarial nodes?: rotate! In ACM [ACM05], pages 704–713. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Seiden:2000:GGR

[Sei00] Steven S. Seiden. A guessing game and randomized online algorithms. In ACM [ACM00], pages 592–601. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p592-seiden/p592-seiden.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p592-seiden/.

ACM order number 508000.

Sellie:2009:ELR

[Sel09] Linda Sellie. Exact learning of random DNF over the uniform distribution. In ACM [ACM09], pages 45–54. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Shapira:2006:CCT

[Sha06] Asaf Shapira. A combinatorial characterization of the testable graph properties: it’s all about regularity. In ACM [ACM06], pages 251–260. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Shapira:2009:GCT

[Sha09] Asaf Shapira. Green’s conjecture and testing linear-invariant properties. In ACM [ACM09], pages 159–166. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Sherstov:2007:SAD

[She07] Alexander A. Sherstov. Separating AC^{0} from depth-2 majority circuits. In ACM [ACM07], pages 294–301. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Sherstov:2008:PMM

[She08] Alexander A. Sherstov. The pattern matrix method for lower bounds on quantum communication. In ACM [ACM08], pages 85–94. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Shi:2005:TNC

[Shi05] Yaoyun Shi. Tensor norms and the classical communication complexity of nonlocal quantum measurement. In ACM [ACM05], pages 460–467.

ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Shparlinski:2001:SPA

[Shp01] Igor E. Shparlinski. Sparse polynomial approximation in finite fields. In ACM [ACM01], pages 209–215. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Shpilka:2007:IDA

[Shp07] Amir Shpilka. Interpolation of depth-3 arithmetic circuits with two multiplication gates. In ACM [ACM07], pages 284–293. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Sivakumar:2002:ADC

[Siv02] D. Sivakumar. Algorithmic derandomization via complexity theory. In ACM [ACM02], pages 619–626. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Sanjeev:2001:LMA

[SK01] Arora Sanjeev and Ravi Kannan. Learning mixtures of arbitrary Gaussians. In ACM [ACM01], pages 247–257. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Singh:2007:AMB

[SL07] Mohit Singh and Lap Chi Lau. Approximating minimum bounded degree spanning trees to within one of optimal. In ACM [ACM07], pages 661–670.

ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Sly:2009:RPM

[Sly09] Allan Sly. Reconstruction for the Potts model. In ACM [ACM09], pages 581–590. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Smyth:2002:RIT

[Smy02] Clifford Smyth. Reimer’s inequality and Tardos’ conjecture. In ACM [ACM02], pages 218–221. ISBN 1-58113-495-

9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Sima:2001:CCT

[S̊O01] Jirí S̊íma and Pekka Orponen.

Computing with continuoustime Liapunov systems. In ACM [ACM01], pages 722–731. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Srinivasan:2000:VSI

[Sri00] Aravind Srinivasan. The value of strong inapproximability results for clique. In ACM [ACM00], pages 144–152. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p144-srinivasan/p144-srinivasan.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p144-srinivasan/

. ACM order number 508000.

Schaefer:2001:DSG

[SS01] Marcus Schaefer and Daniel Stefankovic. Decidability of string graphs. In ACM [ACM01], pages 241–246. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Saks:2002:SLB

[SS02] Michael Saks and Xiaodong Sun. Space lower bounds for distance approximation in the data stream model. In ACM [ACM02], pages 360–369. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Santha:2004:QCQ

[SS04] Miklos Santha and Mario Szegedy. Quantum and classical query complexities of local search are polynomially related. In ACM [ACM04], pages 494–501. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Spielman:2008:GSE

[SS08] Daniel A. Spielman and Nikhil Srivastava. Graph sparsification by effective resistances. In ACM [ACM08], pages 563–568. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Schaefer:2002:RSG

[SSŠ02a] Marcus Schaefer, Eric Sedgwick, and Daniel Štefankovič. Recognizing string graphs in NP. In ACM [ACM02], pages 1–6. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Scharbrodt:2002:NAC

[SSS02b] Mark Scharbrodt, Thomas Schickinger, and Angelika Steger. A new average case analysis for completion time scheduling. In ACM [ACM02], pages 170–178. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Samorodnitsky:2000:PCN

[ST00] Alex Samorodnitsky and Luca Trevisan. A PCP characterization of NP with optimal amortized query complexity. In ACM [ACM00], pages 191– 199. ISBN 1-58113-184-4. URL

http://www.acm.org/pubs/
articles/proceedings/stoc/
335305/p191-samorodnitsky/
p191-samorodnitsky.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p191-samorodnitsky/

. ACM order number 508000.

Spielman:2001:SAA

[ST01] Daniel Spielman and Shang-Hua Teng. Smoothed analysis of algorithms: why the simplex algorithm usually takes polynomial time. In ACM [ACM01], pages 296–305. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Spielman:2004:NLT

[ST04] Daniel A. Spielman and Shang-Hua Teng. Nearly-linear time algorithms for graph partitioning, graph sparsification, and solving linear systems. In ACM [ACM04], pages 81–90. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Samorodnitsky:2006:GUI

[ST06] Alex Samorodnitsky and Luca Trevisan. Gowers uniformity, influence of variables, and PCPs. In ACM [ACM06], pages 11–20. ISBN 1-59593-134-1. LCCN QA75.5 .A22

                                        2006. URL http://portal.
                                        acm.org/citation.cfm?id=

1132516. ACM order number 508060.

Schoenebeck:2007:TIG

[STT07] Grant Schoenebeck, Luca Trevisan, and Madhur Tulsiani. Tight integrality gaps for Lovász–Schrijver LP relaxations of vertex cover and max cut. In ACM [ACM07], pages 302–310. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Shaltiel:2007:LEU

[SU07] Ronen Shaltiel and Christopher Umans. Low-end uniform hardness vs. randomness tradeoffs for AM. In ACM [ACM07], pages 430–439. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Sanghvi:2005:RCT

[SV05a] Saurabh Sanghvi and Salil Vadhan. The round complexity of two-party random selection. In ACM [ACM05], pages 338–347. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Schmidt:2005:PTQ

[SV05b] Arthur Schmidt and Ulrich Vollmer. Polynomial time quantum algorithm for the computation of the unit group of a number field. In ACM [ACM05], pages 475–480. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Shaltiel:2008:HAP

[SV08a] Ronen Shaltiel and Emanuele Viola. Hardness amplification proofs require majority. In ACM [ACM08], pages 589–598. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Shpilka:2008:ROP

[SV08b] Amir Shpilka and Ilya Volkovich.

Read-once polynomial identity testing. In ACM [ACM08], pages 507–516. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Skopalik:2008:IPN

[SV08c] Alexander Skopalik and Berthold Vöcking. Inapproximability of pure Nash equilibria. In ACM [ACM08], pages 355–364. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Sauerhoff:2003:TST

[SW03] Martin Sauerhoff and Philipp Woelfel. Time-space tradeoff lower bounds for integer multiplication and graphs of arithmetic functions. In ACM [ACM03], pages 186–195. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Shpilka:2004:DHT

[SW04] Amir Shpilka and Avi Wigderson. Derandomizing homomorphism testing in general groups. In ACM [ACM04], pages 427–435. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Szegedy:2006:DPS

[Sze06] Mario Szegedy. The DLT priority sampling is essentially optimal. In ACM [ACM06], pages 150–158. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

                                        /portal.acm.org/citation.
                                        cfm?id=1132516. ACM order

number 508060.

Talwar:2004:BEA

[Tal04] Kunal Talwar. Bypassing the embedding: algorithms for low dimensional metrics. In ACM [ACM04], pages 281–290. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Tardos:2003:OPF

[Tar03] Gábor Tardos. Optimal probabilistic fingerprint codes. In ACM [ACM03], pages 116–125. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Tardos:2004:NG

[Tar04] Éva Tardos. Network games. In ACM [ACM04], pages 341–342. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Thorup:2000:NOF

[Tho00] Mikkel Thorup. Near-optimal fully-dynamic graph connectivity. In ACM [ACM00], pages 343–350. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p343-thorup/p343-thorup.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p343-thorup/.

ACM order number 508000. Thorup:2001:FDM

[Tho01] Mikkel Thorup. Fully-dynamic min-cut. In ACM [ACM01], pages 224–230. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Thorup:2003:IPQ

[Tho03a] Mikkel Thorup. Integer priority queues with decrease key in constant time and the single source shortest paths problem. In ACM [ACM03], pages 149–158. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Thorup:2003:SED

[Tho03b] Mikkel Thorup. Space efficient dynamic stabbing with fast queries. In ACM [ACM03], pages 649–658. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Thorup:2005:WCU

[Tho05] Mikkel Thorup. Worst-case update times for fully-dynamic allpairs shortest paths. In ACM [ACM05], pages 112–119. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Thorup:2008:MWC

[Tho08]  Mikkel Thorup.  Minimum

k-way cuts via deterministic greedy tree packing. In ACM [ACM08], pages 159–166. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Tokuyama:2001:MPO

[Tok01] Takeshi Tokuyama. Minimax parametric optimization problems and multi-dimensional parametric searching. In ACM [ACM01], pages 75–83. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Trakhtenbrot:2001:ACH

[Tra01] Boris Trakhtenbrot. Automata, circuits and hybrids: facets of continuous time. In ACM [ACM01], pages 754–755. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Trevisan:2001:NAR

[Tre01] Luca Trevisan. Non-approximability results for optimization problems on bounded degree instances. In ACM [ACM01], pages 453–461. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Trevisan:2005:UAH

[Tre05] Luca Trevisan. On uniform amplification of hardness in NP.

In ACM [ACM05], pages 31– 38. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Trevisan:2009:MCS

[Tre09] Luca Trevisan. Max cut and the smallest eigenvalue. In ACM [ACM09], pages 263–272. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Trifonov:2005:LLL

[Tri05] Vladimir Trifonov. An O(log n log log n)

space algorithm for undirected st-connectivity. In ACM [ACM05], pages 626–633. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Ta-Shma:2009:SSE

[TS09] Amnon Ta-Shma. Short seed extractors against quantum storage. In ACM [ACM09], pages 401–408. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Ta-Shma:2001:LLC

[TSUZ01] Amnon Ta-Shma, Christopher Umans, and David Zuckerman. Loss-less condensers, unbalanced expanders, and extractors. In ACM [ACM01], pages 143–152. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Ta-Shma:2001:EC

[TSZ01] Amnon Ta-Shma and David Zuckerman. Extractor codes. In ACM [ACM01], pages 193–199. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Tulsiani:2009:CGR

[Tul09] Madhur Tulsiani. CSP gaps and reductions in the Lasserre hierarchy. In ACM [ACM09], pages 303–312. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Tao:2005:RMS

[TV05] Terence Tao and Van Vu. On random ±1 matrices: singularity and determinant. In ACM [ACM05], pages 431–440. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Talwar:2007:BAW

[TW07] Kunal Talwar and Udi Wieder. Balanced allocations: the weighted case. In ACM [ACM07], pages 256–265. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Thorup:2001:ADO

[TZ01] Mikkel Thorup and Uri Zwick. Approximate distance oracles. In ACM [ACM01], pages 183– 192. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Umans:2002:PRG

[Uma02] Christopher Umans. Pseudorandom generators for all hardnesses. In ACM [ACM02], pages 627–634. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

Umans:2008:FPF

[Uma08] Christopher Umans. Fast polynomial factorization and modular composition in small characteristic. In ACM [ACM08], pages 481–490. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Vadhan:2000:TIP

[Vad00] Salil Vadhan. On transformation of interactive proofs that preserve the prover’s complexity. In ACM [ACM00], pages 200–207. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p200-vadhan/p200-vadhan.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p200-vadhan/.

ACM order number 508000. vonAhn:2005:CTP

[vAHL05] Luis von Ahn, Nicholas Hopper, and John Langford. Covert twoparty computation. In ACM [ACM05], pages 513–522. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Valiant:2001:QCC

[Val01] Leslie G. Valiant. Quantum computers that can be simulated classically in polynomial time. In ACM [ACM01], pages 114–123. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Valiant:2008:TSP

[Val08] Paul Valiant. Testing symmetric properties of distributions. In ACM [ACM08], pages 383–392. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Vatan:2001:DFP

[Vat01] Farrokh Vatan. Distribution functions of probabilistic automata. In ACM [ACM01], pages 684–693. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

vanDam:2000:STU

[vDMMS00] Wim van Dam, Frédéic Magniez, Michele Mosca, and Miklos Santha. Self-testing of universal and fault-tolerant sets of quantum gates. In ACM [ACM00], pages 688–696. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p688-van_dam/p688-van_dam.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/335305/p688-van_dam/.

ACM order number 508000.

Viola:2009:BPL

[Vio09] Emanuele Viola. Bit-probe lower bounds for succinct data structures. In ACM [ACM09], pages 475–482. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Vocking:2001:AOP

[Vöc01] Berthold Vöcking. Almost optimal permutation routing on hypercubes. In ACM [ACM01], pages 530–539. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Vondrak:2008:OAS

[Von08] Jan Vondrak. Optimal approximation for the Submodular Welfare Problem in the value oracle model. In ACM [ACM08], pages 67–74. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

Vu:2007:CNR

[VT07] Van H. Vu and Terence Tao.

The condition number of a randomly perturbed matrix. In ACM [ACM07], pages 248–255. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Vu:2005:SNR

[Vu05] V. H. Vu. Spectral norm of random matrices. In ACM [ACM05], pages 423–430. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Vassilevska:2006:FMW

[VW06] Virginia Vassilevska and Ryan Williams. Finding a maximum weight triangle in n^{3−Δ} time, with applications. In ACM [ACM06], pages 225–231. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Vassilevska:2009:FMC

[VW09] Virginia Vassilevska and Ryan Williams. Finding, minimizing, and counting weighted subgraphs. In ACM [ACM09], pages 455–464. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Vassilevska:2007:APB

[VWY07] Virginia Vassilevska, Ryan Williams, and Raphael Yuster. All-pairs bottleneck paths for general graphs in truly subcubic time. In ACM [ACM07], pages 585–589. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Vygen:2000:DMC

[Vyg00] Jens Vygen. On dual minimum cost flow algorithms (extended abstract). In ACM [ACM00], pages 117–125. ISBN 1-58113-184-4. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/335305/
p117-vygen/p117-vygen.pdf;
 http://www.acm.org/pubs/
citations/proceedings/stoc/
335305/p117-vygen/. ACM

order number 508000.

Watrous:2001:QAS

[Wat01] John Watrous. Quantum algorithms for solvable groups. In ACM [ACM01], pages 60– 67. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Watrous:2006:ZKA

[Wat06] John Watrous. Zero-knowledge against quantum attacks. In ACM [ACM06], pages 296–305. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.

cfm?id=1132516. ACM order

number 508060.

Cho:2006:PFD

[wCG06] Sung woo Cho and Ashish Goel. Pricing for fairness: distributed resource allocation for multiple objectives. In ACM [ACM06], pages 197–204. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Wee:2005:OPF

[Wee05] Hoeteck Wee. On obfuscating point functions. In ACM [ACM05], pages 523–532. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

Weitz:2006:CIS

[Wei06] Dror Weitz. Counting independent sets up to the tree threshold. In ACM [ACM06], pages 140–149. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:/

/portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Wigderson:2004:DTB

[Wig04] Avi Wigderson. Depth through breadth, or why should we attend talks in other areas? In ACM [ACM04], page 579. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Wigderson:2009:WLV

[Wig09] Avi Wigderson. The work of Leslie Valiant. In ACM [ACM09], pages 1–2. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Wang:2001:ETE

[WW01] Li-San Wang and Tandy Warnow. Estimating true evolutionary distances between genomes. In ACM [ACM01], pages 637–646. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Wu:2007:PRD

[WZ07] Fang Wu and Li Zhang. Proportional response dynamics leads to market equilibrium. In ACM [ACM07], pages 354–363. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Yao:2001:SPC

[Yao01] Andrew Chi-Chih Yao. Some perspectives on computational complexity. In ACM [ACM01], page 600. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

Yao:2003:PQF

[Yao03] Andrew Chi-Chih Yao. On the power of quantum fingerprinting. In ACM [ACM03], pages 77–81. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Yao:2004:GEQ

[Yao04] Andrew Chi-Chih Yao. Graph entropy and quantum sorting problems. In ACM [ACM04], pages 112–117. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

Yekhanin:2007:TQL

[Yek07] Sergey Yekhanin. Towards 3query locally decodable codes of subexponential length. In ACM [ACM07], pages 266–274. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

Yoshida:2009:ICT

[YYI09] Yuichi Yoshida, Masaki Yamamoto, and Hiro Ito. An improved constant-time approximation algorithm for maximum matchings. In ACM [ACM09], pages 225–234. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Zhang:2006:NUL

[Zha06] Shengyu Zhang. New upper and lower bounds for randomized and quantum local search. In ACM [ACM06], pages 634–643. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.

Zuckerman:2006:LDE

[Zuc06] David Zuckerman. Linear degree extractors and the inapproximability of max clique and chromatic number. In ACM [ACM06], pages 681–690. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http:

//portal.acm.org/citation.
cfm?id=1132516. ACM order

number 508060.